### v1

# AF Embeddability of the C*-Algebra of a Deaconu-Renault Groupoid

Abstract.
We study Deaconu-Renault groupoids corresponding to surjective local homeomorphisms on locally compact, Hausdorff, second countable, totally disconnected spaces, and we characterise when the C*-algebras of these groupoids are AF embeddable. Our main result generalises theorems in the literature for graphs and for crossed products of commutative C*-algebras by the integers. We give a condition on the surjective local homeomorphism that characterises the AF embeddability of the C*-algebra of the associated Deaconu-Renault groupoid. In order to prove our main result, we generalise an application of Matui’s HK conjecture by Farsi, Kumjian, Pask, and Sims, that gives an isomorphism from the homology groups of AF algebras and the corresponding K-theory. In this paper, we give an explicit formula for this isomorphism, and we show that this isomorphism is an order isomorphism.
Key words and phrases:
elementary groupoid, AF groupoid
2020 Mathematics Subject Classification:
22A22(primary), 46L05 (secondary)
This research was funded by the Marsden Fund of the Royal Society of New Zealand (grant number 18-VUW-056)
1. Introduction
A problem that has received much interest during the last decades is the classification of C*-algebras. This problem consists of finding an invariant that identifies all C*-algebras of a certain class up to isomorphism. One of the first results in the area is Glimm’s [22] classification of some uniformly hyperfinite C*-algebras. Glimm’s result was interesting because uniformly hyperfinite algebras are the C*-algebraic version of the $\text{II}_{1}$ factors for von Neumann algebras, but unlike the von Neumann case [42] (see also [12, Theorem V.1.13]), uniformly hyperfinite C*-algebras are not all isomorphic.
Dixmier [15] then generalised this classification by studying a class of C*-algebras called matroid C*-algebras. In [6], Bratteli defined a more general class of C*-algebras, called approximately finite dimensional C*-algebras (or AF algebras). An AF algebra is the completion of an increasing union of finite dimensional C*-algebras. Elliott [19, Theorem 4.3] (see [13, Theorem IV.4.3] for a more modern formulation) classified all AF algebras by their K-theory.
Since Elliott’s classification of AF algebras, different types of C*-algebras have been classified by invariants. Several stably finite C*-algebras [23, 24] are classifiable. Kirchberg [30] and Phillips [43] classified all unital, simple, separable, nuclear and purely infinite C*-algebras (called Kirchberg algebras) satisfying the Universal Coefficient Theorem by their K-theory. See Strung’s book on the classification of C*-algebras [56] for more details.
In certain types of C*-algebras, a dichotomy exists between purely infinite and stably finite. In other words, for certain classes of C*-algebras, a C*-algebra is either stably finite or purely infinite. This dichotomy holds for reduced C*-algebras of certain minimal, topological principal, ample groupoids, Bönicke and Li [5, Theorem D], Rainone and Sims [47, Theorem 7.4], and Ma [37]. Additionally to the classification programme, the possibility of finding a dichotomy for large classes of C*-algebras motivates us to study purely infinite and stably finite C*-algebras, and related properties.
Pimsner and Voiculescu [44] proved that the irrational rotation C*-algebra is embedded into the C*-algebra from [18], which is AF and whose $K_{0}$-group is isomorphic to $\mathbb{Z}+\theta\mathbb{Z}$. This embedding led to the classication of irrational rotation C*-algebras in terms of the angle of rotation. Effros [17] observed the need of investigating subalgebras of AF algebras, since virtually nothing was known about them at the time. Such algebras are denoted to be AF embeddable (or AFE algebras).
Since the formulation of the conjecture that a C*-algebra is AF embeddable if, and only if, it is separable, exact, and quasidiagonal (see [4, Section 7]), the problem of characterising AFE algebras has received much attention. Pimsner [45] characterised the AFE property for crossed products of commutative C*-algebras by the integers. Brown [8] generalised this result to crossed products of AF algebras to the integers, and that the notions of stably finite, quasidiagonal and AF embeddable are equivalent for these crossed products. Brown [7] proved that the crossed product of UHF algebra by $\mathbb{Z}^{k}$ is AF embeddable. Similar results are given in [25, 34, 35, 38]. Liu [36] described the AF embeddability of separable exact quasidiagonal C*-algebras in terms of a condition on the C*-algebras representations. Schafhauser [52, Theorem 6.7] characterised the AF embeddability for topological graph algebras with no sinks, where the notions of stably finite, quasidiagonal and AF embeddable are equivalent. Topological graph algebras are very general and cover the C*-algebras of Deaconu-Renault groupoids, which we study in this paper. So, the equivalence of these notions holds for the C*-algebras we are interested.
In the context of graph algebras, there are several results in the literature studying C*-algebras that are AF, AF embeddable or purely infinite. For example, graph loops characterises when a graph algebra is approximately finite dimensional [32]. Similarly, Schafhauser [54] describes when a graph algebra is AF embeddable, while [3, Proposition 5.3] gives a graph condition that makes the corresponding C*-algebra purely infinite. Szymański [57] showed that every stable, purely infinite, simple and classifiable C*-algebra with $K_{1}$ torsion-free is a graph algebra.
Kumjian and Pask [31] study higher rank graph algebras and give sufficient conditions that make the C*-algebras AF or purely infinite. Clark, an Huef, and Sims [11] characterised the AF embeddability of C*-algebras of row-finite and cofinal $k$-graphs with now sources, for $k=2$. Schafhauser [53, Corollary 6.2] extendded this result for an arbitrary $k$.
In this paper we study C*-algebras of Deaconu-Renault groupoids. These algebras generalise graph algebras [33] and are examples of topological graph algebras [27, Proposition 10.9]. In Subsection 2.1, we describe this class of groupoids in more detail. The main contribution of this paper is the description of the AF embeddability of the C*-algebras of Deaconu-Renault groupoids by a corresponding groupoid property. Our result is analogous to similar results for graph algebras [54], for crossed products of unital commutative C*-algebras by the integers [9, Theorem 11.5], [45, Theorem 9], and for higher rank graphs [11], [53, Corollary 6.2], cited above. In particular, we apply the same strategy from [11], by applying [8, Theorem 0.2] to describe AF embeddability in terms of a K-theory condition, and them analysing this condition to find a more concrete condition which can be described by an equation. We state our main result below.
Theorem 1.
Let $\sigma$ be a surjective local homeomorphism on a locally compact, Hausdorff, second countable, totally disconnected space $X$. Denote by $\mathcal{G}$ be the Deaconu-Renault groupoid corresponding to $\sigma$. Then the following are equivalent:
  1. (i)
    $C^{*}(\mathcal{G})$ is AF embeddable,
  2. (ii)
    $C^{*}(\mathcal{G})$ is quasidiagonal,
  3. (iii)
    $C^{*}(\mathcal{G})$ is stably finite,
  4. (iv)
    $\mathrm{Im}(\sigma_{\ast}-\mathrm{id})\cap C_{c}(X,\mathbb{N})=\{0\}$,
where the map $\sigma_{\ast}:C_{c}(X,\mathbb{Z})\rightarrow C_{c}(X,\mathbb{Z})$ is defined by
$\sigma_{\ast}(f)(x)=\sum_{y:\sigma(y)=x}f(y)\hskip 20.0pt\text{for $x\in X$, $f\in C_{c}(X,\mathbb{Z})$.}$
Note that C*-algebras of Deaconu-Renault groupoids are examples of topological graph algebras. When $X$ is compact, then the corresponding topological graph algebra is unital and, in this case, Theorem 1 is an example of [52, Theorem 6.7], that characterises when a unital topological graph algebra is AF embeddable.
The construction of the Deaconu-Renault groupoid $\mathcal{G}$ depends on the surjective local homeomorphism $\sigma$. The contribution of Theorem 1 is that it provides a condition on $\sigma$ that is equivalent to the AF embeddability of the C*-algebra of the corresponding Deaconu-Renault groupoid.
In order to prove Theorem 1, we apply the following result by Brown [8], that gives an equivalent condition to AF embeddability of crossed products of AF algebras by the integers:
Definition.
If $B$ is an AF algebra and $\beta\in\mathrm{Aut}(B)$, then we denote by $H_{\beta}$ the subgroup of $K_{0}(B)$ given by all elements of the form $K_{0}(\beta)(x)-x$ for $x\in K_{0}(B)$.
Theorem 1.1.
(Brown) [8, Theorem 0.2] If $B$ is an AF algebra and $\beta\in\mathrm{Aut}(B)$, then the following are equivalent:
  1. (a)
    $B\rtimes_{\beta}\mathbb{Z}$ is AF embeddable,
  2. (b)
    $B\rtimes_{\beta}\mathbb{Z}$ is quasidiagonal,
  3. (c)
    $B\rtimes_{\beta}\mathbb{Z}$ is stably finite,
  4. (d)
    $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$.
Note that AF embeddability, quasidiagonality and the stable finite properties are preserved by stable isomorphisms. If we choose an AF algebra $B$ such that $B\rtimes_{\beta}\mathbb{Z}$ is stably isomorphic to $C^{*}(\mathcal{G})$, then we can apply Brown’s theorem to get a characterisation of the AF embeddability of $C^{*}(\mathcal{G})$. Indeed, we let $c:\mathcal{G}\rightarrow\mathbb{Z}$ be the canonical cocycle, and we let $B=C^{*}(\mathcal{G}(c))$ be the C*-algebra of the skew product $\mathcal{G}(c)$. This C*-algebra is AF by [20, Lemma 6.1]. Theorem III.5.7 of [48] gives the isomorphism
$B\cong C^{*}(\mathcal{G})\rtimes_{\alpha^{c}}\mathbb{T},$
where $\alpha^{c}$ is the action of $\mathbb{T}$ on $C^{*}(\mathcal{G})$ corresponding to the cocycle $c$. We let $\beta$ be an automorphism on $B$ corresponding to the dual map $\widehat{\alpha}^{c}$ with respect to the isomorphism above. By applying the crossed products on both sides of the equivalence above with the respective maps, we get the isomorphism $B\rtimes_{\beta}\mathbb{Z}\cong C^{*}(\mathcal{G})\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$, and by Takai duality, we have that $B\rtimes_{\beta}\mathbb{Z}$ is stably isomorphic to $C^{*}(\mathcal{G})$. So, we apply Brown’s theorem to $C^{*}(\mathcal{G})$ and we get the equivalence of items (i)–(iii) of Theorem 1 with item (d) of Brown’s theorem.
This equivalence gives a characterisation of when the C*-algebra is AF embeddable. Note that item (d) depends on the K-theory, which is very abstract. We use the theory of homology groups of groupoids in order to show that (d) is equivalent to condition (iv) of Theorem 1. By applying results from [20, 39] on homology groups, we get the commutative diagram below.
[…]
The existence of the isomorphism $K_{0}(B)\rightarrow H_{0}(\mathcal{G}(c))$ is given by [20, Corollary 5.2], which is an application of Matui’s HK conjecture [39, 40, 41], and is such that $[1_{V}]_{0}\mapsto[1_{V}]$ for every compact open $V\subset\mathcal{G}^{(0)}$. The groupoid $c^{-1}(0)$ is the kernel of the cocycle $c:\mathcal{G}\rightarrow\mathbb{Z}$ defined by $c(x,k,y)=k$. We define the isomorphism $H_{0}(\rho_{\varphi}):H_{0}(\mathcal{G})(c))\rightarrow H_{0}(c^{-1}(0))$ in Section 6. The homomorphisms $K_{0}(\beta)$ and $[\widetilde{\beta}]_{\mathcal{G}(c)}$ are group homomorphisms induced by $\beta$, and the group homomorphism $[\sigma_{\ast}]$ is induced by the map $\sigma_{\ast}$ of Theorem 1.
By studying the diagram and by applying properties of homology groups of groupoids, we can show that items (d) and (iv) are equivalent, proving Theorem 1.
We need to know the positive elements of $K_{0}(B)$ in order to study item (d). So, when we apply the commutative diagram, we need to guarantee that the isomorphisms above takes positive elements from one group to the another. Although [20, Corollary 5.2] gives the existence of the group isomorphism $K_{0}(B)\cong H_{0}(\mathcal{G}(c))$, we need this isomorphism to be an ordered group isomorphism. We prove that this isomorphism is indeed a ordered group isomorphism using different techniques, and we give an explicit formula for the map, stated as follows:
Definition.
Let $\sigma:X\rightarrow Y$ be a surjective local homeomorphism on locally compact, Hausdorff, second countable, totally disconnected spaces, and let $\varphi:Y\rightarrow X$ be a continuous section of $\sigma$. We define the map $\mathrm{tr}_{\varphi}:\mathcal{P}_{\infty}(C^{*}(R(\sigma)))\rightarrow C_{c}(X,\mathbb{N})$ by
  • $\mathrm{tr}_{\varphi}(p)(x)=1_{X_{\varphi}}(x)\displaystyle\sum_{y:\sigma(x)=\sigma(y)}p(y)$,       for $p\in C^{*}(R(\sigma))$, $x\in X$, and
  • $\mathrm{tr}_{\varphi}(p)(x)=\displaystyle\sum_{i=1}^{n}\mathrm{tr}_{\varphi}(p_{ii})(x)$,                  for $p\in\mathcal{P}_{n}(C^{*}(R(\sigma)))$, $n\geq 1$,
where we identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$ by applying the $j$ map from [48, Proposition II.4.2] and [55, Proposition 9.3.3].
Theorem 2.
Let $G$ be an AF groupoid, written as the inductive limit $G=\displaystyle\lim_{\rightarrow}G_{n}$ of elementary groupoids. Then, for every $n$, there exists a locally compact, Hausdorff, second countable, totally disconnected space $Y_{n}$, and there is a surjective local homeomorphism $\sigma_{n}:G_{n}^{(0)}\rightarrow Y_{n}$ such that $G_{n}=R(\sigma_{n})$. Moreover,
  1. 1.
    there exists a unique isomorphism $\nu:K_{0}(C^{*}(G))\rightarrow H_{0}(G)$ defined by
    $\nu([p]_{0})=[\mathrm{tr}_{\varphi_{n}}(p)]$
    for $p\in\mathcal{P}_{n}(C^{*}(G_{n}))$, and such that $\varphi_{n}$ is an arbitrary continuous section of $\sigma_{n}$;
  2. 2.
    the map $\nu$ is an isomorphism of ordered groups, i.e., $\nu(K_{0}(C^{*}(G))^{+})=H_{0}(G)^{+}$; and
  3. 3.
    the isomorphism $\nu$ is precisely the map given by [20, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
This theorem also generalises Matui’s [39, Theorem 4.10], which also gave an order isomorphism for AF groupoids with compact unit spaces, but did not provide a formula for the isomorphism. So, we extend Matui’s result to AF groupoids with non-compact unit spaces. Moreover, by showing that $H_{0}(G)^{+}\cong K_{0}(C^{*}(G))^{+}$, Theorem 2 suggests a stronger version of Matui’s HK conjecture, which considered only group homomorphisms.
We organise the paper as follows: in Section 2, we define Deaconu-Renault groupoids, inductive limits and AF groupoids, presenting some results from the literature. In Section 3, we state the isomorphism by [48, Proposition II.5.1] between the C*-algebra of a skew-product groupoid and its corresponding crossed product. We also define the map $\beta$, used later in the commutative diagram. Then we study the theory of homology groups for groupoids in Section 4. In Section 5, we prove Theorem 2, giving an isomorphism between the homology groups for AF groupoids and the K-theory of the C*-algebras of these groupoids. Similar results have been proved in [20, Corollary 5.2], [39, Theorem 4.10]. Our innovation is that we give an explicit formula for this isomorphism, and we show that it preserves positive elements from one group to another. In Section 6 we apply the results from Sections 4 and 5 to prove the commutative diagram (1.1). In Section 7 we prove Theorem 1, our main result, that describes under certain conditions when the C*-algebra of a Deaconu-Renault groupoid is AF embeddable. Finally, we show in Section 8 that our main theorem generalises some known results in the literature.
The author would like to thank Astrid an Huef and Lisa Orloff Clark for their supervision, Camila F. Sehnem for the helpful discussions, and Nathan Brownlowe, Carla Farsi, and Hung Pham for their suggestions.
2. Preliminaries
2.1. Deaconu-Renault groupoids
Given the topological spaces $X$ and $Y$, a map $\sigma:X\rightarrow Y$ is a local homeomorphism if every $x\in X$ has an open neighbourhood $\mathcal{U}$ such that $\sigma(\mathcal{U})$ is open and $\sigma|_{\mathcal{U}}:\mathcal{U}\rightarrow\sigma(\mathcal{U})$ is a homeomorphism. A continuous map $\varphi:Y\rightarrow X$ is a continuous section of $\sigma$ if $y=\sigma(\varphi(y))$ for all $y$. Using the argument from the proof of [20, Theorem 4.10], one can show that every surjective local homeomorphism has a continuous section.
Let $X$ be a locally compact, Hausdorff, second countable space, and $\sigma:X\rightarrow X$ is a surjective local homeomorphism. We define the set
$\mathcal{G}=\{(x,k,y)\in X\times\mathbb{Z}\times X:\exists m,n\in\mathbb{N},k=m-n,\sigma^{m}(x)=\sigma^{n}(y)\},$
and we endow it with the range and source maps $r,s:\mathcal{G}\rightarrow\mathcal{G}$ with $r(x,k,y)=(x,0,x)$ and $s(x,k,y)=(y,0,y)$. Let $\mathcal{G}^{(2)}=\{((x,k_{1},y),(y,k_{2},z)):(x,k_{1},y),(y,k_{2},z)\in\mathcal{G}\}$ and we define the product $\mathcal{G}^{(2)}\rightarrow\mathcal{G}$ and inverse operations $\mathcal{G}\rightarrow\mathcal{G}$ by
$(x,k_{1},y)(y,k_{2},z)=(x,k_{1}+k_{2},z),\text{\hskip 10.0pt}(x,k,y)^{-1}=(y,-k,x).$
With these operations, $\mathcal{G}$ becomes a groupoid. The unit space is $\mathcal{G}^{(0)}=\{(x,0,x):x\in X\}$, which is identified with $X$ by the map $(x,0,x)\mapsto x$. For the open subsets $A,B\subset X$ and for the natural numbers $m,n$ such that $\sigma|_{A}^{m}$ and $\sigma|_{B}^{n}$ are injective with $\sigma^{m}(A)=\sigma^{n}(B)$, define the subset
$\mathcal{U}_{A,B}^{m,n}=\{(x,m-n,y)\in A\times\mathbb{Z}\times B:\sigma^{m}(x)=\sigma^{n}(y)\}.$
Here we use the notation from [58]. It follows from [14, Theorem 1] that the sets $\mathcal{U}_{A,B}^{m,n}$ generate a topology that makes $\mathcal{G}$ a topological groupoids which is locally compact, Hausdorff, second countable, étale. We denote the groupoid $\mathcal{G}$ endowed with this topology a Deaconu-Renault groupoid.
The first example of this groupoid was described in Renault’s book [48, Section III.2] by a class of groupoids whose C*-algebras are Cuntz algebras, where $X$ is the set of sequences $(x_{i})_{i\in\mathbb{N}}$, with $x_{i}\in\{0,1,\dots,n\}$ (or $x_{i}\in\mathbb{N}$ for the Cuntz algebra $\mathcal{O}_{\infty}$), and $\sigma$ was given by $\sigma(x_{0}x_{1}x_{2}\dots)=x_{1}x_{2}\dots$ Deaconu [14] generalises Renault’s definition by considering compact Hausdorff spaces $X$ and assumes that $\sigma:X\rightarrow X$ is a covering map. Under these assumptions, he proved that when $\sigma$ is a homeomorphism, then $C^{*}(\mathcal{G})$ is isomorphic to the crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$. Anantharaman-Delaroche [1, Example 1.2 (c)] generalised it to the case where $\sigma$ is a surjective local homeomorphism, and $X$ is locally compact. She also gave a sufficient conditions to make the C*-algebra of this groupoid purely infinite [1, Propositions 4.2 and 4.3]. Then Renault [49] generalised this definition by considering a map $\sigma$ that is defined only on an open subset of $X$. For the purpose of our paper, here we consider the definition by Anantharaman-Delaroche, and we impose the additional conditions on the space $X$ to be Hausdorff and second countable.
By [49, Proposition 2.9], the groupoid is also amenable. Note that there are different notions of amenability for groupoids: topological amenable and measurewise amenable. By Theorem 3.3.7 of [2], these two notions coincide for groupoids that are locally compact, Hausdorff and étale. In this case, [2, Proposition 6.1.8] implies that the reduced and full C*-algebras of the groupoid coincide.
2.2. Inductive limits
An inductive sequence in a category $\mathscr{C}$ is a sequence $\{A_{n}\}_{n=1}^{\infty}$ of objects in $\mathscr{C}$ and a sequence $\varphi_{n}:A_{n}\rightarrow A_{n+1}$ of morphisms in $\mathscr{C}$, usually written
[…]
For $m>n$ we shall also consider the composed morphisms
$\varphi_{m,n}=\varphi_{m-1}\circ\varphi_{m-2}\circ\cdots\circ\varphi_{n}:A_{n}\rightarrow A_{m},$
which, together with the morphisms $\varphi_{n}$, are called the connecting morphisms (or connecting maps). An inductive limit of the inductive sequence (2.1) in a category $\mathscr{C}$ is a system $(A,\{\mu_{n}\}_{n=1}^{\infty})$, where $A$ is an object in $\mathscr{C}$, $\mu_{n}:A_{n}\rightarrow A$ is a morphism in $\mathscr{C}$ for each $n\geq 1$, and where the following two conditions hold.
  1. (i)
    The diagram
[…]
    commutes for each positive integer $n$.
  2. (ii)
    If $(B,\{\lambda_{n}\}_{n=1}^{\infty})$ is a system, where $B$ is an object in $\mathscr{C}$, $\lambda_{n}:A_{n}\rightarrow B$ is a morphism in $\mathscr{C}$ for each $n$, and where $\lambda_{n}=\lambda_{n+1}\circ\varphi_{n}$ for all $n\geq 1$, then there is one and only one morphism $\lambda:A\rightarrow B$ making the diagram
[…]
    commutative for each $n$.
In this text, will call the morphisms $\mu_{n}$ the inclusion morphisms of $A$
Inductive limits, when they exist, are essentially unique in the sense that if $(A,\{\mu_{n}\})$ and $(B,\{\lambda_{n}\})$ are inductive limits of the same inductive sequence, then there is an isomorphism $\lambda:A\rightarrow B$ making the diagram of item (ii) above commutative. See [51, page 93] for details. Because of the essential uniqueness of inductive limits (when they exist), we shall refer to the inductive limit (rather than an inductive limit).
Consider an AF algebra $A=\overline{\bigcup_{n=0}^{\infty}A_{n}}$, where each $A_{n}\subset A_{n+1}$ is finite dimensional. Then $A=\displaystyle\lim_{\longrightarrow}A_{n}$ and $K_{0}(A)=\displaystyle\lim_{\longrightarrow}K_{0}(A_{n})$.
Lemma 2.1.
Let $A$ be an object in a category $\mathscr{C}$ such that $A$ is the inductive limit of an inductive sequence with connecting morphisms $i_{n}:A_{n}\rightarrow A_{n+1}$ and inclusion morphisms $\alpha_{n}:A_{n}\rightarrow A$. Similarly, let $B$ be an inductive limit in $\mathscr{C}$ with connecting morphisms $j_{n}:B_{n}\rightarrow B_{n+1}$ and inclusion morphisms $\beta_{n}:B_{n}\rightarrow B$. Let $\{\mu_{n}\}_{n=0}^{\infty}$ be a sequence of morphisms $\mu_{n}:A_{n}\rightarrow B_{n}$ satisfying $\mu_{n+1}\circ i_{n}=j_{n}\circ\mu_{n}$. Then there exists a unique morphism $\mu:A\rightarrow B$ making the diagram below commutative.
[…]
Moreover, if each $\mu_{n}$ is an isomorphism, we have that $\mu$ is an isomorphism.
2.3. AF groupoids
There are different definitions for AF groupoids in the literature. The first was given by Renault [48, Definition III.1.1] in his book, and he proved in [48, Proposition III.1.15] that AF groupoids determine AF algebras up to isomorphism. Because of this result, we can see AF groupoids as the groupoid analogs of AF algebras. Other definitions for such groupoids have been introduced in the literature [50], [21, Definition 3.7], [39, Definition 2.2], and [20, Definition 4.9]. In [10] there is a proof of the equivalence of the definitions [48, Definition III.1.1] and [20, Definition 4.9]. In this paper, we consider the Definition 4.9 of [20], by Farsi, Kumjian, Pask, and Sims.
Definition 2.2.
Let $X,Y$ be locally compact, Hausdorff, second countable, totally disconnected spaces. Let $\sigma:X\rightarrow Y$ be a surjective local homeomorphism. Define the set
$R(\sigma)=\{(u,v)\in X\times X:\sigma(u)=\sigma(v)\},$
and equip it with the operations $(u,v)(v,w)=(u,w)$ and $(u,v)^{-1}=(v,u)$. Then $R(\sigma)$ becomes a groupoid. We endow this set with the subspace topology from $X\times X$. Then $R(\sigma)$ is locally compact, Hausdorff, second countable, étale, and totally disconnected.
Definition 2.3.
We say that a locally compact, Hausdorff, second countable, totally disconnected groupoid is elementary if it is isomorphic to the groupoid $R(\sigma)$ for some surjective local homeomorphism $\sigma:X\rightarrow Y$ between locally compact, Hausdorff, second countable, totally disconnected spaces. A groupoid is AF if it is the increasing union of open elementary groupoids with the same unit space.
Note that every AF groupoids is locally compact, Hausdorff, second countable, étale, totally disconnected, and amenable. If $G$ an AF groupoid is the increasing union of the open elementary subgroupoids $G_{n}$. Then $G$ is the inductive limit $G=\displaystyle\lim_{\longrightarrow}G_{n}$ [48, Section III.1].
Remark 2.4.
Given an elementary groupoid $G\cong R(\sigma)$ with $\sigma:X\rightarrow Y$, one we apply the isomorphism to find a local homeomorphism $\widetilde{\sigma}:G^{(0)}\rightarrow Y$ such that $G=R(\widetilde{\sigma})$. So, we can suppose that $G=R(\sigma)$ without loss of generality.
Example 2.5.
Let $X$ be a locally compact, Hausdorff, second countable, totally disconnected space, and let $\sigma:X\rightarrow X$ be a surjective local homeomorphism. Define the set
$c^{-1}(0)=\{(x,y)\in X^{2}:\sigma^{n}(x)=\sigma^{n}(y)\text{ forsome }n\in\mathbb{N}\},$
and equip this set with the operations $(x,y)(y,z)=(x,z),(x,y)^{-1}=(y,x)$ and with the range a source maps $r,s:c^{-1}(0)\rightarrow X$ given by $r(x,y)=x$ and $s(x,y)=y$. We also equip the set with the subspace topology from $X\times X$. Then $c^{-1}(0)$ is an AF groupoids. In fact, it is the increasing union of the open elementary subgroupoids $R(\sigma^{n})$.
We use the notation $c^{-1}(0)$ because of the following. If $\mathcal{G}$ is the Deaconu-Renault groupoid and $c:\mathcal{G}\rightarrow\mathbb{Z}$ is the canonical cocycle, then $c^{-1}(0)$ is isomorphic to the subgroupoid $\{(x,0,y)\in\mathcal{G}\}$ of $\mathcal{G}$. The isomorphism is defined by $(x,y)\mapsto(x,0,y)$.
3. Skew-products and crossed products
As explained in the introduction, we use an isomorphism between the C*-algebras a skew-product groupoid and the corresponding crossed product to understand when the C*-algebra of a Deaconu-Renault groupoid $\mathcal{G}$ is AFE. This isomorphism allows us to apply the results of Brown’s Theorem (Theorem 1.1). In this section, we use Renault’s book [48] as a reference.
Definition 3.1.
[48, Definition I.1.6] Let $G$ be a groupoid and $c:G\rightarrow\mathbb{Z}$ a homomorphism. Define the sets
$G(c)=\{(g,a):g\in G,a\in\mathbb{Z}\}\hskip 10.0pt\text{and}\hskip 10.0ptG(c)^{(0)}=G^{(0)}\times\mathbb{Z}.$
Define the maps $r,s:G(c)\rightarrow G^{(0)}\times\mathbb{Z}$ by $r(g,a)=(r(g),a)$ and $s(g,a)=(s(g),a+c(g))$. Consider the set
(3.1) $G(c)^{(2)}=\{((g,a),(h,a+c(g))):(g,h)\in G^{(2)},a\in\mathbb{Z}\},$
and define the product $G(c)^{(2)}\rightarrow G(c)$ and the inverse $G(c)\rightarrow G(c)$ by
(3.2) $(g,a)(h,a+c(g))=(gh,a)\hskip 15.0pt\text{and}\hskip 15.0pt(g,a)^{-1}=(g^{-1},a+c(g)),$
respectively. Then $G(c)$ becomes a groupoid, called a skew-product.
If $c$ is a continuous cocycle (i.e., if $c:G\rightarrow\mathbb{Z}$ is a continuous homomorphism) and we equip $G(c)$ with the induced topology from $G\times\mathbb{Z}$, then $G(c)$ inherits the usual topogical properties of $G$. Recall that we assume that $G$ is locally compact, Hausdorff, second countable and étale.
Lemma 3.2.
Equip $G(c)$ with the topology induced from $G\times\mathbb{Z}$. Then $G(c)$ is also a locally compact, Hausdorff, second countable, étale groupoid. Moreover, if $G$ is totally disconnected, then $G(c)$ is totally disconnected, too. If $G$ is measurewise amenable, then $G(c)$ is also measurewise amenable.
Note that can identify the unity space $G(c)^{(0)}$ with $\mathcal{X}\times\mathbb{Z}$. For $f,f_{1},f_{2}\in C_{c}(G(c))$ and $(g,a)\in G\times\mathbb{Z}$, the operations on $C_{c}(G(c))$ are given by
$(f_{1}\cdot f_{2})(g,a)=\sum_{g_{1}g_{2}=g}f_{1}(g_{1},a)f_{2}(g_{2},a+c(g_{1})),\text{ and }f^{*}(g,a)=\overline{f(g^{-1},a+c(g))}.$
Assume that $G$ is a locally compact, Hausdorff, second countable, étale, measurewise amenable groupoid, and suppose that $c:G\rightarrow\mathbb{Z}$ is a continuous cocycle. We define a dynamical system $\alpha^{c}$ induced by $c$. For every $z\in\mathbb{T}$, $f\in C_{c}(G)$ and $g\in G$, we set $\alpha_{z}^{c}(f)(g)=z^{c(g)}f(g)$. By Proposition II.5.1 of Renault’s book [48], there exists a unique dynamical system defined by the equation above. We make an abuse of notation by denoting this dynamical system $\alpha^{c}:\mathbb{T}\rightarrow\mathrm{Aut}(C^{*}(G))$. The following proposition by Renault [48, Proposition II.5.7] gives an isomorphism between the C*-algebra of skew-product groupoids and the corresponding crossed products. For an introduction to crossed products, see Chapters 1 and 2 of Dana William’s book [59].
Proposition 3.3.
There exists a unique isomorphism $\rho\rtimes u:C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rightarrow C^{*}(G(c))$ such that
[…]
for $f\in C_{c}(\mathbb{T},C_{c}(G))$ and $(g,a)\in G(c)$. Moreover, the inverse of $\rho\rtimes u$ is given by
$(\rho\rtimes u)^{-1}(F)(z)(g)$ $=\frac{1}{2\pi}\sum_{a\in\mathbb{Z}}z^{c(g)+a}F(g,a),$
for $F\in C_{c}(G(c))$, $z\in\mathbb{T}$, $g\in G$.
Now we find a formula for the action $\beta$ on $C^{*}(G(c))$ induced by the dual action $\widehat{\alpha}^{c}$. We will apply this formula in Chapter 7 for Deaconu-Renault groupoids in order to understand in more detail the condition $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$, given by Brown’s theorem.
Proposition 3.4.
Let $\rho\rtimes u:C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rightarrow C^{*}(G(c))$ be the isomorphism of Proposition 3.3. Then there exists a unique dynamical system $\beta:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G(c)))$ such that
$\beta(F)(g,a)=F(g,a+1)\hskip 15.0pt\text{for $F\in C_{c}(G(c))$, $(g,a)\in G(c)$,}$
and $\beta=(\rho\rtimes u)\circ\widehat{\alpha}^{c}\circ(\rho\rtimes u)^{-1}$ for all $n\in\mathbb{Z}$. Moreover, $C^{*}(G(c))\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic.
Proof.
Let $\widehat{\alpha}^{c}:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T})$ be the dual action, which is given by $\widehat{\alpha}^{c}_{n}(f)(z)=z^{-n}f(z)$, for $f\in C(\mathbb{T},C^{*}(G))$, $z\in\mathbb{T}$, $n\in\mathbb{Z}$. For each $n\in\mathbb{Z}$, let $\beta_{n}=(\rho\rtimes u)^{-1}\circ\widehat{\alpha}^{c}_{n}\circ\rho\rtimes u$. Since $\rho\rtimes u$ is an isomorphism and $\widehat{\alpha}^{c}$ is an automorphism, we have that $\beta:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G(c)))$ is a dynamical system.
Fix $n\in\mathbb{Z}$. By continuity of $\beta_{n}$, finding the formula of this automorphism on $C_{c}(G(c))$ is sufficient to determine $\beta_{n}$ uniquely. Let $F\in C_{c}(G(c))$ and $(g,a)\in G(c)$. Then, Proposition 3.3,
$\beta_{n}(F)(g,a)$ $=(\rho\rtimes u)\circ\widehat{\alpha^{c}}_{n}\circ(\rho\rtimes u)^{-1}(F)(g,a)$
[…]
$=\frac{1}{2\pi}\int_{\mathbb{T}}z^{-c(g)-a-n}\sum_{b\in\mathbb{Z}}z^{c(g)+b}F(g,b)dz$
[…]
$=F(g,a+n).$
By definition, we have that $\rho\rtimes u\circ\widehat{\alpha}^{c}=\beta\circ\rho\rtimes u$. Then $\rho\rtimes u$ is an equivariant isomorphism from $(C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T},\widehat{\alpha}^{c},\mathbb{Z})$ onto $(C^{*}(G(c)),\beta,\mathbb{Z})$. By [59, Lemma 2.65], $C^{*}(G(c))\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic. ∎
4. Homology Groups and K-theory
Our main result, Theorem 1 is an application of Brown’s theorem (Theorem 1.1), which characterises the AF embeddability of a C*-algebra $B\rtimes_{\beta}\mathbb{Z}$ in terms of the K-theory of $B$. Note that condition (iv) is very abstract so, in order to provide a more concrete condition, we need to understand the K-theory of the groupoid C*-algebra $C^{*}(\mathcal{G}(c))$.
The definition of K-theory is very abstract and not very manageable. So, we are going to apply an isomorphism between the K-theory $K_{0}(B)$ and its corresponding homology group $H_{0}(\mathcal{G}(c))$, which allows us to apply several helpful results from the theory of homology groups that will allow us to understand what condition (iv) in Brown’s theorem tells us about the groupoid $\mathcal{G}$.
There is already in the literature an isomorphism between the K-theory and the corresponding homology group for AF algebras, described in [20, Corollary 5.2] by Farsi, Kumjian, Pask and Sims. This isomorphism is not sufficient to study condition (iv) of Brown’s theorem in terms of homology groups because it is not clear if this isomorphism preserves positive elements. So, we are going to extend [20, Corollary 5.2] by showing in Theorem 2 that the isomorphism preserves positive elements, and we give a more explicit formula for the isomorphism. Our idea was to consider the building blocks $R(\sigma)$ of AF groupoids and show that the C*-algebras $C^{*}(R(\sigma))$ are approximated by C*-algebras of the form $C_{0}(V,M_{n})$, whose K-groups are much easier to understand. Theorem 2 also generalises [39, Theorem 4.10] which considers only AF groupoids with compact unit spaces.
4.1. Homology in algebra
First we present the definition of homology groups from abstract algebra. A possible reference is [16, Section 17.1].
Definition 4.1.
Let $\mathcal{C}$ be a sequence of abelian group homomorphisms
[…]
such that $d_{n}\circ d_{n+1}=0$ for all $n\geq 0$. The sequence is called a chain complex. If $\mathcal{C}$ is a chain complex, its $n^{\text{th}}$ homology group is the quotient group $\ker d_{n}/\mathrm{im\hphantom{.}}d_{n+1}$, and is denoted by $H_{n}(\mathcal{C})$.
Definition 4.2.
Let $\mathcal{A}=\{A_{n}\}_{n\geq 0}$ and =$\mathcal{B}=\{B_{n}\}_{n\geq 0}$ be chain complexes. A homomorphism of complexes $\alpha:\mathcal{A}\rightarrow\mathcal{B}$ is a set of homomorphisms $\alpha_{n}:A_{n}\rightarrow B_{n}$ such that for every $n$ the following diagram commutes:
[…]
Proposition 4.3.
A homomorphism $\alpha:\mathcal{A}\rightarrow\mathcal{B}$ of chain complexes induces a group homomorphism $H_{n}(\alpha):H_{n}(\mathcal{A})\rightarrow H_{n}(\mathcal{B})$, for $n\geq 0$, given by
(4.1) $H_{n}(\alpha)([a])=[\alpha_{n}(a)],\text{\hskip 10.0ptfor $a\in A_{n}$ with $d_{n}(a)=0$.}$
4.2. Homology for groupoids
We define homology groups for Haudorff, second countable, ample groupoids, as in [20, 39]. We fix $X,Y$ to be locally compact, Hausdorff, second countable, totally disconnected spaces.
Here $C_{c}(X,\mathbb{Z})$ denotes the set of continuous and compactly supported functions on $X$ assuming integer values. Given $f\in C_{c}(X,\mathbb{Z})$ and a local homeomorphism $\sigma:X\rightarrow Y$, define for every $y\in Y$,
$\sigma_{\ast}(f)(y)=\sum_{x:\sigma(x)=y}f(x).$
This equation defines a function $\sigma_{\ast}(f)\in C_{c}(X,\mathbb{Z})$. For $n\geq 1$, we define the space of composable $n$-tuples by
$G^{(n)}=\{(g_{1},\dots,g_{n})\in G^{n}:s(g_{i})=r(g_{i+1})\text{for }1\leq i<n\}.$
We let $d_{0}^{(1)}=s$, $d_{1}^{(1)}=r$, and for $n\geq 2$ we define
$d_{i}^{(n)}(g_{1},\ldots,g_{n})=\begin{cases}(g_{2},\ldots,g_{n})&i=0,\\ (g_{1},\ldots,g_{i-1},g_{i}g_{i+1},g_{i+2},\ldots,g_{n})&1\leq i\leq n-1\\ (g_{1},\ldots,g_{n-1})&i=n.\end{cases}$
Since $G$ is étale, the maps $d_{i}^{(n)}$ are local homeomorphisms. We use these maps to construct a chain complex. For $n\geq 1$, we define $\partial_{n}:C_{c}(G^{(n)},\mathbb{Z})\rightarrow C_{c}(G^{(n-1)},\mathbb{Z})$ by
$\partial_{n}=\sum_{i=0}^{n}(-1)^{i}(d_{i}^{(n)})_{\ast}.$
We set $\partial_{0}:C_{c}(G^{(0)},\mathbb{Z})\rightarrow 0$. The sequence
[…]
(also denoted $(C_{c}(G^{(\ast)},\mathbb{Z}),\partial_{\ast})$) is a chain complex. This is true because, for $n\geq 1$,
$\partial_{n}\partial_{n+1}=\sum_{i=0}^{n}\sum_{j=0}^{n+1}(-1)^{i+j}(d_{i}^{(n)}d_{j}^{(n+1)})_{\ast}=0,$
which can be proved by showing that $d_{i}^{(n)}d_{j}^{(n+1)}=d_{j-1}^{(n)}d_{i}^{(n+1)}$ for all $10\leq i<j\leq n+$ 1, dividing the prove into different cases.
Definition 4.4.
For $n\geq 0$, let $H_{n}(G)$ be the $n^{\text{th}}$ homology group of the chain complex $(C_{c}(G^{(\ast)},\mathbb{Z}),\partial_{\ast})$. In addition, we define the set of positive elements
$H_{0}(G)^{+}=\{[f]\in H_{0}(G):f(x)\geq 0\text{ for all }x\in G^{(0)}\}.$
The operations on $H_{0}(G)$ are given by $[f_{1}]+[f_{2}]=[f_{1}+f_{2}]$ and $-[f]=[-f]$.
The group $H_{0}(G)$ is not necessarily an ordered group because the intesection $H_{0}(G)^{+}\cap(-H_{0}(G)^{+})$ may be different from $\{0\}$. See [39, Remark 3.2].
Given a homomorphism $\rho:G\rightarrow H$ between étale groupoids, we define $\rho^{(0)}:G^{(0)}\rightarrow H^{(0)}$ by $\rho^{(0)}=\rho|_{G^{(0)}}$. For $n\geq 1$, we define the map $\rho^{(n)}:G^{(n)}\rightarrow H^{(n)}$ by
$\rho^{(n)}(g_{1},\ldots,g_{n})=(\rho(g_{1}),\ldots,\rho(g_{n})),\text{\hskip 10.0ptfor $(g_{1},\ldots,g_{n})\in G^{(n)}$.}$
We prove the lemma below, stated in [39, Subsection 3.2] for étale groupoids.
Lemma 4.5.
Let $G,H$ be locally compact étale groupoids and let $\rho:G\rightarrow H$ be a homomorphism. Then the following are equivalent:
  1. (i)
    $\rho^{(0)}$ is a local homeomorphism,
  2. (ii)
    $\rho$ is a local homeomorphism,
  3. (iii)
    $\rho^{(n)}$ is a local homeomorphism for all $n\in\mathbb{N}$.
Proof.
The implication $\mathrm{(iii)}\Rightarrow\mathrm{(ii)}\Rightarrow\mathrm{(i)}$ is straightforward. To prove $\mathrm{(i)}\Rightarrow\mathrm{(ii)}$, first we show that $\rho$ is an open map. Let $\mathcal{U}$ be an open bisection such that the restriction $\rho|_{r(\mathcal{U})}^{(0)}:r(\mathcal{U})\rightarrow\rho(r(\mathcal{U}))$ is a homeomorphism. Let $V\subset\overline{V}\subset\mathcal{U}$ be an open set, and let $g_{n}$ be a sequence in $\mathcal{U}\setminus V$ converging to an $g\in\mathcal{U}\setminus V$. Since $\mathcal{U}$ is étale, we have $r(g)\notin r(V)$. Since $\rho$ is injective in $\mathcal{U}$, we have $r\circ\rho(g)\notin\rho\circ r(\mathcal{V})$, which implies that $\rho(g)\notin\rho(V)$. Then $\rho(V)$ is open. It is straightforward that $\rho$ is injective on open bisections. Let $\mathcal{U}_{1}\subset G$, $V_{1}\subset H$ be open bisections such that $\rho(\mathcal{U}_{1})\subset V_{1}$. One can show that, locally, $\rho$ can be identified with the map $r|_{V_{1}}^{-1}\circ\rho\circ r|_{\mathcal{U}_{1}}$.
To show $\mathrm{(ii)}\Rightarrow\mathrm{(iii)}$ for $n\geq 2$, choose open bisections $\mathcal{U}_{1},\dots,\mathcal{U}_{n}$ such that $r(\mathcal{U}_{i})=s(\mathcal{U}_{i+1})$ for $i=1,\dots,n-1$, and show that $\rho^{n}(\mathcal{U}\cap G^{n})=\rho^{n}(\mathcal{U})\cap H^{(n)}$, where
$\rho^{n}=\underbrace{\rho\times\dots\times\rho}_{\text{$n$ times}}.$
Since $\rho^{(n)}$ is continuous and injective on $\mathcal{U}\cap G^{(n)}$, then $\rho$ is a local homeomorphism. ∎
Proposition 4.6.
Let $\rho:\rightarrow H$ be a homomorphism that is also a local homeomorphism. Then, for each $n\geq 0$, $\rho$ induces a homomorphism $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ defined by
$H_{n}(\rho)([f])=[\rho_{\ast}^{(n)}(f)]\hskip 7.0pt\text{for }f\in C_{c}(G^{(n)},\mathbb{Z}).$
Remark 4.7.
In this paper, for a function $\tau:C_{c}(G^{(0)},\mathbb{Z})\rightarrow C_{c}(H^{(0)},\mathbb{Z})$ we use the notation $[\tau]$ to denote the function $[\tau]:H_{0}(G)\rightarrow H_{0}(H)$ given by $[\tau]([f])=[\tau(f)]$. Moreover, for a homomorphism $\rho:G\rightarrow H$ that is also a local homeomorphism, we have $H_{0}(\rho)=[\rho_{\ast}^{(0)}]$.
An important tool in the study of homology groups is the notion of homological similarity, that can be applied to show that two groupoids have isomorphic homology groups.
Definition 4.8.
  1. (i)
    Two homomorphisms $\rho,\sigma$ from $G$ to $H$ are said to be similar if there exists a continuous map $\theta:G^{(0)}\rightarrow H$ such that
    (4.2) $\theta(r(g))\rho(g)=\sigma(g)\theta(s(g)),$
    for all $g\in G$.
  2. (ii)
    $G$ and $H$ are said to be homologically similar if there exist homomorphisms $\rho:G\rightarrow H$ and $\sigma:H\rightarrow G$ that are local homeomorphisms and such that $\sigma\circ\rho$ is similar to $\mathrm{id}_{G}$ and $\rho\circ\sigma$ is similar to $\mathrm{id}_{H}$.
Proposition 4.9.
[39, Proposition 3.5] Let $n\geq 1$. If $\rho,\sigma:G\rightarrow H$ are similar homomorphisms and local homeomorphisms, then $H_{n}(\rho)=H_{n}(\sigma)$.
Corollary 4.10.
[39, Proposition 3.5] If $G$ and $H$ are homologically similar, then they have isomorphic homology groups. In particular, if $n\geq 1$, $\rho:G\rightarrow H$ and $\sigma:H\rightarrow G$ are local homeomorphisms and homomorphisms satisfying (4.2), then $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ is an isomorphism with inverse $H_{n}(\sigma)$. Moreover, both $H_{0}(\rho)$ and $H_{0}(\sigma)$ preserve positive elements.
4.3. Homology for AF groupoids
Like K-theory, homology groups are continuous with respect to inductive limits. We start this subsection by looking at the building blocks of AF groupoids, i.e., groupoids of the form $R(\sigma)$ and finding a property that helps us find the equivalence in $H_{0}(R(\sigma))$.
Lemma 4.11.
Let $\sigma:X\rightarrow Y$ be a surjective local homeomorphism. Then two functions $f_{1},f_{2}\in C_{c}(X,\mathbb{N})$ are equivalent in $H_{0}(R(\sigma))$ if, and only if,
(4.3) $\sum_{u:\sigma(u)=\sigma(x)}f_{1}(u)=\sum_{u:\sigma(u)=\sigma(x)}f_{2}(u)\hskip 15.0pt\text{for all $x\in X$.}$
Proof.
First, let us define a map $T:C_{c}(X,\mathbb{N})\rightarrow C_{c}(X,\mathbb{Z})$ which is going to be used in the proof. Fix a continuous section $\varphi:Y\rightarrow X$ of $\sigma$. Let $T:C_{c}(X,\mathbb{N})\rightarrow C_{c}(X,\mathbb{N})$ be given by
(4.4) $T(f)(x)=1_{X_{\varphi}}(x)\sum_{u:\sigma(u)=\sigma(x)}f(u),$
where $X_{\varphi}=\varphi(Y)$. To show that $T$ is well-defined, Fist we consider functions of the form $1_{V}$, where $V\subset X$ be a compact open subset on which $\sigma$ is injective. We claim that $T(1_{V})=1_{\varphi(\sigma(V))}$. Let $x\in X$. Suppose that $T(1_{V})(x)\neq 0$. Equation (4.4) implies that $x\in X_{\varphi}$ and that there exists an $u\in X_{\varphi}$ such that $u\in V$ and $\sigma(u)=\sigma(x)$. Since $\sigma$ is injective on $V$, such $u$ is unique, which implies that $x=\varphi(\sigma(u))$. Thus $1_{\varphi(\sigma(V))}(x)=1_{V}(x)$.
Now suppose that $T(1_{V})(x)=0$. Then $x\notin X_{\varphi}$ or $\sum_{u:\sigma(u)=\sigma(x)}1_{V}(u)=0$. If $x\notin X_{\varphi}$, then $x\notin\varphi(\sigma(V))$ because $\varphi(\sigma(V))\subset X_{\varphi}$. Then $01_{\varphi(\sigma(V))}(x)=$ 0. If $\sum_{u:\sigma(u)=\sigma(x)}1_{V}(u)=0$, then $\varphi(\sigma(x))\notin V$, which implies that $T(1_{V})(x)=1_{\varphi(\sigma(V))}(x)=0$. Therefore $T(1_{V})=1_{\varphi(\sigma(V))}$.
Note that every function $f\in C_{c}(X,\mathbb{N})$ is spanned by function of the form $1_{V}$, as considered above. Since $T$ is a semigroup homomorphism, then $T(f)\in C_{c}(X,\mathbb{N})$. Therefore, $T$ is well-defined.
Let $f_{1},f_{2}\in C_{c}(X,\mathbb{N})$ be equivalent in $H_{0}(R(\sigma))$. Then there exists $F\in C_{c}(R(\sigma),\mathbb{Z})$ such that, for all $u\in X$,
$f_{1}(u)$ $=f_{2}(u)+\partial_{1}F(u)$
$=f_{2}(u)+s_{\ast}(F)(u)-r_{\ast}(F)(u)$
$=f_{2}(u)+\sum_{v:\sigma(v)=\sigma(u)}[F(v,u)-F(u,v)].$
Fix $x\in X$. Then
$\sum_{u:\sigma(u)=\sigma(x)}f_{1}(u)$ $=\sum_{u:\sigma(u)=\sigma(x)}f_{2}(u)+\sum_{\begin{subarray}{c}u,v\\ \sigma(u)=\sigma(x)\\ \sigma(v)=\sigma(x)\end{subarray}}(F(v,u)-F(u,v))$
$=\sum_{u:\sigma(u)=\sigma(x)}f_{2}(u)+\sum_{\begin{subarray}{c}u,v\\ \sigma(u)=\sigma(x)\\ \sigma(v)=\sigma(x)\end{subarray}}F(v,u)-\sum_{\begin{subarray}{c}u,v\\ \sigma(u)=\sigma(x)\\ \sigma(v)=\sigma(x)\end{subarray}}(F(u,v).$
Note that the second and the third sums of the right-hand side are equal. Then (4.3) holds.
Conversely, suppose that (4.3) holds. Define the set
$K=\{x\in X:x\in\mathrm{supp\hphantom{.}}f_{1}\cup\mathrm{supp\hphantom{.}}f_{2}\text{\hskip 7.0pt and \hskip 7.0pt}T(f_{1})(\varphi(\sigma(x)))\neq 0\}.$
This set is compact because it is the intesection of the compact set $\mathrm{supp\hphantom{.}}f_{1}\cup\mathrm{supp\hphantom{.}}f_{2}$ and the closed set $(T(f_{1})\circ\varphi\circ\sigma)^{-1}(\mathbb{Z}\setminus\{0\})$.
Let $F\in C_{c}(R(\sigma),\mathbb{Z})$ be defined by
$F(x,y)$ =$=\begin{cases}\displaystyle\frac{f_{1}(y)f_{2}(x)}{T(f_{1})(\varphi(\sigma(x)))}&\text{if }(x,y)\in K\times\mathrm{supp\hphantom{.}}f_{2},\\ 0&\text{otherwise.}\end{cases}$
Then, for $x\in K$, we have
$\sum_{y:\sigma(y)=\sigma(x)}F(x,y)$ $=\frac{f_{2}(x)}{T(f_{1})(\varphi(\sigma(x)))}\sum_{y:\sigma(y)=\sigma(x)}f_{1}(y)$
$=\frac{f_{2}(x)}{T(f_{1})(\varphi(\sigma(x)))}T(f_{1})(\varphi(\sigma(x)))$
(4.5) $=f_{2}(x).$
If $x\in X\setminus K$, we have
(4.6) $\sum_{y:\sigma(y)=\sigma(x)}F(x,y)=0=f_{2}(x).$
By similar arguments, we have
(4.7) $\sum_{y:\sigma(y)=\sigma(x)}{F(y,x)}=f_{1}(x)\hskip 20.0pt\text{for $x\in X$.}$
The equations (4.5), (4.6) and (4.7) imply that, for all $x\in X$, $f_{1}(x)=f_{2}(x)+\partial_{1}F(x).$ Therefore $f_{1}\sim f_{2}$. ∎
The proof of the following lemma is straightforward, and this lemma will be applied in the proof of Theorem 1.
Lemma 4.12.
Let $G=\displaystyle\lim_{\rightarrow}G_{n}$. For each $n$, let $\partial_{1}^{(n)}:C_{c}(G_{n},\mathbb{Z})\rightarrow C_{c}(G^{(0)},\mathbb{Z})$ be defined by $\partial_{1}^{(n)}=(s|_{G_{n}})_{\ast}-(r|_{G_{n}})_{\ast}$. Analogously, let $\partial_{1}:C_{c}(G,\mathbb{Z})\rightarrow C_{c}(G^{(0)},\mathbb{Z})$ be given by $\partial_{1}=s_{\ast}-r_{\ast}$. Let $F\in C_{c}(G_{n},\mathbb{Z})$. Then for all $k\geq 0$, we have $\partial_{1}F=\partial_{1}^{(n)}F=\partial_{1}^{(n+k)}F.$
Now we show that $H_{0}$ is continuous with respect to inductive limits.
Lemma 4.13.
Let $G=\displaystyle\lim_{\rightarrow}G_{n}$. Then $H_{0}(G)=\displaystyle\lim_{\rightarrow}H_{0}(G_{n})$ as an inductive limit of groups. Moreover, if each $H_{0}(G_{n})$ is an ordered group, then $H_{0}(G)$ is an ordered group.
Proof.
[…]
(4.8) $\lambda_{n}=\lambda_{n+1}\circ i_{n}.$
If the groups $H_{0}(G_{n})$. are ordered, suppose that $B$ is ordered and that the maps $\lambda_{n}$ preserves positive elements. Let $\lambda:H_{0}(G)\rightarrow B$ be defined by $\lambda([f])=\lambda_{0}(f)$. The equations (4.8) imply that $\lambda$ is well-defined. Indeed, let $f_{1},f_{2}\in C_{c}(G^{(0)},\mathbb{Z})$ be such that $f_{1},f_{2}$ are equivalent in $H_{0}(G_{n})$. Then there exists $F\in C_{c}(G_{n},\mathbb{Z})$ such that $f_{1}=f_{2}+\partial_{1}F$. Then there is an $n$ such that $F\in C_{c}(G_{n},\mathbb{Z})$. Then, by Lemma 4.12, $f_{1}=f_{2}+\partial_{1}^{(n)}F$. Applying (4.8) multiply times, we have
$\lambda_{0}(f_{1})=\lambda_{1}\circ i_{0}(f_{1})=\lambda_{1}([f_{1}]_{1})=\lambda_{2}([f_{1}]_{2})=\dots=\lambda_{n}([f_{1}]_{n})=\lambda_{n}([f_{2}]_{n})=\dots=\lambda_{0}(f_{2}).$
Thus $\lambda$ is well-defined. It is straightforward to show that $\lambda$ is the only homomorphism such that $\lambda\circ\alpha_{n}=\lambda_{n}$. Therefore $H_{0}(G)=\displaystyle\lim_{\rightarrow}H_{0}(G_{n})$. If each $H_{0}(G_{n})$ is an ordered group, then $H_{0}(G)$ is an ordered group. ∎
Now we show that, for AF groupoids, the zeroth homology group is an ordered group.
Proposition 4.14.
Let $G$ be an AF groupoid. Then $H_{0}(G)$ is an ordered group.
Proof.
We begin by showing that, for a surjective local homeomorphism $\sigma:X\rightarrow Y$ of locally compact, Hausdorff, second countable, totally disconnected spaces, the group $H_{0}(R(\sigma))$ is an ordered group. Indeed, it is straightforward from the definition that $H_{0}(R(\sigma))=H_{0}(R(\sigma))^{+}-H_{0}(R(\sigma))^{+}$. We now prove the intersection
$H_{0}(R(\sigma))^{+}\cap(-H_{0}(R(\sigma))^{+})=\{0\}.$
Choose a function $f\in C_{c}(X,\mathbb{Z})$ with $[f]\in H_{0}(R(\sigma))^{+}\cap(-H_{0}(R(\sigma))^{+})$. Then there are two functions $f_{1},f_{2}\in C_{c}(X,\mathbb{Z})$ assuming non-negative values, such that $[f]=[f_{1}]=[-f_{2}].$ Choose $x\in X$. By Lemma 4.11,
$0\leq f_{1}(x)\leq\sum_{u:\sigma(u)=\sigma(x)}f_{1}(u)=-\sum_{u:\sigma(u)=\sigma(x)}f_{2}(u)\leq-f_{2}(x)\leq 0$
Then $f_{1}(x)=f_{2}(x)=0$. Since $x$ is arbitrary, we have $f_{1}=f_{2}=0$. This implies that $H_{0}(R(\sigma))^{+}\cap(-H_{0}(R(\sigma))^{+})=\{0\}$. Therefore, $H_{0}(R(\sigma))$ is an ordered group.
Now let $G$ be an AF groupoid. Then there exists a sequence of surjective local homeomorphisms $\sigma_{n}:G^{(0)}\rightarrow Y_{n}$, where $G^{(0)}$ and $Y_{n}$ are locally compact, Hausdorff, second countable and totally disconnected, such that $G=\displaystyle\lim_{\rightarrow}R(\sigma_{n})$. Since each $H_{0}(R(\sigma_{n}))$ is an ordered group, then Lemma 4.13 implies that $H_{0}(G)$ is an ordered group. ∎
5. K-theory for AF Groupoids
Here we extend the isomorphism [20, Corollary 5.2] between the K-theory and the homology group for AF groupoids, showing that the map preserves positive elements, and giving a formula for the isomorphism.
Although it is often difficult to find the K-theory of an arbitrary C*-algebra, finding the K-theory of $M_{n}$ is a easy. For this case, we get the isomorphism $K_{0}(M_{n})\rightarrow\mathbb{Z}$ given by $[p]_{0}\mapsto\mathrm{tr}\hphantom{.}p$, where $\mathrm{tr}\hphantom{.}p$ is the matrix trace of $p\in\mathcal{P}_{\infty}(M_{n})$. Similarly, for a locally compact, Hausdorff, second countable, totally disconnected space $X$, the K-theory of $C_{0}(X,M_{n})$ is also described in terms of a trace operator.
To find the K-theory of an AF groupoid, we first consider the building blocks of the groupoid. In other words, we consider an elementary groupoid of the form $R(\sigma)$. We show that $R(\sigma)$ is locally approximated by groupoids of the form $C_{0}(V,M_{n})$. Then we apply approximation to find a formula for the $K_{0}$-group of $C^{*}(R(\sigma))$ and an isomorphism from this group to the corresponding homology group.
Throughout this section, we write $[f_{jk}]$ instead of $[f_{jk}]_{j,k=1,\dots,n}$ to simplify our notation. We also assume that the topological spaces $X,Y,X_{n}$ are locally compact, Hausdorff, second countable, and totally disconnected. The map $\sigma:X\rightarrow Y$ is a surjective local homeomorphism.
Remark 5.1.
Unless otherwise stated, we assume that $X$ has a compact open subset $X_{0}$ such that $\sigma^{-1}(\sigma(X_{0}))=X_{0}$ and $\sigma$ is injective on $X\setminus X_{0}$. Later on, we prove that an arbitrary elementary groupoid is the inductive limit of groupoids of the form $R(\sigma)$ where this condition is satisfied.
Lemma 5.2.
Let $\varphi:Y\rightarrow X$ be a continuous section of $\sigma$, and let $y\in Y$. Then $y$ has a clopen neighbourhood $W$, there is a positive integer $n$, and there are disjoint subsets $\mathcal{U}_{1},\dots,\mathcal{U}_{n}\subset X$ such that
  1. (i)
    $\sigma^{-1}(W)=\bigcup_{i=1}^{n}\mathcal{U}_{i}$,
  2. (ii)
    $\sigma|_{\mathcal{U}_{i}}:\mathcal{U}_{i}\rightarrow W$ is a homeomorphism for all $i$, and
  3. (iii)
    $\sigma^{-1}(W)\cap X_{\varphi}=\mathcal{U}_{1}$.
If $y\in\sigma(X_{0})$, then $W$ and the sets $\mathcal{U}_{i}$ are compact. If $y\in Y\setminus\sigma(X_{0})$, we let $n=1$, $W=Y\setminus\sigma(X_{0})$ and $\mathcal{U}_{1}=X\setminus X_{0}$.
Proof.
The proof of the case $y\in Y\setminus\sigma(X_{0})$ follows directly from the fact that $\sigma$ is injective on $X\setminus X_{0}$. Suppose $y\in\sigma(X_{0})$. Since $\sigma$ is a local homeomorphism, $X_{0}$ is compact and $\sigma^{-1}(\sigma(X_{0}))=X_{0}$, then $y$ has a compact open neighbourhood $W$ and there are finite disjoint compact open subsets $\mathcal{O}_{1},\dots,\mathcal{O}_{n}$ with
$\sigma^{-1}(W)=\bigcup_{i=1}^{n}\mathcal{O}_{i},$
and such that $\sigma|_{\mathcal{O}_{i}}:\mathcal{O}_{i}\rightarrow W$ is a homeormorphism for all $i$.
Define the set $\mathcal{U}_{1}=\sigma^{-1}(W)\cap X_{\varphi}$. Then $\sigma$ is injective on $\mathcal{U}_{1}$. Moreover, $\sigma(\mathcal{U}_{1})=W$ because $\sigma(\sigma^{-1}(X)\cap X_{\varphi}))=W\cap Y=W$. Therefore, $\sigma|_{\mathcal{U}_{1}}:\mathcal{U}_{1}\rightarrow W$ is a homeomorphism. By rearranging the sets $\mathcal{O}_{1},\dots,\mathcal{O}_{n}$ if necessary, let $m\leq n$ be such that $\mathcal{U}_{1}\cap\mathcal{O}_{i}=\emptyset\Leftrightarrow j>m$. Then we have the disjoint union
$\mathcal{U}_{1}=\bigcup_{i=1}^{m}\mathcal{U}_{i}\cap\mathcal{O}_{i}.$
For all $i,j=1,\dots,m$, let $\Omega_{ij}=\sigma^{-1}(\sigma(\mathcal{U}_{1}\cap\mathcal{O}_{j}))\cap\mathcal{O}_{i}$. Then these sets are disjoint. For all $i$, we have
(5.1) $\bigcup_{i=1}^{m}\Omega_{ii}=\bigcup_{i=1}^{m}\sigma^{-1}(\sigma(\mathcal{U}_{1}\cap\mathcal{O}_{i}))\cap\mathcal{O}_{i}=\bigcup_{i=1}^{m}\mathcal{U}_{1}\cap\mathcal{O}_{i}=\mathcal{U}_{1},$
because $\sigma$ is injective on each $\mathcal{O}_{i}$. Similarly,
(5.2) $\bigcup_{j=1}^{m}\Omega_{ij}=\sigma^{-1}\left(\sigma\left(\bigcup_{j=1}^{m}\mathcal{U}_{1}\cap\mathcal{O}_{j}\right)\right)\cap\mathcal{O}_{i}=\sigma^{-1}(\sigma(\mathcal{U}_{1}))\cap\mathcal{O}_{i}=\sigma^{-1}(W)\cap\mathcal{O}_{i}=\mathcal{O}_{i}.$
Also,
(5.3) $\sigma(\Omega_{ij})=\sigma\left(\sigma^{-1}\left(\sigma(\mathcal{U}_{1}\cap\mathcal{O}_{j})\right)\cap\mathcal{O}_{i}\right)=\sigma(\mathcal{U}_{1}\cap\mathcal{O}_{j})=\sigma(\Omega_{1j}).$
Since $\sigma$ is injective on $X_{\varphi}$, (5.1) implies that $\Omega_{ij}\cap X_{\varphi}=\emptyset$ for $i\geq j$.
For $i=2,\dots,m$, let $\mathcal{U}_{i}=\bigcup_{j=1}^{m}\Omega_{ij}$. For $i=m+1,\dots,n$, let $\mathcal{U}_{i}=\mathcal{O}_{i}$. Then equations (5.1), (5.2), (5.3) imply items (i)–(iii). ∎
Lemma 5.3.
Let $p,q\in C_{0}(X,M_{n})$ be two projections. Then $p,q\in C_{c}(X,\mathbb{N})$ and $\mathrm{tr}\hphantom{.}p,\mathrm{tr}\hphantom{.}q\in C_{c}(X,\mathbb{N})$. Moreover, the following are equivalent:
  1. (i)
    $p\sim q$,
  2. (ii)
    $\mathrm{tr}\hphantom{.}p=\mathrm{tr}\hphantom{.}q$,
  3. (iii)
    For every compact subset $K\subset X$ such that $p,q$ vanish outside $K$, there is a $u\in C_{c}(X,M_{n})$ supported on $K$ such that
[…]
Proof.
Let $p\in C_{0}(X,M_{n})$ be a projection.Let $K\subset X_{\varphi}$ be a compact subset such that $\|p(x)\|<1$ for every $x\notin K$. It follows from [51, Propositions 2.2.4 and 2.2.7] that $\mathrm{tr}\hphantom{.}p(x)=0$ for all $x\notin K$. Then $\mathrm{tr}\hphantom{.}p\in C_{c}(X_{\varphi},\mathbb{N})$ and $p\in C_{c}(X,\mathbb{N})$.
The implication $\mathrm{(i)}\Rightarrow\mathrm{(ii)}$ is straightforward. We show $\mathrm{(ii)}\Rightarrow\mathrm{(iii)}$. Suppose $p\sim q$. Choose a compact subset $K$ containing the supports of $p$ and $q$. By continuity, there is a collection of elements $x_{1},\ldots,x_{m}\in K$ and a partition $V_{1},\ldots,V_{m}$ of $K$ by disjoint compact open subsets such that, for each $i=1,\ldots,m$, we have
(5.4) $x_{i}\in V_{i},\hskip 10.0pt\|p(y)-p(x_{i})\|<\tfrac{1}{2},\text{\hskip 10.0ptand }\|q(y)-q(x_{i})\|<\tfrac{1}{2},\text{\hskip 10.0ptfor }y\in V_{i}.$
Let $i=1,\ldots,m$. By assumption, $\mathrm{tr}(p(x_{i}))=\mathrm{tr}(q(x_{i}))$. Then there exists an unitary $v_{i}\in M_{n}$ such that $p(x_{i})=v_{i}^{*}q(x_{i})v_{i}$.
Let
$\widetilde{p}=\displaystyle\sum_{i=1}^{m}p(x_{i})1_{V_{i}},\text{\hskip 10.0pt}\widetilde{q}=\displaystyle\sum_{i=1}^{m}q(x_{i})1_{V_{i}}\text{\hskip 10.0ptand\hskip 10.0pt}v=\sum_{i=1}^{m}v_{i}1_{V_{i}}.$
Then $\widetilde{p}$ and $\widetilde{q}$ are projections in $C_{0}(X,M_{n})$. Also, $\widetilde{p}=v^{*}\widetilde{q}v$ and $v^{*}v=vv^{*}=\mathrm{id}1_{K}$. Considering the unital C*-algebra $C(K,M_{n})\subset C_{0}(X,M_{n})$, it follows from [51, Propositions 2.2.4 and 2.2.7] that there exists $u_{p},u_{q}\in C_{0}(X,M_{n})$ with support in $K$ such that
[…]
Note that $u_{p}$ and $u_{q}$ are unitary in $A$, not necessarily in $C_{0}(X,M_{n})$. Then
$u_{p}^{*}u_{p}=u_{p}u_{p}^{*}=\mathrm{id}1_{K}\hskip 15.0pt\text{and}\hskip 15.0ptu_{q}^{*}u_{q}=u_{q}u_{q}^{*}=\mathrm{id}1_{K}.$
Let $u=u_{q}vu_{p}$. Then $u^{*}u=uu^{*}=\mathrm{id}1_{K}$, and
$u^{*}qu=u_{p}^{*}v^{*}u_{q}^{*}qu_{q}vu_{p}=u_{p}^{*}v^{*}\widetilde{q}vu_{p}=u_{p}^{*}\widetilde{p}u_{p}=p.$
Therefore (iii) holds. To show $\mathrm{(iii)}\Rightarrow\mathrm{(i)}$, let $v=qu$. Then $v^{*}v=p$ and $vv^{*}=q$. ∎
In the following proposition we show that the C*-algebra can be locally isomorphic to C*-algebras of the form $C_{0}(V,M_{n})$, where we can use the trace operator to describe equivalence classes of projections. We also apply this isomorphism to induce a trace operator that describes the equivalence of projections in $\mathcal{P}_{\infty}(C^{*}(R(\sigma)))$.
Proposition 5.4.
Let $\varphi:Y\rightarrow X$ be a continuous section of $\sigma$. Then there is a positive integer $N$, there are disjoint open sets $V_{1},\dots,V_{N}$ with $X_{\varphi}=V_{1}\cup\dots\cup V_{N}$, there are positive integers $n_{1},\dots,n_{N}$, and there is an isomorphism $\mu:C^{*}(R(\sigma))\rightarrow\bigoplus_{i=1}^{N}C_{0}(V_{i},M_{n_{i}})$ such that
  1. (i)
    for $f\in C^{*}(R(\sigma))$ and $x\in X_{\varphi}$, we have
    $\mathrm{tr}\hphantom{.}\mu(f)(x)=\sum_{u:\sigma(u)=\sigma(x)}f(u),$
    where we apply Renualt’s $j$ map [48, Proposition II.4.2] to identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$.
  2. (ii)
    For a positive integer $k$, for $F\in M_{k}(C^{*}(R(\sigma)))$, and for $x\in X_{\varphi}$, we have
    $\mathrm{tr}\hphantom{.}\mu_{k}(F)(x)=\sum_{i=1}^{k}\mathrm{tr}\hphantom{.}\mu(F_{ii})(x),$
    where $\mu_{k}:M_{k}(C^{*}(R(\sigma)))\rightarrow\bigoplus_{i=1}^{N}C_{0}(V_{i},M_{kn_{i}})$ is the isomorphism given by
    $\mu_{k}(F)(x)=[\mu(F_{rs})(x)],\hskip 15.0pt\text{for }F=[F_{rs}]\in M_{k}(C^{*}(R(\sigma))),\text{ }x\in X_{\varphi}.$
For the next items, we fix positive integers $k,l$, and choose projections $p\in\mathcal{P}_{k}(C^{*}(R(\sigma)))$, $q\in\mathcal{P}_{l}(C^{*}(R(\sigma)))$. Then
  1. (iii)
    $\mathrm{tr}\hphantom{.}\mu_{k}(p)\in C_{c}(X_{\varphi},\mathbb{N})$.
  2. (iv)
    $\mathrm{tr}\hphantom{.}\mu_{k+l}(p\oplus q)=\mathrm{tr}\hphantom{.}\mu_{k}(p)+\mathrm{tr}\hphantom{.}\mu_{l}(q)$.
  3. (v)
    $p\sim q\Leftrightarrow\mathrm{tr}\hphantom{.}\mu_{k}(p)=\mathrm{tr}\hphantom{.}\mu_{l}(q).$
  4. (vi)
    $\mathrm{tr}\hphantom{.}\mu_{k}(p)=0\Rightarrow p=0_{k}$.
Proof.
Let $y\in Y$. By Lemma 5.2, $y$ has a clopen neighbourhood $\widetilde{W}_{y}$, there is a positive integer $n_{y}$, and there are $n_{y}$ disjoint clopen subsets $\widetilde{\mathcal{U}}_{1}^{(y)},\widetilde{\mathcal{U}}_{2}^{(y)},\dots,\widetilde{\mathcal{U}}_{n_{y}}^{(y)}\subset X$ such that
  1. (a)
    $\sigma^{-1}(\widetilde{W}_{y})=\bigcup_{j=1}^{n_{y}}\widetilde{\mathcal{U}}_{j}^{(y)}$,
  2. (b)
    $\sigma|_{\widetilde{\mathcal{U}}_{j}^{(y)}}:\widetilde{\mathcal{U}}_{j}^{(y)}\rightarrow\widetilde{W}_{y}$ is a homeomorphism for all $j$, and
  3. (c)
    $\sigma^{-1}(\widetilde{W}_{y})\cap X_{\varphi}=\widetilde{\mathcal{U}}_{1}^{(y)}.$
Since $\sigma(X_{0})$ is compact and Hausdorff, we choose $N-1$ elements $y_{1},\dots,y_{N-1}\in\sigma(X_{0})$ such that $\sigma(X_{0})=\widetilde{W}_{y_{1}}\cup\dots\cup\widetilde{W}_{y_{N-1}}$ and such that for all $i,j=1,\dots,N-1$,
(5.5) $y_{i}\in\widetilde{W}_{y_{j}}\Leftrightarrow i=j.$
Let $W_{1}=\widetilde{W}_{1}$ and, for $i=2,\dots,N-1$, define $W_{i}=\widetilde{W}_{i}\setminus\bigcup_{j=1}^{i-1}W_{j}$. It follows from property (5.5) that the sets $W_{i}$ are non-empty. Moreover, $\sigma(X_{0})$ is the disjoint union $\sigma(X_{0})=W_{1}\cup\dots\cup W_{N-1}$.
If $\sigma(X_{0})=Y$, we make an abuse of notation and replace $N-1$ by $N$. Otherwise, let $y_{N}\in Y\setminus\sigma(X_{0})$, and set $n_{N}=n_{y_{N}}=1$, $W_{N}=Y\setminus\sigma(X_{0})$, and $\mathcal{U}_{1}^{(y_{N})}=X\setminus X_{0}$. Then properties (a), (b) and (c) also hold by replacing $\widetilde{W}_{y}$, $n_{y}$ and $\widetilde{\mathcal{U}}_{j}^{(y)}$ by $W_{N}$, $n_{N}$ and $\mathcal{U}_{1}^{(y_{n})}$, respectively. In either case, we have the disjoint union
$Y=\sigma(X_{0})\cup(Y\setminus\sigma(X_{0}))=W_{1}\cup\dots\cup W_{N}.$
Define $n_{i}=n_{y_{i}}$ and $V_{i}=\mathcal{U}_{1}^{(y_{i})}$ for all $i=1,\dots,N$. Note that $V_{1},\dots,V_{N}\subset X_{\varphi}$ by property (c). Since $\sigma$ is injective on $X_{\varphi}$, and since $W_{1},\dots,W_{N}$ are disjoint, then $V_{1},\dots,V_{N}$ are disjoint subsets. Moreover,
$X_{\varphi}=\sigma^{-1}(Y)\cap X_{\varphi}=\sigma^{-1}\left(\bigcup_{i=1}^{N}W_{i}\right)\cap X_{\varphi}=\bigcup_{i=1}^{N}\sigma^{-1}(W_{i})\cap X_{\varphi}=\bigcup_{i=1}^{N}V_{i}.$
Now we use the sets above to define the map $\mu$ from $C^{*}(R(\sigma))$ to $\bigoplus_{i=1}^{N}C_{0}(V_{i},M_{n_{i}})$ by
(5.6) $\mu(f)(x)_{rs}=f(\sigma|_{\mathcal{U}_{r}^{(y_{i})}}^{-1}(\sigma(x)),\sigma|_{\mathcal{U}_{s}^{(y_{i})}}^{-1}(\sigma(x))),$
for $x\in V_{i}$, $r,s=1,\dots,n_{i}$, $i=1,\dots,n_{i}$, $f\in C^{*}(R(\sigma))$.
It is straightforward that $\mu$ is well-defined. We show that $\mu$ is an $\ast$-isomorphism.
  • $\mu$ is a homomorphism
    In order to simplify our notation, given $i=1,\dots,N$, $x\in\sigma^{-1}(W_{i})$, and $r=1,\dots,n_{i}$, we denote $x_{r}=\sigma|_{\mathcal{U}_{r}^{(y_{i})}}^{-1}(\sigma(x))$. Note that $\sigma^{-1}(\{\sigma(x)\})=\{x_{1},\dots,x_{n_{i}}\}.$ Let $f,f_{1},f_{2}\in C^{*}(R(\sigma))$ and let $x\in V_{i}$ with $i=1,\dots,N$. Then, for $r,s=1,\dots,n_{i}$,
    $\mu(f^{*})(x)_{rs}=f^{*}(x_{r},x_{s})=\overline{f(x_{s},x_{r})}=(\mu(f)(x)^{*})_{rs}.$
    By properties (a) and (b), we have that an element $u\in X$ satisfies $\sigma(u)=\sigma(x)$ if, and only if, $u=x_{j}$ for some $j=1,\dots,n_{i}$. Then
    $\mu(f_{1}\cdot f_{2})(x)_{rs}=f_{1}\cdot f_{2}(x_{r},x_{s})=\sum_{t=1}^{n_{i}}f_{1}(x_{r},x_{t})f_{2}(x_{t},x_{s})=(\mu(f_{1})\mu(f_{2}))(x)_{rs}.$
    Thus $\mu$ is a homomorphism.
  • $\mu$ is an isometry
    By [55, Proposition 9.3.1], for all $x\in X$, the $\ast$-representation $\pi_{x}:C_{c}(R(\sigma))\rightarrow B(\ell^{2}(R(\sigma)_{x}))$ is given by
    $\pi_{x}(f)\delta_{(v,x)}=\sum_{(u,v)\in R(\sigma)}f(u,v)\delta_{(u,x)}=\sum_{u:\sigma(u)=\sigma(x)}f(u,v)\delta_{(u,x)},$
    for $f\in C_{c}(R(\sigma))$ and $v\in X$ with $\sigma(v)=\sigma(x)$.
    Let $x\in X_{\varphi}$. Then there exists a unique $i=1,\dots,N$ such that $x\in V_{i}$. To simplify our notation, we let $x_{r}=\sigma_{\mathcal{U}_{r}^{(y_{i})}}^{-1}(\sigma(x))$ for all $r=1,\dots,n_{i}$. Note that $\sigma^{-1}(\sigma(x))=\{x_{1},\dots,x_{n_{i}}\}$.
    Let $T_{x}:\mathbb{C}^{n_{i}}\rightarrow\ell^{2}(R(\sigma)_{x})$ be the isometry defined by $T_{x}e_{r}=\delta_{(x_{r},x)}$ for $r=1,\dots,n_{i}$. Then, for all $s=1,\dots,n_{i}$, we have
    $\pi_{x}(f)T_{x}e_{s}$ $=\pi_{x}(f)\delta_{(x_{s},x)}$
    $=\sum_{u:\sigma(u)=\sigma(x)}f(u,x_{s})\delta_{(u,x)}$
    $=\sum_{r=1}^{n_{i}}f(x_{r},x_{s})\delta_{(x_{r},x)}$
    $=T_{x}\sum_{r=1}^{n_{1}}f(x_{r},x_{s})e_{r}$
    $=T_{x}\mu(f)(x)e_{s}.$
    This implies that $\|\pi_{x}(f)\|=\|\mu(f)(x)\|$ for all $x\in X_{\varphi}$.
    Given $x\in X$, let $u=\varphi(\sigma(x))\in X_{\varphi}$. [55, Proposition 9.3.1] implies that there exists a unitary $U_{(x,u)}:\ell^{2}(R(\sigma)_{u})\rightarrow\ell^{2}(R(\sigma)_{x})$ such that $\pi_{x}=U_{(x,u)}\pi_{u}U_{(x,u)}^{*}$. Then $\|\pi_{x}(f)\|=\|\pi_{u}(f)\|$. Therefore,
    $\|f\|=\sup_{x\in X}\|\pi_{x}(f)\|=\sup_{x\in X_{\varphi}}\|\pi_{x}(f)\|=\sup_{x\in X_{\varphi}}\|\mu(f)(x)\|=\|\mu(f)\|.$
    Therefore the restriction of $\mu$ to $C_{c}(R(\sigma))$ is an isometry. By continuity, $\mu$ is an isometry.
  • $\mu$ is an isomoprism
    Since $\mu$ is an isometry, it follows that it is injective. We show that $\mu$ is surjective. Let $F\in\bigoplus_{i=1}^{N}C_{c}(V_{i},M_{n_{i}})$. Define $f\in C_{c}(R(\sigma))$ by
    $f(x_{r},x_{s})=F(x)_{rs}\hskip 15.0pt\text{for $x\in V_{i}$, $r,s=1,\dots,n_{i}$.}$
    Note that $f$ is continuous because it is the composition of continuous functions. Let $K\subset X_{\varphi}$ by a compact subset containing the support of $F$, then the support of $f$ is a subset of $[\sigma^{-1}(\sigma(K))\times\sigma^{-1}(\sigma(K))]\cap R(\sigma)$, which is compact. This implies that $f\in C_{c}(R(\sigma))$. Moreover, $\mu(f)=F$. So far we have that the restriction
    $\mu|_{C_{c}(R(\sigma))}:C_{c}(R(\sigma))\rightarrow\bigoplus_{i=1}^{N}C_{c}(V_{i},M_{n_{i}})$
    is surjective. Since $\bigoplus_{i=1}^{N}C_{c}(V_{i},M_{n_{i}})$ is dense in $\bigoplus_{i=1}^{N}C_{0}(V_{i},M_{n_{i}})$ and since $\mu$ is an isometry, then $\mu$ is surjective.
Now we show property (i). Let $f\in C_{c}(R(\sigma))$, and $x\in V_{i}$ with $i=1,\dots,N$. Since $\sigma^{-1}(x)=\{x_{1},\dots,x_{n_{i}}\}$, then
$\mathrm{tr}\hphantom{.}\mu(f)(x)=\sum_{l=1}^{n_{i}}f(x_{l},x_{l})=\sum_{u:\sigma(u)=\sigma(x)}f(u).$
Now we prove property (ii). Let $k$ be a positive integer, and let $F=[F_{rs}]\in M_{k}(C^{*}(R(\sigma)))$, and fix $x\in V_{i}$ with $i=1,\dots,N$. Then
$\mathrm{tr}\hphantom{.}\mu_{k}(F)(x)=\sum_{q=1}^{kn_{i}}\mu_{k}(F)_{qq}(x)=\sum_{j=0}^{k-1}\sum_{l=1}^{n_{i}}\mu_{k}(F)_{n_{i}j+l,n_{i}j+l}(x)=\sum_{i=1}^{k}\sum_{l=1}^{n_{i}}\mu(F_{ii})_{ll}(x).$
Thus property (ii) holds.
Next, we prove item (iii). Let $k$ be a positive integer, and fix $p\in\mathcal{P}_{k}(C^{*}(R(\sigma)))$. Note that $C^{*}(R(\sigma))\subset C(R(\sigma))$, so it follows that, for all $i=1,\dots,k$, $p_{ii}\in C(R(\sigma))$. given $x\in X_{\varphi}$, items (i) and (ii) imply that
$\mathrm{tr}\hphantom{.}\mu_{k}(p)(x)=\sum_{i=1}^{k}\mathrm{tr}\hphantom{.}\mu(p_{ii})(x)=\sum_{i=1}^{k}\sum_{u:\sigma(u)=\sigma(x)}p_{ii}(u).$
The equation above implies that $\mathrm{tr}\hphantom{.}\mu_{k}(p)$ is a composition of continuous functions, then $\mathrm{tr}\hphantom{.}\mu_{k}(p)$ is a continuous function. It follows from Lemma 5.3 that $\mathrm{tr}\hphantom{.}\mu_{k}(p)\in C_{c}(X_{\varphi},\mathbb{N})$.
Now we show item (iv). Let $x\in X_{\varphi}$. By definition of $\mu$,
$\mu_{k+l}(p\oplus q)(x)=\mu_{k}(p)(x)\oplus\mu_{l}(q)(x).$
Then item (ii) implies that $\mathrm{tr}\hphantom{.}\mu(p\oplus q)(x)=\mathrm{tr}\hphantom{.}\mu_{k}(p)(x)+\mathrm{tr}\hphantom{.}\mu_{k}(q)(x).$
Next, we show (v). Suppose $p\sim q$. Then $p\oplus 0_{l},$ $q\oplus 0_{k}\in\mathcal{P}_{l}(C^{*}(R(\sigma)))$ are equivalent. Given $x\in X_{\varphi}$, then $\mu_{k+l}(p\oplus 0_{l}),\mu_{k+l}(q\oplus 0_{k})$ are equivalent in $M_{k+l}$. This fact and item (iv) imply that
$\mathrm{tr}\hphantom{.}\mu_{k}(p)(x)=\mathrm{tr}\hphantom{.}\mu_{k+l}(p\oplus 0_{l})(x)=\mathrm{tr}\hphantom{.}\mu_{k+l}(q\oplus 0_{k})(x)=\mathrm{tr}\hphantom{.}\mu_{l}(q)(x).$
Conversely, assume that $\mathrm{tr}\hphantom{.}\mu_{k}(p)=\mathrm{tr}\hphantom{.}\mu_{l}(q)$. By item (iv), $\mathrm{tr}\hphantom{.}\mu_{k+l}(p\oplus 0_{l})=\mathrm{tr}\hphantom{.}\mu_{k+l}(q\oplus 0_{k})$. Hence, by Proposition 5.3, $\mu_{k+l}(p\oplus 0_{l})\sim\mu_{k+l}(q\oplus_{k})$. Since $\mu_{k+l}$ is an isomorphism, we have $p\oplus 0_{l}\sim q\oplus 0_{k}$ and then $p\sim q$.
Finally, we prove item (vi). Suppose $\mathrm{tr}\hphantom{.}\mu_{k}(p)=0$. Then, for all $x\in X_{\varphi}$, $0=\mathrm{tr}\hphantom{.}\mu_{k}(p)(x)=\dim\mu_{k}(p(x))\mathbb{C}^{k}.$ So $\mu_{k}(p(x))=0$. Since $x$ is arbitrary and $\mu_{k}$ is an isomorphism, we have $p=0$. ∎
We use the isomorphism $\mu$ to define a trace operator on $\mathcal{P}_{\infty}(C^{*}(R(\sigma)))$ that characterises the $K_{0}$-group of $C^{*}(R(\sigma))$. Given a continuous section $\varphi:Y\rightarrow X$ of $\sigma$, we define $\mathrm{tr}_{\varphi}:\mathcal{P}_{\infty}(C^{*}(R(\sigma)))\rightarrow C(X,\mathbb{N})$ by
$\mathrm{tr}_{\varphi}\hphantom{.}p(x)=\begin{cases}\mathrm{tr}\hphantom{.}\mu_{k}(p)(x)&\text{if }x\in X_{\varphi},\\ 0&\text{otherwise,}\end{cases}$
for $p\in\mathcal{P}_{k}(C^{*}(R(\sigma)))$, $k\geq 1$, and $x\in X$. Note that $\mathrm{tr}_{\varphi}\hphantom{.}p$ is the extension of $\mathrm{tr}\hphantom{.}\mu_{k}(p)$ to a larger domain $X\supset X_{\varphi}$. We do this because later we define the isomorphism from $K_{0}(C^{*}(R(\sigma)))$ to $H_{0}(R(\sigma))$ by $[p]_{0}\mapsto[\mathrm{tr}_{\varphi}\hphantom{.}p]$ for $p\in\mathcal{P}_{\infty}(C^{*}(R(\sigma)))$. In order to study the equivalence class of $\mathrm{tr}_{\varphi}\hphantom{.}p$ in the homology group of $H_{0}(R(\sigma))$, the function $\mathrm{tr}_{\varphi}\hphantom{.}p$ needs to be defined in the unit space $X$ of $R(\sigma)$.
Remark 5.5.
Note that, if the projection $p$ is in $C^{*}(R(\sigma))$, then for an arbitrary continuous section $\varphi$ of $\sigma$, $x\in X$, and $f\in C^{*}(R(\sigma))$
$\mathrm{tr}_{\varphi}\hphantom{.}p(x)=1_{X_{\varphi}}(x)\sum_{u:\sigma(u)=\sigma(x)}f(u).$
The following lemma shows the connection of the trace operator $\mathrm{tr}_{\varphi}$ with the homology group $H_{0}(R(\sigma))$.
Lemma 5.6.
Let $\varphi,\psi$ be two continuous sections of $\sigma$. Then
  1. (i)
    Given $p\in\mathcal{P}_{\infty}(C^{*}(R(\sigma)))$, $\mathrm{tr}_{\varphi}\hphantom{.}p\sim\mathrm{tr}_{\psi}\hphantom{.}p$.
  2. (ii)
    For a compact open set $V\subset X$, we have $1_{V}\sim\mathrm{tr}_{\varphi}1_{V}$.
Proof.
We begin by proving (i). Fix $p\in\mathcal{P}_{\infty}(C^{*}(R(\sigma)))$, then there exists a positive integer $k$ with $p\in\mathcal{P}_{k}(C^{*}(R(\sigma)))$. Given $x\in X$, set $x_{\varphi}=\varphi(\sigma(x))$ and $x_{\psi}=\psi(\sigma(x))$. Then $x_{\varphi}$ (resp. $x_{\psi}$) is the only element in $X_{\varphi}$ (resp. $X_{\psi}$) with $\sigma(x_{\varphi})=\sigma(x_{\psi})=\sigma(x)$. By definition of $\mathrm{tr}_{\varphi}\hphantom{.}p$ and by Proposition 5.4, we have
$\sum_{u:\sigma(u)=\sigma(x)}\mathrm{tr}_{\varphi}\hphantom{.}p(u)$ $=\mathrm{tr}_{\varphi}\hphantom{.}p(x_{\varphi})=\sum_{i=1}^{k}\mathrm{tr}_{\varphi}\hphantom{.}p_{ii}(x_{\varphi})=\sum_{i=1}^{k}\sum_{v:\sigma(v)=\sigma(x_{\varphi})}p_{ii}(v)$
$=\sum_{i=1}^{k}\sum_{v:\sigma(v)=\sigma(x_{\psi})}p_{ii}(v)=\mathrm{tr}_{\psi}\hphantom{.}p(x_{\psi})=\sum_{u:\sigma(u)=\sigma(x)}\mathrm{tr}_{\psi}\hphantom{.}p(u).$
Since $x$ is arbitrary, it follows from Lemma 4.11 that $\mathrm{tr}_{\varphi}\hphantom{.}p\sim\mathrm{tr}_{\psi}\hphantom{.}p$.
Now we prove item (ii). Let $V\subset X$ a compact open subset, let $x\in X$, and set $x_{\varphi}=\varphi(\sigma(x))$. Then
$\sum_{u:\sigma(u)=\sigma(x)}\mathrm{tr}_{\varphi}\hphantom{.}1_{V}(u)=\mathrm{tr}_{\varphi}\hphantom{.}1_{V}(x_{\varphi})=\sum_{u:\sigma(u)=\sigma(x_{\varphi})}1_{V}(u)=\sum_{u:\sigma(u)=\sigma(x)}1_{V}(u).$
Since $x$ is arbitrary, Lemma 4.11 implies that $\mathrm{tr}_{\varphi}\hphantom{.}1_{V}\sim 1_{V}$. ∎
The following theorem gives an isomorphism of the ordered groups $K_{0}(C^{*}(G))$ and $H_{0}(G)$. Corollary 5.2 by Farsi, Kumjian, Pask and Sims [20] and Matui’s [39, Theorem 4.10] prove isomorphisms of homology groups of groupoids and the corresponding $K_{0}$-group. However, [20, Corollary 5.2] does not prove whether such isomorphism preseves positive elements from one group to another, and Matui only considers groupoids with compact unit spaces, without providing a detailed proof. Our theorem generalises both results, and we describe the isomorphism by an explicit formula.
Theorem 5.7.
Let $G$ be an AF groupoid, written as the inductive limit $G=\displaystyle\lim_{\rightarrow}G_{n}$ of elementary groupoids. Then, for every $n$, there exists a locally compact, Hausdorff, second countable, totally disconntect space $Y_{n}$, and there is a surjective local homeomorphism $\sigma_{n}:G_{n}^{(0)}\rightarrow Y_{n}$ such that $G_{n}=R(\sigma_{n})$. Moreover,
  1. (1)
    there exists a unique isomorphism $\nu:K_{0}(C^{*}(G))\rightarrow H_{0}(G)$ defined by
    $\nu([p]_{0})=[\mathrm{tr}_{\varphi_{n}}(p)]$
    for $p\in\mathcal{P}_{n}(C^{*}(G_{n}))$, and such that $\varphi_{n}$ is an arbitrary continuous section of $\sigma_{n}$;
  2. (2)
    the map $\nu$ is an isomorphism of ordered groups, i.e., $\nu(K_{0}(C^{*}(G))^{+})=H_{0}(G)^{+}$; and
  3. (3)
    the isomorphism $\nu$ is precisely the map given by [20, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
We divide the proof into three parts:
  1. Part 1
    First we study the groupoid $R(\sigma)$ where $\sigma$ satisfies the conditions of the beginning of the chapter, i.e., that there exists a compact subset $X_{0}$ such that $\sigma^{-1}(\sigma(X_{0}))=X_{0}$ and that $\sigma$ is injective on $X\setminus X_{0}$.
  2. Part 2
    Then we generalise the result to a groupoid of the form $R(\sigma)$, where $X$ no longer need to necessarily satisfy the conditions of Part 1.
  3. Part 3
    Finally, we prove the isomorphism for an AF groupoid $G$.
Proof of Part 1.
Fix a continuous section $\varphi$ of $\sigma$. Given $p\in\mathcal{P}_{\infty}(C^{*}(R(\sigma)))$, let $\nu([p]_{0})=[\mathrm{tr}_{\varphi}\hphantom{.}p]$. It follows from item (v) of Proposition 5.4 that $\nu([p]_{0})$ is well-defined. Moreover, item (iv) of the same proposition implies that for a projection $q\in\mathcal{P}_{\infty}(C^{*}(R(\sigma)))$,
$\nu([p]_{0}+[q]_{0})=\nu([p\oplus q]_{0})=[\mathrm{tr}_{\varphi}(p\oplus q)]=[\mathrm{tr}_{\varphi}(p)]+[\mathrm{tr}_{\varphi}(q)]=\nu([p]_{0})+\nu([q]_{0}).$
Then we extend $\nu$ uniquely to a group homomorphism $\nu:K_{0}(C^{*}(R(\sigma)))\rightarrow H_{0}(R(\sigma))$.
We claim that $\nu$ is an isomorphism. Let $x\in K_{0}(C^{*}(R(\sigma)))$ be such that $\nu(x)=0$. Then there are $p,q\in\mathcal{P}_{\infty}(C^{*}(R(\sigma)))$ such that $x=[p]_{0}-[q]_{0}$. Then
$[\mathrm{tr}_{\varphi}\hphantom{.}p]=\nu([p]_{0})=\nu([q]_{0})=[\mathrm{tr}_{\varphi}\hphantom{.}q],$
by definition of the homomorphism $\nu$. It follows from Lemma 4.11 that, for every $x\in X$,
$\sum_{u:\sigma(u)=\sigma(x)}\mathrm{tr}_{\varphi}\hphantom{.}p(u)=\sum_{u:\sigma(u)=\sigma(x)}\mathrm{tr}_{\varphi}\hphantom{.}q(u).$
This equivalence implies that
$\mathrm{tr}_{\varphi}\hphantom{.}p(\varphi(\sigma(x)))=\mathrm{tr}_{\varphi}\hphantom{.}q(\varphi(\sigma(x))).$
Since $\mathrm{tr}_{\varphi}\hphantom{.}p$ and $\mathrm{tr}_{\varphi}\hphantom{.}q$ have support in $X_{\varphi}$, then $\mathrm{tr}_{\varphi}\hphantom{.}p=\mathrm{tr}_{\varphi}\hphantom{.}q$. It follows from item (v) of Proposition 5.4 that $p\sim q$. Hence, $x=0$ and therefore $\nu$ is injective. We show that $\nu$ is also surjective. Given a compact open subset $V\subset X$, Lemma 5.6 implies that
$[1_{V}]=[\mathrm{tr}_{\varphi}\hphantom{.}1_{V}]=\nu([1_{V}]_{0}).$
Since the elements of the form $[1_{V}]$ generate $H_{0}(R(\sigma))$, it follows that $\nu$ is a group isomorphism.
Next, we prove that $\nu$ is an isomorphism of ordered groups. Item (iii) of Proposition 5.4 implies that $\nu(K_{0}(C^{*}(R(\sigma))^{+})\subset H_{0}(R(\sigma))^{+}$. Let $f\in C_{c}(X,\mathbb{N})$. Then there are compact open sets $V_{1},\dots,V_{n}\subset X$, and non-negative integers $a_{1},\dots,a_{n}$ such that $f=a_{1}1_{V_{1}}+\dots+a_{n}1_{V_{n}}$. Then
$[f]=\sum_{i=1}^{n}a_{i}[1_{V_{i}}]=\sum_{i=1}^{n}a_{i}\nu([1_{V_{i}}]_{0})=\nu\left(\sum_{i=1}^{n}a_{i}[1_{V_{i}}]_{0}\right).$
Then $\nu(K_{0}(C^{*}(R(\sigma)))^{+})=H_{0}(R(\sigma))^{+}$. This completes the proof of Part 1. ∎
We need the following lemma before proving Part 2. Now we no longer assume that $\sigma$ satisfies that conditions of Part 1.
Lemma 5.8.
Let $\sigma:X\rightarrow Y$ be a surjective local homeomorphism with continuous section $\varphi:Y\rightarrow X$. Let $X=\bigcup_{k=1}^{\infty}X_{k}$ be an increasing union of compact open sets such that $\varphi(\sigma(X_{k}))\subset X_{k}$. For each $k$, define the map $\sigma_{k}:X\rightarrow\sigma(X_{k})\sqcup(X\setminus X_{k})$ by
$\sigma_{k}(x)=\begin{cases}\sigma(x)&\text{if }x\in X_{k},\\ x&\text{otherwise.}\end{cases}$
Define also the map $\varphi_{k}:\sigma(X_{k})\sqcup(X\setminus X_{k})\rightarrow X$ by
$\varphi_{k}(y)=\begin{cases}\varphi(y)&\text{if }y\in\sigma(X_{k}),\\ y&\text{otherwise.}\end{cases}$
Then
  1. (i)
    $\sigma_{k}$ is a surjective local homeomorphism with continuous section $\varphi_{k}$.
  2. (ii)
    $R(\sigma)$ is an inductive limit of the form $R(\sigma)=\displaystyle\lim_{\rightarrow}R(\sigma_{k})$.
  3. (iii)
    For $k\geq 1$, $p\in\mathcal{P}_{\infty}(C^{*}(R(\sigma_{k}))$, we have $\mathrm{tr}_{\varphi}\hphantom{.}p=\mathrm{tr}_{\varphi_{l}}\hphantom{.}p$ for all $l\geq k$, where $\mathrm{tr}_{\varphi}\hphantom{.}p$ is given by
    $\mathrm{tr}_{\varphi}\hphantom{.}p(x)=1_{X_{\varphi}}(x)\sum_{i=1}^{n}\sum_{u:\sigma(u)=\sigma(x)}p_{ii}(u),\hskip 15.0pt\text{for $x\in X$.}$
Proof.
Since $\sigma$ is a surjective local homeomorphism and $X_{k}$ is open, then the restriction $\sigma_{k}|_{X_{k}}:X_{k}\rightarrow\sigma(X_{k})$ is a surjective local homeormorphism. The restriction $\sigma_{k}$ on $X\setminus X_{k}$ is the identity. Thus, $\sigma_{k}$ is a surjective local homeomorphism. It is straightforward that $\varphi_{k}$ is a continuous section of $\sigma_{k}$. This proves (i).
Next we prove (ii). Let $(x,y)\in R(\sigma)$. Since $X$ is covered by the union of the sets $X_{k}$, then there exists a $k$ such that $x,y\in X_{k}$. By definition of $\sigma_{k}$, we have that $\sigma_{k}(x)=\sigma_{k}(y)$, and then $(x,y)\in R(\sigma_{k})$. Therefore, $R(\sigma)\subset\bigcup_{k=1}^{\infty}R(\sigma_{k})$. By similar arguments, we have that $\bigcup_{k=1}^{\infty}R(\sigma_{k})\subset R(\sigma)$. It is straightforward that $R(\sigma_{k})$ is an open subset of $R(\sigma_{k+1})$ for all $k$. Therefore, (ii) holds.
Finally, we show (iii). Let $k\geq 1$, and let $p\in\mathcal{P}_{\infty}(C_{c}(R(\sigma_{k})))$. Then there exists an $n$ such that $p\in\mathcal{P}_{n}(C_{c}(R(\sigma)))$. By increasing $n$ if necessary, we have $p\in\mathcal{P}_{n}(C_{c}(R(\sigma_{n})))$ and that the support of $p$ is included in $X_{n}$. Then for any $l\geq n$ and $x\in X$
$\mathrm{tr}_{\varphi_{l}}\hphantom{.}p(x)=1_{X_{\varphi_{l}}}(x)\sum_{i=1}^{n}\sum_{u:\sigma_{l}(u)=\sigma_{l}(x)}p_{ii}(u)=1_{X_{\varphi_{l}}}(x)\sum_{i=1}^{n}\sum_{u\in X_{l}:\sigma_{l}(u)=\sigma_{l}(x)}p_{ii}(u)$
while
$\mathrm{tr}_{\varphi}\hphantom{.}p(x)=1_{X_{\varphi}}(x)\sum_{i=1}^{n}\sum_{u:\sigma(u)=\sigma(x)}p_{ii}(u)=1_{X_{\varphi}}(x)\sum_{i=1}^{n}\sum_{u\in X_{l}:\sigma(u)=\sigma(x)}p_{ii}(u)$
Therefore $\mathrm{tr}_{\varphi}\hphantom{.}p(x)=\mathrm{tr}_{\varphi_{l}}\hphantom{.}p(x)$.
Now let $p\in\mathcal{P}_{n}(C^{*}(R(\sigma)))$. Since $C^{*}(R(\sigma))=\displaystyle\lim_{\rightarrow}C^{*}(R(\sigma_{m}))$, then $K_{0}(C^{*}(R(\sigma)))=\displaystyle\lim_{\rightarrow}K_{0}(C^{*}(R(\sigma_{m})))$ as ordered groups. Then there is $q\in\mathcal{P}_{\infty}(C^{*}(R(\sigma_{m})))$ with $p\sim q$. By Part 1 and by item (v) of Proposition 5.4, we have $\mathrm{tr}_{\varphi_{m}}\hphantom{.}p=\mathrm{tr}_{\varphi_{m}}\hphantom{.}q$. By applying the formula of item 3 for the isomorphism $K_{0}(C^{*}(R(\sigma_{m})))\cong H_{0}(R(\sigma_{m}))$, we can assume that $q\in C_{c}(R(\sigma_{m}))$ without loss of generality. Then $q\in C_{c}(R(\sigma))$. Hence, there exists a $k$ such that $\mathrm{tr}_{\varphi}\hphantom{.}q=\mathrm{tr}_{\varphi_{l}}\hphantom{.}q$ for all $l\geq k$. Then
(5.7) $\mathrm{tr}_{\varphi}\hphantom{.}q=\mathrm{tr}_{\varphi_{l}}\hphantom{.}q=\mathrm{tr}_{\varphi_{l}}\hphantom{.}p.$
Let $x\in X$. Choose $j\geq l$ such that $X_{j}$ contains both $x$ and $\sigma(x)$. Then
$\mathrm{tr}_{\varphi}\hphantom{.}p(x)=1_{X_{\varphi}}(x)\sum_{i=1}^{n}\sum_{u:\sigma_{j}(u)=\sigma_{j}(x)}p_{ii}(u)=\mathrm{tr}_{\varphi_{j}}\hphantom{.}p(x)=\mathrm{tr}_{\varphi_{l}}\hphantom{.}p(x).\qed$
Proof of Part 2.
Fix a continuous section $\varphi$ of $\sigma$. Since $X$ is second countable and totally disconnected, then we write $X=\bigcup_{k=1}^{\infty}\widetilde{X}_{k}$ as the increasing union of compact open sets $\widetilde{X}_{k}$. For each $k$, let $X_{k}=\widetilde{X_{k}}\cap\varphi(\sigma(X_{k}))$. Then $X=\bigcup_{k=1}^{\infty}X_{k}$ is the increasing union of the compact open sets $X_{k}$, with $\varphi(\sigma(X_{k}))\subset X_{k}$. Lemma 5.8 gives a sequence of surjective local homeomorphisms $\sigma_{k}:X_{k}\rightarrow\sigma(X_{k})\sqcup(X\setminus X_{k})$ satisfying the properties of the lemma.
By Part 1, we have the isomorphism $\nu_{k}:K_{0}(C^{*}(R(\sigma_{k})))\rightarrow H_{0}(R(\sigma_{k}))$ of ordered groups given by $\nu_{k}([p]_{0,k})=[\mathrm{tr}_{\varphi_{k}}\hphantom{.}p]_{k}$ for $p\in\mathcal{P}_{\infty}(C^{*}(R(\sigma_{k})))$. This isomorphism is such that $[1_{V}]_{0,k}\mapsto[1_{V}]_{k}$ for all compact open sets $V\subset X$.
Denote $[\cdot]_{0,k}$, $[\cdot]_{0}$, $[\cdot]_{k}$ and []delimited-[][$\cdot$][ ⋅ ] the equivalence classes in the groups
$K_{0}(C^{*}(R(\sigma_{k}))),\hskip 10.0ptK_{0}(C^{*}(R(\sigma))),\hskip 10.0ptH_{0}(R(\sigma_{k}))\hskip 10.0pt\text{ and }\hskip 10.0ptH_{0}(R(\sigma)),$
respectively. Recall that $K_{0}(C^{*}(R(\sigma)))=\displaystyle\lim_{\rightarrow}K_{0}(C^{*}(R(\sigma_{k}))$ and $H_{0}(R(\sigma))=\displaystyle\lim_{\rightarrow}H_{0}(R(\sigma_{k}))$. For $k\geq 1$, let
$i_{k}:K_{0}(C^{*}(R(\sigma_{k})))\rightarrow K_{0}(C^{*}(R(\sigma_{k+1})))\hskip 5.0pt\text{and}\hskip 5.0ptj_{k}:H_{0}(R(\sigma_{k}))\rightarrow H_{0}(R(\sigma_{k+1}))$
be the connecting morphisms, and let $\alpha_{k}:K_{0}(C^{*}(R(\sigma_{k})))\rightarrow K_{0}(C^{*}(R(\sigma)))$ and $\beta_{k}:H_{0}(R(\sigma_{k}))\rightarrow H_{0}(R(\sigma))$ be the inclusion morphisms such that
$i_{k}([p]_{0,k})=[p]_{0,k+1}\hskip 10.0pt\text{and}\hskip 10.0pt\alpha_{k}([p]_{0,k})=[p]_{0}\hskip 10.0pt\text{for }p\in\mathcal{P}_{\infty}(C^{*}(R(\sigma_{k}))),$
$j_{k}([f]_{k})=[f]_{k+1}\hskip 13.0pt\text{and}\hskip 17.0pt\beta_{k}([f]_{k})=[f]\hskip 15.0pt\text{for }f\in C(X,\mathbb{Z}).\hskip 40.0pt$
The morphisms $i_{k}$, $j_{k}$, $\alpha_{k}$, $\beta_{k}$ preserve positive elements from one group to another. It is straightforward that $j_{k}\circ\nu_{k}([1_{V}]_{0,k})=\nu_{k+1}\circ i_{k}([1_{V}]_{0,k})$ for the elements $V$ compact open. Since the elements $[1_{V}]_{0,k}$ generate $K_{0}(C^{*}(R(\sigma_{k})))$, we have $j_{k}\circ\nu_{k}=\nu_{k+1}\circ i_{k}$. Lemma 2.1 gives an isomorphism $\nu:K_{0}(C^{*}(R(\sigma)))\rightarrow H_{0}(R(\sigma))$ that makes the diagram below commutative for all $k$.
${K_{0}(C^{*}(R(\sigma_{k}))}$${H_{0}(R(\sigma_{k}))}$${K_{0}(C^{*}(R(\sigma))}$${H_{0}(R(\sigma))}$$\scriptstyle{\nu_{k}}$$\scriptstyle{\alpha_{k}}$$\scriptstyle{\beta_{k}}$$\scriptstyle{\nu}$
Then, for all compact open subset $V\subset X$ we have that
$\nu([1_{V}]_{0})=\nu\circ\alpha_{k}([1_{V}]_{0,k})=\beta_{k}\circ\nu_{k}([1_{V}]_{0,k})=\beta_{k}([1_{V}]_{k})=[1_{V}].$
Now let $n\geq 1$ and $p\in\mathcal{P}_{n}(C^{*}(R(\sigma)))\subset K_{0}(C^{*}(R(\sigma)))^{+}$. By Lemma 5.8, there exists a $k$ such that $\mathrm{tr}_{\varphi_{k}}\hphantom{.}p=\mathrm{tr}_{\varphi}\hphantom{.}p$. It follows from Part 1 that
(5.8) $\nu([p]_{0})=\nu(\alpha_{k}([p]_{0,k})=\beta_{k}\circ\nu_{k}([p]_{0,k})=\beta_{k}([\mathrm{tr}_{\varphi_{k}}\hphantom{.}p]_{k})=[\mathrm{tr}_{\varphi_{k}}\hphantom{.}p]=[\mathrm{tr}_{\varphi}\hphantom{.}p].$
This completes the proof of Part 2. ∎
Proof of Part 3.
The proof is analogous to Part 2. Let $G$ be an AF groupoid. Write $G=\displaystyle\lim_{\rightarrow}R(\sigma_{n})$, where $\sigma:G^{(0)}\rightarrow Y_{n}$ is a surjective local homeomorphism of totally disconnected spaces. As in Part 2, let $i_{n}:K_{0}(C^{*}(R(\sigma_{n})))\rightarrow K_{0}(C^{*}(R(\sigma_{n+1})))$ and $j_{n}:H_{0}(R(\sigma_{n}))\rightarrow H_{0}(R(\sigma_{n+1}))$ be the inclusion maps. By Part 2, we have an isomorphism $\nu_{n}:K_{0}(C^{*}(R(\sigma_{n})))\rightarrow H_{0}(R(\sigma_{n}))$ satisfying the properties 1–3 of this theorem for all $n$. Lemma 2.1 gives an isomorphism $\nu:K_{0}(C^{*}(G))\rightarrow H_{0}(G)$ of ordered groups such that the we have the following commutative diagram for all $n$, where $\alpha_{n}$ and $\beta_{n}$ are the inclusion maps.
${K_{0}(C^{*}(R(\sigma_{n}))}$${H_{0}(R(\sigma_{n}))}$${K_{0}(C^{*}(G)}$${H_{0}(G)}$$\scriptstyle{\nu_{n}}$$\scriptstyle{\alpha_{n}}$$\scriptstyle{\beta_{n}}$$\scriptstyle{\nu}$
Using the same ideas to prove (5.8) in Part 2, we have that $\nu$ satisfies properties 1–3. This completes the proof of the theorem. ∎
Corollary 5.9.
Given an AF groupoid $G$, $K_{0}(C^{*}(G))^{+}$ is generated by elements of the form $[1_{V}]_{0}$, where $V\subset G^{(0)}$ are compact open subsets.
Proof.
Recall that $H_{0}(G)^{+}=\{[f]:f\in C_{c}(X,\mathbb{N})\}$. Since $G^{(0)}$ is locally compact, Hausdorff, second countable, and totally disconnected, then every $f\in C_{c}(G^{(0)},\mathbb{N})$ is a linear combination of the form $f=\alpha_{1}1_{V_{1}}+\dots+\alpha_{n}1_{V_{n}}$, where all $\alpha_{i}$ non-negative integers, and $V_{i}$ are compact open. This implies that the classes of the form $[1_{V}]$, for $V\subset G^{(0)}$ compact open, generate $H_{0}(G)^{+}$. By Theorem 5.7, there is an ordered group isomorphism $K_{0}(C^{*}(G))\rightarrow H_{0}(G))^{+}$ that maps $[1_{V}]_{0}$ to $[1_{V}]$. Since the equivalence classes $[1_{V}]$, for $V\subset G^{(0)}$ compact open, generate $H_{0}(G)^{+}$, then the corresponding equivalence classes $[1_{V}]_{0}$ generate $K_{0}(C^{*}(G))^{+}$. ∎
6. The Commutative Diagram
Recall from the introduction that we want to find a group which is isomorphic to $K_{0}(B)$ and more manageable, where $\mathcal{G}$ is a Deaconu-Renault groupoid, $B$ is the C*-algebra of the skew product groupoid $\mathcal{G}(c)$, and $c:\mathcal{G}\rightarrow\mathbb{Z}$ is the continuous cocycle given by $c(x,k,y)=k$, called the canonical cocycle. By studying this new group which is isomorhic to $K_{0}(B)$, we will find a condition on $\sigma$ that is equivalent to the condition $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$ from Theorem 1.1.
We prove the diagram below by using the techniques for homology groups described in Section 4, and by applying isomorphism from Theorem 5.7.
[…]
The action $\beta$ is described in Proposition 3.4. For a Deaconu-Renault groupoid $\mathcal{G}$, the action $\beta\in\mathrm{Aut}(B)$ is such that $\beta(f)(x,k,y,a)=f(x,k,y,a+1)$ for $(x,k,y,a)\in\mathcal{G}(c)$. We prove the first part of the diagram using Theorem 5.7, and we prove the second part by applying results on homology groups for groupoids.
[…]
6.1. Part 1 of the diagram
First we define the map $\widetilde{\beta}$ of the diagram. This is a map $\widetilde{\beta}:C_{c}(X\times\mathbb{Z},\mathbb{Z})\rightarrow C_{c}(X\times\mathbb{Z},\mathbb{Z})$ given by $\widetilde{\beta}(f)(x,a)=f(x,a+1)$, for $f\in C_{c}(X\times\mathbb{Z},\mathbb{Z})$ and $(x,a)\in X\times\mathbb{Z}$. Note that $\widetilde{\beta}$ is the restriction of $\beta$ to $C_{c}(X\times\mathbb{Z},\mathbb{Z})$. Now we study the ordered group homomorphism induced by $\widetilde{\beta}$.
Lemma 6.1.
The map $[\widetilde{\beta}]_{\mathcal{G}(c)}:H_{0}(\mathcal{G}(c))\rightarrow H_{0}(\mathcal{G}(c))$ is well-defined and is an ordered group homomorphism.
Proof.
Recall from Remark 4.7 that $[\widetilde{\beta}]_{\mathcal{G}(c)}$ if given by $[\widetilde{\beta}]_{\mathcal{G}(c)}([f]_{\mathcal{G}(c)})=[\widetilde{\beta}(f)]_{\mathcal{G}(c)}$. Let $f_{1},f_{2}\in C_{c}(X\times\mathbb{Z},\mathbb{Z})$ be equivalent in $H_{0}(\mathcal{G}(c))$. Then there is an $F\in C_{c}(\mathcal{G}(c),\mathbb{Z})$ such that $f_{1}=f_{2}+\partial_{1}F$. Given $(x,a)\in X\times\mathbb{Z}$, we have
$\beta(\partial_{1}F)(x,a)$ $=\partial_{1}F(x,a+1)$
$=s_{\ast}(F)(x,a+1)-r_{\ast}(F)(x,a+1)$
$=\sum_{(g,b):s(g,b)=(x,a+1)}F(g,b)-\sum_{(g,b):r(g,b)=(x,a+1)}F(g,b)$
$=\sum_{g\in\mathcal{G}:s(g)=x}F(g,a+1-c(g))-\sum_{g\in\mathcal{G}:r(g)=x}F(g,a+1)$
$=\sum_{g\in\mathcal{G}:s(g)=x}\beta(F)(g,a-c(g))-\sum_{g\in\mathcal{G}:r(g)=x}\beta(F)(g,a)$
$=s_{\ast}(\beta(F))(x,a)-r_{\ast}(\beta(F))(x,a)$
$=\partial_{1}(\beta(F))(x,a).$
Thus, $\widetilde{\beta}(f_{1})=\widetilde{\beta}(f_{2})+\partial_{1}(\beta(F))$. Note that $\beta(F)\in C_{c}(\mathcal{G}(c),\mathbb{Z})$. Then $\beta(f_{1})$ and $\beta(f_{2})$ are equivalent and therefore $[\widetilde{\beta}]_{\mathcal{G}(c)}$ is well-defined. It straightforward from the definition that $[\widetilde{\beta}]_{\mathcal{G}(c)}$ is an ordered group homomorphism. ∎
Proposition 6.2.
The first part of the diagram (1.1) is commutative.
Proof.
Since $H_{0}(\mathcal{G}(c))$ is generated by elements of the form $[1_{V\times\{a\}}]_{\mathcal{G}(c)}$ for $V\subset X$ compact open and $a\in\mathbb{Z}$, then we only need to prove that, for all $V\subset X$ compact open and $a\in\mathbb{Z}$,
$K_{0}(\beta)\circ\mu([1_{V\times\{a\}}]_{\mathcal{G}(c)})=\mu\circ[\widetilde{\beta}]_{\mathcal{G}(c)}([1_{V\times\{a\}}]_{\mathcal{G}(c)}).$
So, fix $V$ and $a$. Then
$K_{0}(\beta)\circ\mu([1_{V\times\{a\}}]_{\mathcal{G}(c)})=K_{0}(\beta)([1_{V\times\{a\}}]_{K_{0}(B)})=[\beta(1_{V\times\{a\}})]_{K_{0}(B)}=\mu\circ[\widetilde{\beta}]_{\mathcal{G}(c)}([1_{V\times\{a\}}]_{\mathcal{G}(c)}).$
Therefore the result holds. ∎
6.2. Part 2 of the diagram
In the second part of the diagram, we have an isomorphism between homology groups of different groupoids. We will obtain this isomorphism by applying the techniques of homological similarity of groupoids. Let us define the groupoid $c^{-1}(0)$.
Until the rest of this section, we fix $\varphi:X\rightarrow X$ be a continuous section of the surjective local homeomorphism $\sigma:X\rightarrow X$ that defines $\mathcal{G}$. Let us define the map that induces the similarity of the groupoids $\mathcal{G}(c)$ and $c^{-1}(0)$.
Lemma 6.3.
There exists a homomorphism $\rho_{\varphi}:\mathcal{G}(c)\rightarrow c^{-1}(0)$ given by
$\rho_{\varphi}(x,k,y,a)=(\varphi^{a}(x),\varphi^{a+k}(y)).$
Moreover, $\rho_{\varphi}$ is a local homeomorphism.
Proof.
First we show that $\rho_{\varphi}$ is well-defined. Let $(x,k,y,a)\in\mathcal{G}(c)$, and let $m,n\in\mathbb{N}$ be such that $\sigma^{m}(x)=\sigma^{n}(y)$ and $k=m-n$. We divide the proof into four cases, considering the signs of $a$ and $a+k$. Suppose $a\geq 0$ and $a+k\geq 0$. Then $\sigma^{a}(\varphi^{a}(x))=x$ and $\sigma^{a+k}(\varphi^{a+k}(y))=y$. This implies that
$\sigma^{m+a}(\varphi^{a}(x))$ $=\sigma^{m}(\sigma^{a}(\varphi^{a}(x)))=\sigma^{m}(x)=\sigma^{n}(y),\text{\hskip 10.0ptand}$
$\sigma^{m+a}(\varphi^{a+k}(y))$ $=\sigma^{n+a+k}(\varphi^{a+k}(y))=\sigma^{n}(\sigma^{a+k}(\varphi^{a+k}(y)))=\sigma^{n}(y).$
Hence, $(\varphi^{a}(x),\varphi^{a+k}(y))\in c^{-1}(0)$. The proof for the remaining cases is analogous. Then $\rho_{\varphi}$ is well-defined.
Next we show that $\rho_{\varphi}$ is a homomorphism. Given $(x,k,y,a),(y,l,z,a+k)\in\mathcal{G}(c)$, we have
$\rho_{\varphi}((x,k,y,a)(y,l,z,a+k))$ $=\rho_{\varphi}(x,k+l,z,a)$
$=(\varphi^{a}(x),\varphi^{a+k+l}(z))$
$=(\varphi^{a}(x),\varphi^{a+k}(y))(\varphi^{a+k}(y),\varphi^{a+k+l}(z))$
$=\rho_{\varphi}(x,k,y,a)\rho_{\varphi}(y,l,z,a+k).$
Then $\rho_{\varphi}$ is a homomorphism.
Finally, we prove that $\rho_{\varphi}$ is a local homeomorphism. Fix $a\in\mathbb{Z}$. Then the restriction $\rho_{\varphi}^{(0)}|_{X\times\{a\}}:X\times\{a\}\rightarrow X$ given by $(x,a)\mapsto\varphi^{a}(x)$ is a local homeomorphism because $\varphi^{a}$ is a local homeomorphism. By Lemma 4.5, $\rho_{\varphi}$ is a local homeomorphism. ∎
Proposition 6.4.
$H_{0}(\rho_{\varphi}):H_{0}(\mathcal{G}(c))\rightarrow H_{0}(c^{-1}(0))$ is an isomorphism and has inverse $H_{0}(\eta)$, where $\eta:c^{-1}(0)\rightarrow\mathcal{G}(c)$ is defined by $\eta(x,y)=(x,0,y,0)$.
Proof.
It follows from the definition that $\eta$ is a well-defined homomorphism. Note that $\eta^{(0)}$ is defined by $\eta^{(0)}(x)=(x,0)$. Thus $\eta^{(0)}$ is a local homeomorphism. By Lemma 4.5, $\eta$ is also a local homeomorphism. It is straightforward that $\rho_{\varphi}\circ\eta=\mathrm{id}_{c^{-1}(0)}$.
Now we show that $\eta\circ\rho_{\varphi}$ is homologically similar to $\mathrm{id}_{\mathcal{G}(c)}$. Let $\theta:X\times\mathbb{Z}\rightarrow\mathcal{G}(c)$ be defined by $\theta(x,a)=(x,-a,\varphi^{a}(x),a)$. Note that $\theta$ is well-defined. In fact, if $a\geq 0$, we have that $\sigma^{a}(\varphi^{a}(x))=x$. Then $(x,-a,\varphi^{a}(x),a)\in\mathcal{G}(c)$. If $a<0$, then $\varphi^{a}(x)=\sigma^{-a}(x)$, which implies that $(x,-a,\varphi^{a}(x),a)\in\mathcal{G}(c)$. This function is also continuous because it is continuous on each coordinate.
Let $g=(x,k,y,a)\in\mathcal{G}(c)$. Then
$\theta(r(g))\eta\circ\rho_{\varphi}(g)$ $=\theta(x,a)\eta\circ\rho_{\varphi}(x,k,y,a)$
$=(x,-a,\varphi^{a}(x),a)\eta\circ\rho_{\varphi}(x,k,y,a)$
$=(x,-a,\varphi^{a}(x),a)\eta(\varphi^{a}(x),\varphi^{a+k}(y))$
$=(x,-a,\varphi^{a}(x),a)(\varphi^{a}(x),0,\varphi^{a+k}(y),0)$
$=(x,-a,\varphi^{a+k}(y),a),\text{\hskip 7.0pt and}$
$\mathrm{id}_{\mathcal{G}(c)}(g)\theta(s(g))$ $=(x,k,y,a)\theta(y,a+k)$
$=(x,k,y,a)(y,-a-k,\varphi^{a+k}(y),a+k)$
$=(x,-a,\varphi^{a+k}(y),a).$
So $\eta\circ\rho_{\varphi}$ and $\mathrm{id}_{\mathcal{G}(c)}$ are homologically similar. Therefore, $\mathcal{G}(c)$ and $c^{-1}(0)$ are homologically similar. Moreover, by Corollary 4.10, we have an isomorphism $H_{0}(\rho_{\eta}):H_{0}(\mathcal{G}(c))\rightarrow H_{0}(c^{-1}(0))$ with inverse $H_{0}(\eta)$. ∎
Proposition 6.5.
The second part of the diagram (1.1) is commutative.
Proof.
By Proposition 6.4, $H_{0}(\rho_{\varphi}):H_{0}(\mathcal{G}(c))\rightarrow H_{0}(c^{-1}(0))$ is an isomorphism with inverse $H_{0}(\eta)$. In order to show that the diagram above commutes, we prove that $(\rho_{\varphi}^{(0)})_{\ast}\circ\widetilde{\beta}\circ\eta^{(0)}_{\ast}=\sigma_{\ast}$. Let $f\in C_{c}(X,\mathbb{Z})$ and $x\in X$. Note that the restriction $\rho_{\varphi}^{(0)}$ satisfies $\rho_{\varphi}^{(0)}(y,a)=\varphi^{a}(y)$. Then
$(\rho_{\varphi})_{\ast}^{(0)}\circ\widetilde{\beta}\circ\eta_{\ast}^{(0)}(f)(x)$ $=\sum_{(y,a):\rho_{\varphi}(y,a)=x}\widetilde{\beta}\circ\eta_{\ast}^{(0)}(f)(y,a)$
$=\sum_{a\in\mathbb{Z}}\sum_{y:\varphi^{a}(y)=x}\widetilde{\beta}\circ\eta_{\ast}^{(0)}(f)(y,a)$
$=\sum_{a\in\mathbb{Z}}\sum_{y:\varphi^{a}(y)=x}\eta_{\ast}^{(0)}(f)(y,a+1).$
Note that $\eta_{\ast}^{(0)}(f)(y,a+1)=1$ if $a=-1$, and zero otherwise. Then
$(\rho_{\varphi})_{\ast}^{(0)}\circ\widetilde{\beta}\circ\eta_{\ast}^{(0)}(f)(x)$ $=\sum_{y:\varphi^{-1}(y)=x}f(y)=\sum_{y:\sigma(y)=x}f(y)=\sigma_{\ast}(f)(x).$
This implies that $[\sigma_{\ast}]=H_{0}(\rho)[\widetilde{\beta}]H_{0}(\eta)$. Therefore, the diagram commutes. ∎
7. Main Result
The main result of this paper is Theorem 1, which gives a condition on the map $\sigma:X\rightarrow X$, characterising when the corresponding groupoid C*-algebra $C^{*}(\mathcal{G})$ is AF embeddable. We prove this result by applying the commutative diagram of Section 4 in combination with Theorem 1.1.
We begin this section by defining $\sigma_{\ast}(F)$ for functions $F\in C_{c}(c^{-1}(0),\mathbb{Z})$, using the fact that $c^{-1}(0)$ is the increasing union of the sets $R(\sigma^{n})$.
Lemma 7.1.
Fix $n\in\mathbb{N}$ and let $F\in C_{c}(R(\sigma^{n}),\mathbb{Z})$. We define $\sigma_{\ast}(F)\in C_{c}(R(\sigma^{n}),\mathbb{Z})$ by
(7.1) $\sigma_{\ast}(F)(x,y)=\sum_{\begin{subarray}{c}u:\hphantom{f}\sigma(u)=x\\ v:\hphantom{f}\sigma(v)=y\end{subarray}}F(u,v).$
Then $\sigma_{*}(F)$ is well-defined. In particular, if $n\geq 1$, we have $\sigma_{\ast}(F)\in C_{c}(R(\sigma^{n-1}),\mathbb{Z})$.
Proof.
Define the sets
$\mathcal{U}_{A,B}^{n}=\{(x,y):x\in A,y\in B,\sigma^{n}(A)=\sigma^{n}(B)\},$
where $A,B$ are compact open wiht $\sigma^{n}(A)=\sigma^{n}(B)$, and $\sigma^{n}$ is injective on $A$ and $B$. Then one can show that $\sigma(1_{\mathcal{U}_{A,B}^{n,n}})=1_{\mathcal{U}_{\sigma(A),\sigma(B)}^{n-1,n-1}}$. The sets $\mathcal{U}_{A,B}^{n}$ generate the topology of $R(\sigma^{n})$, then the functions $1_{\mathcal{U}_{A,B}^{n,n}}$. By linearity, $\sigma_{\ast}(F)$ is well-defined. ∎
Theorem 1.
Let $\sigma$ be a surjective local homeomorphism on a locally compact, Hausdorff, second countable, totally disconnected space $X$. Denote by $\mathcal{G}$ the Deaconu-Renault groupoid corresponding to $\sigma$. Then the following are equivalent:
  1. (i)
    $C^{*}(\mathcal{G})$ is AF embeddable,
  2. (ii)
    $C^{*}(\mathcal{G})$ is quasidiagonal,
  3. (iii)
    $C^{*}(\mathcal{G})$ is stably finite,
  4. (iv)
    $\mathrm{Im}(\sigma_{\ast}-\mathrm{id})\cap C_{c}(X,\mathbb{N})=\{0\},$
where the map $\sigma_{\ast}:C_{c}(X,\mathbb{Z})\rightarrow C_{c}(X,\mathbb{Z})$ is defined by
$\sigma_{\ast}(f)(x)=\sum_{y:\sigma(y)=x}f(y)\hskip 10.0pt\text{for }x\in X,f\in C_{c}(X,\mathbb{Z}).$
Proof.
Let $c:\mathcal{G}\rightarrow\mathbb{Z}$ be the continuous cocycle given by $c(x,k,y)=k$ for $(x,k,y)\in\mathcal{G}$. By [20, Corollary 5.2], we have that $\mathcal{G}(c)$ is an AF groupoid, and therefore the C*-algebra $C^{*}(G(c))$ is AF. For simplicity of notation, we denote this C*-algebra by $B$.
Let $\beta$ be the action on $B$ corresponding to the dual action $\hat{\alpha}^{c}$ and induced by the isomorphism $B\cong C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}$ from Proposition 3.4. Then Brown’s theorem (Theorem 1.1) implies the equivalent conditions bellow.
  1. (i’)
    $B\rtimes_{\beta}\mathbb{Z}$ is AF embeddable,
  2. (ii’)
    $B\rtimes_{\beta}\mathbb{Z}$ is quasidiagonal,
  3. (iii’)
    $B\rtimes_{\beta}\mathbb{Z}$ is stably finite,
  4. (iv’)
    $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$.
By Proposition 3.4, $B\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(\mathcal{G})\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic. It follows from Takai duality that $C^{*}(\mathcal{G})\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ and $C^{*}(\mathcal{G})$ are stably isomorphic. So, $B\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(\mathcal{G})$ are stably isomorphic. Since AF embeddability, quasidiagonality and stable finiteness are preserved under stable isomorphisms, then we have the equivalence below:
  1. (i)
    $C^{*}(\mathcal{G})$ is AF embeddable,
  2. (ii)
    $C^{*}(\mathcal{G})$ is quasidiagonal,
  3. (iii)
    $C^{*}(\mathcal{G})$ is stably finite,
  4. (iv’)
    $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$.
Now we study condition (iv’). By the commutative diagram (1.1), (iv’) is false if, and only if, there are non-zero $f\in C_{c}(X,\mathbb{Z})$ and $h\in C_{c}(X,\mathbb{N})$ such that
$[\sigma_{\ast}(f)-f]=[h].$
Suppose the equation above holds. Then there exists $F\in C_{c}(c^{-1}(0),\mathbb{Z})$ such that $\sigma_{\ast}(f)-f=h+\partial_{1}F$. Then there exists $n\geq 1$ such that $F\in C_{c}(R(\sigma^{n}),\mathbb{Z})$. Define $\partial_{1}^{(n)}:C_{C}(R(\sigma^{n}),\mathbb{Z})\rightarrow C_{C}(R(\sigma^{n}),\mathbb{Z})$ by $\partial_{1}^{(n)}=(s|_{R(\sigma^{n})})_{\ast}-(r|_{R(\sigma^{n})})_{\ast}$. It is straightforward that $\partial_{1}F=\partial_{1}^{(n)}F$. Hence,
(7.2) $\sigma_{\ast}(f)-f=h+\partial_{1}^{(n)}F.$
Moreover,
$\sigma_{\ast}(\partial^{(n)}_{1}(F))(x)$ $=\sum_{\begin{subarray}{c}y:\sigma(y)=x\\ z:\sigma^{n}(z)=\sigma^{n}(y)\end{subarray}}[F(z,y)-F(y,z)]$
$=\sum_{\begin{subarray}{c}v:\sigma^{n-1}(v)=\sigma^{n-1}(x)\\ y:\sigma(y)=x\\ z:\sigma(z)=v\end{subarray}}[F(z,y)-F(y,z)]$
$=\sum_{v:\sigma^{n-1}(v)=\sigma^{n-1}(x)}[\sigma_{\ast}(F)(v,x)-\sigma_{\ast}(F)(x,v)]$
$=\partial_{1}^{(n-1)}(\sigma_{\ast}(F))(x)$
It follows from Lemma 7.1 that $\sigma_{\ast}(F)\in C_{c}(R(\sigma^{n}))$. By applying $\sigma^{n}_{\ast}$ on both sides of the equation (7.2) and setting $\widetilde{f}=\sigma_{\ast}^{n}(f)$ and $\widehat{h}=\sigma_{\ast}^{n}(h)$, we obtain
$\sigma_{\ast}(\widetilde{f})-\widetilde{f}=\widetilde{h}+\sigma_{\ast}^{n}(\partial_{1}^{(n)}F)=\widetilde{h}.$
We have just proved that $\mathrm{Im}(\sigma_{\ast}-\mathrm{id})\cap C_{c}(X,\mathbb{N})\neq\{0\}$. Now we prove the converse. Suppose there are $f\in C_{c}(X,\mathbb{N})$ and $h\in C_{c}(X,\mathbb{N})$ such that $h\neq 0$ and $\sigma_{\ast}(f)-f=h$. It follows from Lemma 4.11 that $[h]\neq 0$. Then equation (iv’) holds, which implies that $C^{*}(\mathcal{G})$ is not AF embeddable. ∎
We will use the following corollary to show that our main theorem generalises a known theorem for graph algebras.
Corollary 7.2.
The following are equivalent:
  1. (1)
    $C^{*}(\mathcal{G})$ is AF embeddable,
  2. (2)
    $\mathrm{Im}(\sigma_{\ast}-\mathrm{id})\cap C_{c}(X,\mathbb{N})=\{0\},$
  3. (3)
    $\mathrm{Im}(\sigma_{\ast}^{n}-\mathrm{id})\cap C_{c}(X,\mathbb{N})=\{0\}$ for some $n\geq 1$,
  4. (4)
    $\mathrm{Im}(\sigma_{\ast}-\mathrm{id})\cap C_{c}(X,\mathbb{N})=\{0\}$ for all $n\geq 1$.
Proof.
The equivalence $(1)\Leftrightarrow(2)$ follows from Theorem 1, and the implications $(4)\Rightarrow(2)\Rightarrow(3)$ are straightforward. We prove $(3)\Rightarrow(4)$ by contrapositive. Suppose $(4)$ is false, and fix $n$. Then, for some $m\geq 1$, there are nonzero $f\in C_{c}(X,\mathbb{Z}$ and $h\in C_{c}(X,\mathbb{N})$ such that $\sigma_{\ast}(f)-f=h$. Then, for all $k$, we have
$\sigma^{k}_{\ast}(\sigma_{\ast}^{m}(f))-\sigma_{\ast}^{k}(f)=\sigma_{\ast}^{k}(h).$
Let $F=\sigma_{\ast}^{m}(f)$ and $H=h+\sigma_{\ast}(h)+\dots+\sigma_{\ast}^{n-1}(h)$. Note that $H$ is non-negative and nonzero. By a telescoping sum, we have
$\sigma^{n}_{\ast}(F)-F=\sum_{k=0}^{n-1}[\sigma_{\ast}^{k+1}(\sigma_{\ast}^{m}(f))-\sigma_{\ast}^{k}(\sigma_{\ast}^{m}(f))]=\sum_{k=0}^{n-1}\sigma_{\ast}^{k}(h)=H.$
Thus $\mathrm{Im}(\sigma_{\ast}^{n}-\mathrm{id})\cap C_{c}(X,\mathbb{N})\neq\{0\}$. Since $n$ is arbitrary, then $(3)$ is false. Therefore, $(3)\Rightarrow(4)$. ∎
8. Examples
In this section we show that Theorem 1 generalises three known results in the literature that characterise the AF embeddability of graph algebras, topological graph algebras and the crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$.
8.1. Graph algebras
Here we prove Schafhauser’s result on graph algebras [54] as a corollary of Theorem 1.
A (directed) graph is a tuple $E=(E^{0},E^{1},r,s)$ where $E^{0}$ and $E^{1}$ are countable sets equipped with the range and source maps $r,s:E^{1}\rightarrow E^{0}$. The elements of $E^{0}$ are called vertices, while $E^{1}$ contains the edges of the graph. We say that $E$ is row-finite if $r^{-1}(v)$ is finite for every vertex $v$. Given a vertex $w$, we say that $w$ is a source if $r^{-1}(w)=\emptyset$, and call $w$ a sink if $s^{-1}(w)=\emptyset$. A finite or infinite sequence of edges $\mu_{1}\mu_{2}\dots$ is a path if, for every two consecutive edges, we have $s(\mu_{i})=r(\mu_{i+1})$. $E^{\ast}$, $E^{\infty}$ denote the sets of finite and infinite paths, respectively, while $E^{n}$ is the set of paths of length $n$. The range of a path is the range of its first edge, while the source of a finite path is the source of its last element. We denote the length of a path $\mu$ by $|\mu|$. A vertex $v$ is considered path of length zero and $r(v)=s(v)=v$. Given two paths $\mu,\nu$ such that $s(\mu)=r(\nu)$, $\mu\nu$ is the path obtained by concatenating the two sequences. A finite path $\mu$ is a cycle if $r(\mu)=s(\mu)$. This cycle has an entrance if there exists an edge $e$ with $r(e)=r(\mu_{i})$ and $e\neq\mu_{i}$ for some $i$. For an introduction to graph algebras, see the book [46] by Raeburn.
Given a row-finite graph with no sources, [33, Theorem 4.2] gives an isomorphism from the graph algebra $C^{*}(E)$ to the C*-algebra $C^{*}(G_{E})$ defined by $1_{e}\mapsto 1_{Z(e,r(e))}$, where
$G_{E}=\{(x_{0}z,|x_{0}|-|y_{0}|,y_{0}z):x_{0},y_{0}\in E^{\ast},z\in E^{\infty}\},$
and $Z(e,s(e))=\{(ex,1,x):x\in E^{\infty},r(x)=s(e)\}$. The set $G_{E}$ is an example of a Deaconu-Renault groupoid, where the map $\sigma:E^{\infty}\rightarrow E^{\infty}$ is given by
$\sigma(x_{1}x_{2}x_{3}\dots)=x_{2}x_{3}\dots$
and the topology on $E^{\infty}$ is generated by the cylinders $Z(\mu)$ for $\mu\in E^{*}$, where each $Z(\mu)$ is the set of all infinite sequences starting with $\mu$.
The groupoid is almost identical to the groupoid in [33, Definition 2.3], but we made some changes to make it compatible with the notation in Raeburn’s book [46]. In constrast to our definition, in [33] two consecutive edges $x_{i},x_{i+1}$ satisfy $r(x_{i})=s(x_{i+1})$. Also, their groupoid has elements of the form $(x,-k,y)$ instead of $(x,k,y)$.
We want to apply the isomorphism $C^{*}(E)\cong C^{*}(G_{E})$, so we fix $E$ to be a row-finite graph with no sources. Note also that the map $\sigma$ is surjective if, and only if, $E$ has no sinks. In order to apply Theorem 1, we need $\sigma$ to be surjective, so we also assume throughout this subsection that $E$ has no sinks.
Schafhauser proved in [54] that the graph algebra $C^{*}(E)$ is AF embeddable if, and only if, no cycle in $E$ has an entrance. In the next lemmas we study properties of $E^{\infty}$ that will help us to understand when Schafhauser’s condition is false.
Lemma 8.1.
Suppose that no cycle in $E$ has an entrance. Given $x\in E^{\infty}$, then either
  1. (1)
    all edges of $x$ are distinct, or
  2. (2)
    there are unique $n\geq 0,p\geq 1$ such that $\sigma^{n+p}(x)=\sigma^{n}(x)$ and the elements $x_{1},\dots,x_{n+p}$ are distinct.
Proof.
Let $x\in E^{\infty}$. If all edges of $x$ are distinct, it is straightforward that $(2)$ is false. Now suppose that $(1)$ is false. This means that $x$ has repeated edges. Let $n$ be the smallest number such that $x_{n+1}$ appears more than once in $x$. Let $p$ be the smallest number such that $p\geq 1$ and $x_{n+p+1}=x_{n+1}$. Then $x_{1},\dots,x_{n+p}$ are distinct. Let $\mu=x_{1}\dots x_{n}$ and $\nu=x_{n+1}\dots x_{n+p}$. Note that $\nu$ is a cycle. By hypothesis, $\nu$ has no entrance. Then $\nu\nu\nu\dots$ is the only infinite path in $E$ with range $r(\nu)$. This implies that $x=\mu\nu\nu\dots$ and therefore $\sigma^{n+p}(x)=\sigma^{n}(x)$. ∎
Lemma 8.2.
Suppose that no cycle in $E$ has an entrance. Let $x\in E^{\infty}$ be such that condition $(2)$ of Lemma 8.1 holds for $n,p\geq 1$ (note that $n\geq 0$ is assumed). Given $y\in E^{\infty}$ with $\sigma(y)=x$, then it satisfies condition $(2)$ of that lemma for $n+1$ and $p$.
Proof.
As in the proof of Lemma 8.1, we can write $x$ as $x=\mu\nu\nu\nu\dots$ where the finite paths $\mu$ and $\nu$ have lengths $n$ and $p$, respectively, and the edges of $\mu\nu$ are distinct. By hypothesis, we can write $y=e\mu\nu\nu\dots$, where $e\in E^{1}$. If $e$ is an edge in $\mu$, then $\mu$ contains a cycle with an entrance, which is a contradiction. If $e$ is in $\nu$, then $\mu_{1}$ is an entrance of $\nu$, which is a contradiction. Therefore $y$ satisfies condition $(2)$ for $n+1$ and $p$. ∎
Using similar arguments, we can prove the lemma below.
Lemma 8.3.
Suppose that no cycle in $E$ has an entrance. Let $x\in E^{\infty}$ be such that condition $(2)$ of Lemma 8.1 holds for $n=0,p\geq 1$. If $\sigma^{p}(y)=x$ and $y$ satisfies condition $(2)$ of the same lemma for $n=0$, then $y=x$.
The following lemmas study functions in $C_{c}(E^{\infty},\mathbb{Z})$.
Lemma 8.4.
Suppose that no cycle in $E$ has an entrance, and let $x\in E^{\infty}$ have distinct edges. Given $F\in C_{c}(E^{\infty},\mathbb{Z})$, there exists $n\geq 1$ such that $F(\sigma^{k}(x))=0$ for all $k\geq n$.
Proof.
Let $x=x_{1}x_{2}x_{3}\dots$. Then the vertices $r(x_{1}),r(x_{2}),\dots$ are distinct, otherwise $E$ would have a cycle with an entrance. Moreover, $\sigma^{k}(x)\in Z(r(x_{k}))$ for $k\geq 0$. Note that $\{Z(v)\}_{v\in E^{0}}$ is a cover of $\mathrm{supp\hphantom{.}}F$ by open sets. By compactness, the set $\mathrm{supp\hphantom{.}}F$ has a finite subcover $\{Z(v)\}_{v\in I}$. Therefore, there are finitely many elements $k$ such that $F(\sigma^{k}(x))\neq 0$. ∎
By a similar argument, we can prove the following:
Lemma 8.5.
Suppose that no cycle in $E$ has an entrance. Let $\{x^{(k)}\}_{k\geq 1}$ be a sequence in $E^{\infty}$ such that
  • $\sigma(x^{(k+1)})=x^{(k)}$,
  • $r(x^{(1)}),r(x^{(2)}),\dots$ are distinct.
Then there exists an $n\geq 1$ such that $F(x^{(k)})=0$ for all $k\geq n$.
Lemma 8.6.
Let $\mu\in E^{*}$. Then
$\sigma_{\ast}^{|\mu|+1}(1_{Z(\mu)})=\sum_{\begin{subarray}{c}e\in E^{1}\\ r(e)=s(\mu)\end{subarray}}1_{Z(s(e))}.$
Proof.
Let $x\in E^{\infty}$. Then
$\sigma_{\ast}^{|\mu|+1}(1_{Z(\mu)})(x)=\sum_{y:\sigma^{|\mu|+1}(y)=x}1_{Z(\mu)}(y).$
If $11_{Z(\mu)}(y)=$ 1, we must have $y=\mu ex$ for some $e\in E^{1}$. Then
$\sigma_{\ast}^{|\mu|+1}(1_{Z(\mu)}(x)=\sum_{\begin{subarray}{c}e\in E^{1}\\ r(e)=s(\mu)\\ s(e)=r(x)\end{subarray}}1=\sum_{\begin{subarray}{c}e\in E^{1}\\ r(e)=s(\mu)\end{subarray}}1_{Z(s(e))}(x).$
Therefore the result holds. ∎
Now we show Schafhauser’s theorem [54] as an application of Theorem 1 with the further assumptions that the graph is row-finite with no sinks or sources.
Corollary 8.7.
The following are equivalent:
  1. (i)
    $C^{*}(E)$ is AF embeddable,
  2. (ii)
    $C^{*}(E)$ is AF quasidiagonal,
  3. (iii)
    $C^{*}(E)$ is AF stably finite,
  4. (iv)
    $C^{*}(E)$ is finite,
  5. (v)
    No cycle in $E$ has an entrance.
Proof.
Note that $C^{*}(E)\cong C^{*}(G_{E})$. We are going to prove that $C^{*}(E)$ is not AF embeddable if, and only if, $E$ has a cycle with an entrance. First we prove the implication ()($\Rightarrow$)( ⇒ ). So, suppose that $C^{*}(E)$ is not AF embeddable. By Theorem 1, there are two non-zero functions $f\in C_{c}(E^{\infty},\mathbb{Z})$, $h\in C_{c}(E^{\infty},\mathbb{N})$ such that $\sigma_{\ast}(f)-f=h$.
Assume that no cycle in $E$ has an entrance. With this assumption, we will find a contradiction by showing that $h=0$. In order to do this, we will consider the different types of infinite paths $x$ given by Lemma 8.1:
  1. (a)
    If $x$ has distinct edges, then $h(x)=0$
    Let $x\in E^{\infty}$ have distinct edges. Suppose that $f(x)>0$. Since $\sigma_{\ast}(f)(x)\geq f(x)$, there exists $x^{(2)}\in E^{\infty}$ with distinct edges such that $\sigma(x^{(2)})=x$ and $f(x^{(2)})>0$. By repeating this argument, we find a sequence of infinite paths $x^{(k)}$ such that $x^{(1)}=x$ and, for all $k\geq 1$, $\sigma(x^{(k+1)})=x^{(k)}$ and $f(x^{(k)})>0$. Since $x$ has distinct edges, Lemma 8.5 implies that $f(x^{(k)})=0$ for some $k$, which is a contradiction. Therefore, we must have $f(x)\leq 0$. In particular, Lemma 8.1 implies that $y$ has distinct edges for all $y$ with $\sigma(y)=x$. Then $f(y)\leq 0$.
    Now suppose that $f(x)<0$. Then
    $f(\sigma(x))\leq\sigma_{\ast}(f)(\sigma(x))=\sum_{y:\sigma(y)=\sigma(x)}f(y)\leq f(x),$
    because $\sigma(x)$ has distinct edges and thus $f(y)\leq 0$ for all $y$ with $\sigma(y)=\sigma(x)$. By repeating this argument, we have that, for all $k\geq 1$, $f(\sigma^{k}(x))\leq f(x)<0$. This contradicts Lemma 8.4. Therefore, $f(x)=0$ for all $x\in E^{\infty}$ with distinct edges. Since every $y$ with $\sigma(y)=x$ also has distinct edges, it follows that $\sigma_{\ast}(f)(x)=0$. Therefore $h(x)=0$.
  2. (b)
    If $x$ satisfies condition $(2)$ of Lemma 8.1 for $n,p\geq 1$, then $f(x)\leq 0$
    Assume that $f(x)>0$. By similar arguments to the previous item, we can find a sequence $x^{(k)}$ such that $x^{(1)}=x$, $\sigma(x^{(k+1)})=x^{(k)}$, and $f(x^{(k)})>0$. By Lemma 8.5 all $r(x^{(k)})$ are distinct. This implies that $\{Z(r(x^{(k)}))\}_{k\geq 1}$ is an infinite open cover of $\mathrm{supp}\hphantom{.}f$ by disjoint subsets, which is a contradiction. Therefore $f(x)\leq 0$.
  3. (c)
    If $x$ satisfies condition $(2)$ of of Lemma 8.1 for $n=0,p\geq 1$, then $h(x)=0$ and $\sigma_{\ast}^{p}(f)(x)=f(x)$
    Note that by assumption, $\sigma^{p}(x)=x$. By Lemma 8.3,
    $\sigma_{\ast}^{p}(f)(x)=\sum_{y:\sigma^{p}(y)=x}f(y)=\sum_{\begin{subarray}{c}y:\sigma^{p}(y)=x\\ y\neq x\end{subarray}}f(y)+f(x)\leq f(x).$
    Using the same telescoping sums from the proof of Corollary 7.2, we have
    $f(x)\leq\sigma_{\ast}(f)(x)\leq\sigma_{\ast}^{2}(f)(x)\leq\sigma_{\ast}^{p}(f)(x)\leq f(x).$
    This implies that $h(x)=\sigma_{\ast}(f)(x)-f(x)=0$.
  4. (d)
    If $x$ satisfies condition $(2)$ of Lemma 8.1 for $n=1$, $p\geq 1$, then $f(x)=0$
    Let $z_{0}=\sigma^{n}(x)$ and, for every $k\geq 1$, let $z_{k}=\sigma^{kn(2p-1)}(z_{0})$. Note that $\sigma^{p}(z_{0})=z_{0}$. Moreover, $z_{0}$ satisfies condition $(2)$ of Lemma 8.1 for 00 and $p$. Note that $\sigma^{n}(z_{k+1})=z_{k}$. Indeed,
    $\sigma^{n}(z_{k+1})=\sigma^{n}(\sigma^{(k+1)n(2p-1)}(z_{0}))=\sigma^{kn(2p-1)}(\sigma^{2pn}(z_{0}))=\sigma^{kn(2p-1)}(z_{0})=z_{k}.$
    Note that each $z_{k}$ satisfies condition $(2)$ of Lemma 8.1 for 00 and $p$. So, we have
    (8.1) $f(z_{k})\leq\sigma_{\ast}^{n}(f)(z_{k})=\sum_{y:\sigma^{n}(y)=z_{k}}f(y)=f(z_{k+1})+\sum_{\begin{subarray}{c}\sigma^{n}(y)=z_{k}\\ y\neq z_{k+1}\end{subarray}}f(y).$
    But if $y\neq z_{k+1}$ and $\sigma^{n}(y)=z_{k}$, then $y$ satisfies condition $(2)$ of Lemma 8.1 for $n,p\geq 1$. So $f(y)\leq 0$ by item (b) of this proof. Hence, $f(z_{k})\leq f(z_{k+1})$. It is straightforward from the definition that $z_{p}=z_{0}$. This implies that $f(z_{0})=f(z_{1})=\dots=f(z_{p})$. By (8.1), $f(y)=0$ for all $y\neq z_{1}$ such that $\sigma^{n}(y)=z_{0}=\sigma^{n}(x)$. Since $z_{1}$ satisfies condition $(2)$ of 8.1 for 00 and $p$, we have that $x\neq z_{1}$. Therefore $f(x)=0$.
  5. (e)
    For every $x\in E^{\infty}$ with repeated edges, we have $h(x)=0$
    Note that $x$ satisfies condition $(2)$ of Lemma 8.1 for $n\geq 0$ and $p\geq 1$. Also, $f(x)\leq 0$ by (b). Since we covered the case $n=0$ in item $(c)$, we assume here that $n\geq 1$. We claim that $\sigma_{\ast}^{k}(f)(x)=0$. If this were false, we could find a sequence $x^{(k)}$ such that $x^{(1)}=x$, $\sigma(x^{(k+1)})=x^{(k)}$ and $f(x^{(k)})\neq 0$, which would contradict the fact that $f$ is compactly supported. Hence,
    $0=\sigma_{\ast}^{k}(f)(x)\geq\sigma_{\ast}^{k-1}(f)(x)\geq\dots\sigma_{\ast}(f)(x)\geq f(x)\leq 0$
    Therefore, $h(x)=\sigma_{\ast}(f)(x)-f(x)=0$.
Now let $x\in E^{\infty}$. In item (a), we showed that $h(x)=0$ if $x$ has distinct edges; in item (e), we proved that $h(x)=0$ if $x$ has repeated edges. This implies that $h=0$, which is a contradiction. Therefore, $E$ has a cycle with an entrance.
Finally, we show ()($\Leftarrow$)( ⇐ ). Suppose $E$ has a cycle with an entrance. Then there is a finite path $\mu\in E^{*}$ with $s(\mu)=r(\mu)$, and there exists an $e_{1}\in E^{1}$ such that $r(e_{1})=s(\mu)$ and $e_{1}\neq\mu_{1}$. Since $E$ has no sinks, all cycles in $E$ must have length greater than zero.
Note that $0\leq\sigma_{\ast}(1_{Z(\mu)})\leq 1_{Z(s(\mu_{1}))}$. In fact, $\sigma_{\ast}(1_{Z(\mu)})\geq 0$ by definition of $\sigma_{\ast}$. Let $x\in E^{\infty}$. Then
$\sigma_{\ast}(1_{Z(\mu)})(x)=\sum_{y:\sigma(y)=x}1_{Z(\mu)}(y)=1_{Z(\mu)}(\mu_{1}x)1_{Z(s(\mu_{1}))}(x)\leq 1_{Z(s(\mu))}(x).$
By Lemma 8.6, we have
$\sigma_{\ast}^{|\mu|+1}(1_{Z(\mu)})=\sum_{\begin{subarray}{c}e\in E^{1}\\ r(e)=s(\mu)\end{subarray}}1_{Z(s(e))}\geq 1_{Z(s(\mu_{1}))}+1_{Z(s(e_{1}))}\geq\sigma_{\ast}(1_{Z(\mu)})+1_{Z(s(e_{1}))}.$
Then there exists a non-zero function $h\in C_{c}(E^{\infty},\mathbb{N})$ such that
$\sigma_{\ast}^{|\mu|}(\sigma_{\ast}(1_{Z(\mu)}))-\sigma_{\ast}(1_{Z(\mu)})=h.$
By Corollary 7.2, this implies that $C^{*}(E)\cong C^{*}(G_{E})$ is not AF embeddable.
So far we have proved $(i)\Leftrightarrow(v)$. By Theorem 1, we have the equivalence $(i)\Leftrightarrow(ii)\Leftrightarrow(iii)\Leftrightarrow(v)$. The implication $(iii)\Rightarrow(iv)$ is straightforward, and $(iv)\Rightarrow(v)$ follows from the same argument used in [54, Theorem 1.1]. ∎
8.2. The crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$
Now we apply Theorem 1 to show that it generalises Pimsner’s theorem [45, Theorem 9] (see also [9, Theorem 11.5] by Brown). In contrast to these references, we do not assume that $X$ is a compact metric space. Instead, we fix $X$ to be a locally compact, Hausdorff, second countable, and totally disconnected space.
Note that, when $\sigma:X\rightarrow X$ is a homeomorphism, it follows from Deaconu’s paper [14] that $C(X)\rtimes_{\sigma}\mathbb{Z}$ is isomorphic to $C^{*}(\mathcal{G})$.
Definition 8.8.
We say that $\sigma$ compresses an open set $\mathcal{U}\subset X$ if $\sigma(\mathcal{U})\subsetneq\mathcal{U}$. So, the homeomorphism $\sigma$ compresses no compact open sets if the following condition holds for all $\mathcal{U}\subset X$ compact open: $\sigma(\mathcal{U})\subset\mathcal{U}\Rightarrow\sigma(\mathcal{U})=\mathcal{U}.$
The definition above is stated differently in [9]. Recall that our topological space is more specific, being totally disconnected. For our purposes, both definitions are equivalent when $X$ is assumed to be compact. Now we show that Theorem 1 can be applied to prove [45, Theorem 9] (or [9, Theorem 11.5]).
Corollary 8.9.
Let $X$ be a locally compact, Hausdorff, second countable, totally disconnected space. Then the following are equivalent:
  1. (i)
    $C_{0}(X)\rtimes_{\sigma}\mathbb{Z}$ is AF embeddable,
  2. (ii)
    $C_{0}(X)\rtimes_{\sigma}\mathbb{Z}$ is quasidiagonal,
  3. (iii)
    $C_{0}(X)\rtimes_{\sigma}\mathbb{Z}$ is stably finite,
  4. (iv)
    $\sigma$ compresses no compact open sets.
Proof.
Conditions (i)-(iii) are equivalent by Theorem 1. We show the equivalence $(iv)\Leftrightarrow(i)$. Suppose $C_{0}(X)\rtimes_{\sigma}\mathbb{Z}$ is not AF embeddable. By Theorem 1, there are non-zero functions $f\in C_{c}(X,\mathbb{Z})$ and $h\in C_{c}(X,\mathbb{N})$ such that
(8.2) $\sigma_{\ast}(f)-f=h.$
Since $\sigma$ is a homeomorphism, we have $\sigma_{\ast}(f)=f\circ\sigma^{-1}$. By applying $\sigma$ to both sides of (8.2), we have
(8.3) $f-f\circ\sigma=h\circ\sigma.$
Note that $h\circ\sigma\in C_{c}(X,\mathbb{N})$ is non-zero, and $f\geq f\circ\sigma$.
Given a natural number $n$, define the function $H_{n}=\sum_{i=0}^{n}h\circ\sigma^{i}$. Then $H_{n}$ assumes non-negative values. Similarly, define $H(x)$ for every $x\in X$ by
$H(x)=\sum_{k=0}^{\infty}h\circ\sigma^{k}(x).$
It follows from (8.3) that $H(x)$ is bounded. The dominated convergence theorem implies that the equation above defines a continuous function $H:X\rightarrow\mathbb{N}$. Also, $H$ is non-zero.
We are going to find a set that is a compressed by $\sigma$. Since $H$ is non-zero and bounded, there exists an $x_{0}\in X$ be such that $H(x_{0})=h(x_{0})\geq 1$. We claim that $f(x_{0})\leq 0$. Suppose this is false. Then define the compact sets
$A=\{x\in X:f(x)\geq 1\hskip 7.0pt\text{and}\hskip 7.0ptH(x)\geq 1\}\hskip 5.0pt\text{and}\hskip 5.0ptK_{n}=\{x\in X:f(x)\geq 1\hskip 7.0pt\text{and}\hskip 7.0ptH_{n-1}(x)=0\}.$
Both sets are non-empty because $\sigma^{-n}(x_{0})\in K_{n}\cap A$ for $n\geq 1$. By the finite intersection property, there exists an $x_{\infty}\in\cap_{n=1}^{\infty}(A\cap K_{n})$, which implies that $H(x_{\infty})=0$ because $x_{\infty}$ is in the intersection of the sets $K_{n}$. But we also have $H(x_{\infty})\geq 1$ because $x_{\infty}\in A$, which is a contradiction. Therefore, $f(x_{\infty})\leq 0$.
Define the sets
$\mathcal{U}$ =$=\{x:H(x)=h(x)=h(x_{0})>0\hskip 7.0pt\text{and}\hskip 7.0ptf(x)=f(x_{0})\},\hskip 7.0pt\text{and}$
$V$ =$=\{x:H(x)=0\hskip 7.0pt\text{and}\hskip 7.0ptf(x)<f(x_{0})\}.$
Then $\sigma(\mathcal{U})\cup\sigma(V)\subset V$, which implies that $\sigma$ compresses the compact set $\mathcal{U}\cup V$.
Conversely, suppose that $\sigma$ compresses a compact open subset $\mathcal{U}\subset X$. Let $f=-1_{\mathcal{U}}$. Then, by the definition of $\sigma_{\ast}$,
$\sigma_{\ast}(f)-f=-1_{\mathcal{U}}\circ\sigma^{-1}+1_{\mathcal{U}}=1_{\mathcal{U}}-1_{\sigma(\mathcal{U})}.$
Since $\mathcal{U}\subsetneq\sigma(\mathcal{U})$, the function $h=1_{\mathcal{U}\setminus\sigma(\mathcal{U})}\in C_{c}(X,\mathbb{N})$ is nonzero and satisfies $\sigma_{\ast}(f)-f=h$. Therefore, $C_{0}(X)\rtimes_{\sigma}\mathbb{Z}$ is not AF embeddable. ∎
8.3. Topological graph algebras
In this subsection we prove that Theorem 1 is an application of Theorem 6.7 of [52] by Schafhauser. We consider the definition of topological graphs introduced by Katsura in [26] and studied in the subsequent papers [27, 28, 29].
Definition 8.10.
A topological graph $E=(E^{0},E^{1},s,r)$ consists of two locally compact Hausdorff spaces $E^{0},E^{1}$, and two maps $r,s:E^{1}\rightarrow E^{0}$, where $s$ is a local homeomorphism and $r$ is continuous. A $v\in E^{0}$ is called a sink if $s^{-1}(v)=\emptyset$. We say that $E$ is compact if $E^{0},E^{1}$ are compact and $s$ is surjective.
For consistence of notation, we make a change of notation, using $s$ instead of $d$. Here we also assume that the spaces $E^{0}$ and $E^{1}$ are Hausdorff. Note that every directed graph is an example of topological graph. Now define $\varepsilon$-pseudopaths as in Schafhauser’s paper.
Definition 8.11.
Let $E$ be a topological graph, and let $\varepsilon>0$. Let $d$ be a metric on $E^{0}$ compatible with the topology. An $\varepsilon$-pseudopath in $E$ is a finite sequence $\alpha=(e_{1},e_{2},\dots,e_{n})\in E^{1}$ such that for each $i=1,\dots,n-1$, $d(s(e_{i}),r(e_{i+1}))<\varepsilon$. We write $r(\alpha)=r(e_{1})$ and $s(\alpha)=s(e_{n})$. An $\varepsilon$-pseudopath $\alpha$ is called an $\varepsilon$-pseudoloop based at $r(\alpha)$ if $d(r(\alpha),s(\alpha))<\varepsilon$.
Here we make a slight change of notation by writing $\alpha$ as $(e_{1},\dots,e_{n})$ instead of $(e_{n},\dots,e_{1})$ from [52], to make our notation compatible with the notation of graphs. We also write that a $\varepsilon$-pseudoloop $\alpha$ is based at $r(\alpha)$ insted of $s(\alpha)$. Now we state the theorem that characterises the AF embeddability of topological graph algebras, writing $r$ instead $s$ in item (v), to agree with our notation.
Theorem 8.12.
[52, Theorem 6.7] if $E$ is a compact topological graph with no sinks, then the following are equivalent:
  1. (i)
    $C^{*}(E)$ is AF embeddable,
  2. (ii)
    $C^{*}(E)$ is quasidiagonal,
  3. (iii)
    $C^{*}(E)$ is stably finite,
  4. (iv)
    $C^{*}(E)$ is finite,
  5. (v)
    $r$ is injective and for every $v\in E^{0}$ and $\varepsilon>0$, there exists an $\varepsilon$-pseudoloop in $E$ based at $v$.
Note that, for a Deaconu-Renault groupoid $\mathcal{G}$ corresponding to $\sigma:X\rightarrow X$, then $E=(E^{0},E^{1},s,r)\coloneqq(X,X,\sigma,\mathrm{id})$ is a topological graph with $C^{*}(E)\cong C^{*}(\mathcal{G})$ by [27, Proposition 10.9]. For Deaconu-Renault groupoids, we prove the equivalence of items $(i),(ii),(iii)$ and $(v)$ in Theorem 8.12.
Corollary 8.13.
Let $E=(E^{0},E^{1},s,r)\coloneqq(X,X,\sigma,\mathrm{id})$ be the topological graph corresponding to $\mathcal{G}$. Then the following are equivalent:
  1. (i)
    $C^{*}(\mathcal{G})$ is AF embeddable,
  2. (ii)
    $C^{*}(\mathcal{G})$ is quasidiagonal,
  3. (iii)
    $C^{*}(\mathcal{G})$ is stably finite,
  4. (iv)
    $\mathrm{Im}(\sigma_{\ast}-\mathrm{id})\cap C_{c}(X,\mathbb{N})=\{0\}.$
Moreover, consider the following statement:
  1. (v)
    for every $x\in X$ and $\varepsilon>0$, there is an $\varepsilon$-pseudoloop in $E$ based at $x$.
Then $(v)\Rightarrow(iv)$. If $X$ is compact, then all the items above are equivalent.
Proof.
Theorem 1 gives the equivalence $(i)\text{--}(iv)$. First we show $(v)\Rightarrow(iv)$. Suppose $(iv)$ is false. Then there are nonzero functions $f\in C_{c}(X,\mathbb{Z})$ and $h\in C_{c}(X,\mathbb{N})$ such that $\sigma_{\ast}(f)-f=h$. Since $\sigma$ is a local homeomorphism on a totally disconnected space and since $f,\sigma_{\ast}(f)\in C_{c}(X,\mathbb{Z})$, then there are disjoint compact open sets $V_{1},\dots,V_{n}$, there is a number $1\leq m\leq n$, and a surjective function $\rho:\{1,\dots,n\}\rightarrow\{1,\dots,m\}$ such that
  • $\sigma$ is injective on each $V_{i}$,
  • $\sigma(V_{i})=V_{\rho(i)}$,
  • There are integers $a_{1},\dots,a_{n}$ and nonzero integers $b_{1},\dots,b_{m}$ such that
    $\sigma_{\ast}(f)=\sum_{i=1}^{n}a_{i}1_{V_{i}}\hskip 7.0pt\text{and}\hskip 7.0ptf=\sum_{j=1}^{m}b_{j}1_{V_{j}}.$
It is straightforward that, for every $j=1,\dots,m$, $b_{j}=\sum_{i:\rho(i)=j}a_{i}$. We claim that $n>m$. Suppos this is false. Then $\rho$ is a bijection with $a_{j}=b_{\rho^{-1}(j)}$. Since $h$ is nonzero, then there exists an $l$ such that $a_{j}>b_{j}$. This implies that
$\sum_{i=1}^{n}a_{i}>\sum_{i=1}^{n}b_{i}\Rightarrow\sum_{i=1}^{n}b_{\rho^{-1}(i)}=\sum_{i=1}^{n}b_{i}>\sum_{i=1}^{n}b_{i},$
which is a contradiction. Therefore, $n>m$.
[…]
Now assume that $X$ is compact. We prove $(iv)\Rightarrow(v)$. Suppose $(v)$ is false. Then there are $x\in X$ and $\varepsilon>0$ such that there exists no $\varepsilon$-pseudoloop in $E$ based at $x$. By continuity, there exists an open set $V\subset B(x,\varepsilon)$ containing $x$ such that $d(\sigma(u),\sigma(x))<\varepsilon$ for all $u\in V$. Note that
(8.4) $\sigma^{i}(V)\cap V=\emptyset\hskip 7.0pt\text{for all }i\geq 1,$
otherwise $X$ would have an $\varepsilon$-pseudoloop based at $x$. Let $\Omega=\bigcup_{i=0}^{\infty}\sigma^{i}(V)$. Then $\Omega$ is compact open because $X$ is compact and totally disconnected. By 8.4, we have $\sigma(\Omega)\subset\Omega$ and $\sigma(\Omega)\cap V=\emptyset$. So $\sigma(\Omega)\subsetneq\Omega$. Let $f=1_{\sigma(\Omega)}$, define $h=\sigma_{\ast}(f)-f$. Then, for all $y\in X$, we have
$\sigma_{\ast}(f)(y)=\sum_{u:\sigma(u)=y}1_{\sigma(\Omega)}(u)\geq 1_{\Omega}(y)\geq 1_{\sigma(\Omega)}(y)=f(y).$
This implies that $h\in C_{c}(X,\mathbb{N})$. Moreover,
$h=\sigma_{\ast}(f)-f\geq 1_{\Omega}-1_{\sigma(\Omega)}=1_{\Omega\setminus\sigma(\Omega)}\neq 0.$
Therefore, $(iv)$ is false. ∎
Remark 8.14.
When $X$ is not compact, the implication $(iv)\Rightarrow(v)$ might not be true. For example, consider $X=\mathbb{Z}$ and define $\sigma:X\rightarrow X$ by $\sigma(n)=n=1$. Then $C^{*}(\mathcal{G})$ is AF embeddable, but there are no $\varepsilon$-pseudoloops for $20<\varepsilon\leq 1/$ 2.
Remark 8.15.
Note that for graph algebras and for C*-algebras of compact topological graphs, the C*-algebra is AF embeddable if, and only if, it is finite. We do not know if this equivalence for C*-algebras of Deaconu-Renault groupoids. We believe this is an open question.

### v2

@@
-AF Embeddability of the C*-Algebra of a Deaconu-Renault Groupoid
+Positivity of Matui’s HK Conjecture for AF Groupoids
 Abstract.
-We study Deaconu-Renault groupoids corresponding to surjective local homeomorphisms on locally compact, Hausdorff, second countable, totally disconnected spaces, and we characterise when the C*-algebras of these groupoids are AF embeddable.
-Our main result generalises theorems in the literature for graphs and for crossed products of commutative C*-algebras by the integers.
-We give a condition on the surjective local homeomorphism that characterises the AF embeddability of the C*-algebra of the associated Deaconu-Renault groupoid.
-In order to prove our main result, we generalise an application of Matui’s HK conjecture by Farsi, Kumjian, Pask, and Sims, that gives an isomorphism from the homology groups of AF algebras and the corresponding K-theory.
-In this paper, we give an explicit formula for this isomorphism, and we show that this isomorphism is an order isomorphism.
+In this paper we generalise an application of Matui’s HK conjecture by Farsi, Kumjian, Pask, and Sims, that gives an isomorphism from the homology groups of AF groupoids to the corresponding K-theory.
+We give an explicit formula for this isomorphism, and we show that the map is an order isomorphism.
+Since homology groups are equipped with several useful techniques, this map can help us to understand the K-theory of the C*-algebra in more detail.
+To illustrate this, we apply the isomorphism to characterise the AF embeddability of the C*-algebra of Deaconu-Renault groupoids.
 Key words and phrases:
@@
 1. Introduction
-A problem that has received much interest during the last decades is the classification of C*-algebras.
-This problem consists of finding an invariant that identifies all C*-algebras of a certain class up to isomorphism.
-One of the first results in the area is Glimm’s [22] classification of some uniformly hyperfinite C*-algebras.
-Glimm’s result was interesting because uniformly hyperfinite algebras are the C*-algebraic version of the $\text{II}_{1}$ factors for von Neumann algebras, but unlike the von Neumann case [42] (see also [12, Theorem V.1.13]), uniformly hyperfinite C*-algebras are not all isomorphic.
-Dixmier [15] then generalised this classification by studying a class of C*-algebras called matroid C*-algebras.
-In [6], Bratteli defined a more general class of C*-algebras, called approximately finite dimensional C*-algebras (or AF algebras).
-An AF algebra is the completion of an increasing union of finite dimensional C*-algebras.
-Elliott [19, Theorem 4.3] (see [13, Theorem IV.4.3] for a more modern formulation) classified all AF algebras by their K-theory.
-Since Elliott’s classification of AF algebras, different types of C*-algebras have been classified by invariants.
-Several stably finite C*-algebras [23, 24] are classifiable.
-Kirchberg [30] and Phillips [43] classified all unital, simple, separable, nuclear and purely infinite C*-algebras (called Kirchberg algebras) satisfying the Universal Coefficient Theorem by their K-theory.
-See Strung’s book on the classification of C*-algebras [56] for more details.
-In certain types of C*-algebras, a dichotomy exists between purely infinite and stably finite.
-In other words, for certain classes of C*-algebras, a C*-algebra is either stably finite or purely infinite.
-This dichotomy holds for reduced C*-algebras of certain minimal, topological principal, ample groupoids, Bönicke and Li [5, Theorem D], Rainone and Sims [47, Theorem 7.4], and Ma [37].
-Additionally to the classification programme, the possibility of finding a dichotomy for large classes of C*-algebras motivates us to study purely infinite and stably finite C*-algebras, and related properties.
-Pimsner and Voiculescu [44] proved that the irrational rotation C*-algebra is embedded into the C*-algebra from [18], which is AF and whose $K_{0}$-group is isomorphic to $\mathbb{Z}+\theta\mathbb{Z}$.
-This embedding led to the classication of irrational rotation C*-algebras in terms of the angle of rotation.
-Effros [17] observed the need of investigating subalgebras of AF algebras, since virtually nothing was known about them at the time.
-Such algebras are denoted to be AF embeddable (or AFE algebras).
-Since the formulation of the conjecture that a C*-algebra is AF embeddable if, and only if, it is separable, exact, and quasidiagonal (see [4, Section 7]), the problem of characterising AFE algebras has received much attention.
-Pimsner [45] characterised the AFE property for crossed products of commutative C*-algebras by the integers.
-Brown [8] generalised this result to crossed products of AF algebras to the integers, and that the notions of stably finite, quasidiagonal and AF embeddable are equivalent for these crossed products.
-Brown [7] proved that the crossed product of UHF algebra by $\mathbb{Z}^{k}$ is AF embeddable.
-Similar results are given in [25, 34, 35, 38].
-Liu [36] described the AF embeddability of separable exact quasidiagonal C*-algebras in terms of a condition on the C*-algebras representations.
-Schafhauser [52, Theorem 6.7] characterised the AF embeddability for topological graph algebras with no sinks, where the notions of stably finite, quasidiagonal and AF embeddable are equivalent.
-Topological graph algebras are very general and cover the C*-algebras of Deaconu-Renault groupoids, which we study in this paper.
-So, the equivalence of these notions holds for the C*-algebras we are interested.
-In the context of graph algebras, there are several results in the literature studying C*-algebras that are AF, AF embeddable or purely infinite.
-For example, graph loops characterises when a graph algebra is approximately finite dimensional [32].
-Similarly, Schafhauser [54] describes when a graph algebra is AF embeddable, while [3, Proposition 5.3] gives a graph condition that makes the corresponding C*-algebra purely infinite.
-Szymański [57] showed that every stable, purely infinite, simple and classifiable C*-algebra with $K_{1}$ torsion-free is a graph algebra.
-Kumjian and Pask [31] study higher rank graph algebras and give sufficient conditions that make the C*-algebras AF or purely infinite.
-Clark, an Huef, and Sims [11] characterised the AF embeddability of C*-algebras of row-finite and cofinal $k$-graphs with now sources, for $k=2$.
-Schafhauser [53, Corollary 6.2] extendded this result for an arbitrary $k$.
-In this paper we study C*-algebras of Deaconu-Renault groupoids.
-These algebras generalise graph algebras [33] and are examples of topological graph algebras [27, Proposition 10.9].
-In Subsection 2.1, we describe this class of groupoids in more detail.
-The main contribution of this paper is the description of the AF embeddability of the C*-algebras of Deaconu-Renault groupoids by a corresponding groupoid property.
-Our result is analogous to similar results for graph algebras [54], for crossed products of unital commutative C*-algebras by the integers [9, Theorem 11.5], [45, Theorem 9], and for higher rank graphs [11], [53, Corollary 6.2], cited above.
-In particular, we apply the same strategy from [11], by applying [8, Theorem 0.2] to describe AF embeddability in terms of a K-theory condition, and them analysing this condition to find a more concrete condition which can be described by an equation.
-We state our main result below.
+Matui [20] studied how homology groups of groupoids can describe the dynamical properties the topological full groups of a groupoid.
+There, the homology groups $H_{0}(G)$ and $H_{1}(G)$ for étale groupoids over totally disconnected spaces are described by adapting the construction from Crainic and Moerdijk [8].
+In [21, Conjecture 2.6], Matui formulated the HK conjecture, which connects the homology groups with the K-groups of the associated C*-algebras, stated below:
+Conjecture 1.1 (Matui’s HK conjecture).
+Let $G$ be a locally compact, Hausdorff, étale groupoid, such that $G^{(0)}$ is a Cantor set. Suppose that $G$ is both effective and minimal. Then for $j=0,1$ we have
+$K_{j}(C^{*}_{r}(G))\cong\bigoplus_{i=0}^{\infty}H_{2i+j}(G).$
+This conjecture holds for AF groupoids with compact unit spaces [20, Theorem 4.10], and for étale groupoids arising from subshifts of finite type [20].
+Other examples can be found in [3, 22, 23, 25, 27, 41, 44].
+Proietti and Yamashita [28] found an example beyond the original assumptions of the conjecture, while Scarparo [35], Deeley [10], and Ortega and Sanchez [24] provide counterexamples.
+Farsi, Kumjian, Pask, and Sims [12] proved Matui’s HK Conjecture for Deaconu-Renault groupoids of ranks one and two, and for k-graph groupoids of single vertex k-graphs which satisfy a mild joint-coprimality condition.
+Moreover, they generalised Matui’s [20, Theorem 4.10], showing in [12, Corollary 5.2] that the conjecture holds for AF groupoids with non-compact unit spaces.
+Additionally, Farsi et al.
+show that, for an AF groupoid $G$, there exists an isomorphism between the K-group $K_{0}(C^{*}(G))$ and the corresponding homology group $H_{0}(G)$.
+This map is such that $[1_{V}]_{0}\mapsto[1_{V}]$ for every compact open set $V\subset G^{(0)}$.
+However, it is not clear from their paper whether this map is an isomorphism of ordered groups, i.e., that it preservers positive maps from one group to the other.
+In this paper, we show that the map provided in Farsi et al. for AF groupoids is an ordered group isomorphism. Moreover, we give an explicit formula for the isomorphism, as we state below.
+Theorem 1.
+Let $G$ be an AF groupoid, written as the inductive limit $G=\displaystyle\lim_{\rightarrow}G_{n}$ of elementary groupoids.
+Then, for every $n$, there exists a locally compact, Hausdorff, second countable, totally disconnected space $Y_{n}$, and there is a surjective local homeomorphism $\sigma_{n}:G_{n}^{(0)}\rightarrow Y_{n}$ such that $G_{n}=R(\sigma_{n})$.
+Moreover,
+  1. 1.
+    there exists a unique isomorphism $\nu:K_{0}(C^{*}(G))\rightarrow H_{0}(G)$ defined by
+    $\nu([p]_{0})=[\mathrm{tr}_{\varphi_{n}}(p)]$
+    for $p\in\mathcal{P}_{n}(C^{*}(G_{n}))$, and such that $\varphi_{n}$ is an arbitrary continuous section of $\sigma_{n}$;
+  2. 2.
+    the map $\nu$ is an isomorphism of ordered groups, i.e., $\nu(K_{0}(C^{*}(G))^{+})=H_{0}(G)^{+}$; and
+  3. 3.
+    the isomorphism $\nu$ is precisely the map given by [12, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
+The theorem presents two advantages.
+First, K-groups are important in the theory of C*-algebras, since several results rely on them.
+However, K-theory is often too abstract and hard to calculate.
+In contrast, homology groups are more manageable and are equipped with techniques that make them easier to understand.
+So, for the C*-algebra of an AF groupoid, the explicit formula above can be applied to understand the K-theory in more detail.
+Second, in some situations, we need to know the positive elements of the $K_{0}$-group, so an application of Matui’s HK conjecture would be useless unless we guarantee that the isomorphism preserves positive elements from one group to another.
+Note that Theorem 1 brings another question: for other groupoids where Matui’s HK conjecture hold, is the group isomorphism $K_{0}(C^{*}_{r}(G))\rightarrow\bigoplus_{i=0}^{\infty}H_{2i}(G)$ also an ordered group isomorphism?
+In the second part of the paper, we apply Theorem 1 to solve a more practical problem, characterising when the C*-algebra of a large class of Deaconu-Renault groupoids is AF embeddable.
+The example illustrates the power of homology techniques when one needs to manipulate the elements of the $K_{0}$-group.
+We state the theorem as follows:
+Definition.
+Let $\sigma:X\rightarrow Y$ be a surjective local homeomorphism on locally compact, Hausdorff, second countable, totally disconnected spaces, and let $\varphi:Y\rightarrow X$ be a continuous section of $\sigma$.
+We define the map $\mathrm{tr}_{\varphi}:\mathcal{P}_{\infty}(C^{*}(R(\sigma)))\rightarrow C_{c}(X,\mathbb{N})$ by
+  • $\mathrm{tr}_{\varphi}(p)(x)=1_{X_{\varphi}}(x)\displaystyle\sum_{y:\sigma(x)=\sigma(y)}p(y)$,       for $p\in C^{*}(R(\sigma))$, $x\in X$, and
+  • $\mathrm{tr}_{\varphi}(p)(x)=\displaystyle\sum_{i=1}^{n}\mathrm{tr}_{\varphi}(p_{ii})(x)$,                  for $p\in\mathcal{P}_{n}(C^{*}(R(\sigma)))$, $n\geq 1$,
+where we identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$ by applying the $j$ map from [30, Proposition II.4.2] and [40, Proposition 9.3.3].
 Theorem 2.
@@
 $\sigma_{\ast}(f)(x)=\sum_{y:\sigma(y)=x}f(y)\hskip 20.0pt\text{for $x\in X$, $f\in C_{c}(X,\mathbb{Z})$.}$
-Note that C*-algebras of Deaconu-Renault groupoids are examples of topological graph algebras.
-When $X$ is compact, then the corresponding topological graph algebra is unital and, in this case, Theorem 1 is an example of [52, Theorem 6.7], that characterises when a unital topological graph algebra is AF embeddable.
 The construction of the Deaconu-Renault groupoid $\mathcal{G}$ depends on the surjective local homeomorphism $\sigma$.
 The contribution of Theorem 2 is that it provides a condition on $\sigma$ that is equivalent to the AF embeddability of the C*-algebra of the corresponding Deaconu-Renault groupoid.
+This theorem is analogous to similar results for graph algebras [38], for crossed products of unital commutative C*-algebras by the integers [5, Theorem 11.5], [26, Theorem 9].
+Schafhauser [39, Theorem C] studied Cuntz-Pimsner algebras.
+Kumjian and Pask [18] study higher rank graph algebras and give sufficient conditions that make the C*-algebras AF or purely infinite.
+Clark, an Huef, and Sims [7] characterised the AF embeddability of C*-algebras of row-finite and cofinal $k$-graphs with now sources, for $k=2$.
+Schafhauser [37, Corollary 6.2] extendded this result for an arbitrary $k$.
+In particular, we apply the same strategy from [7], by applying [4, Theorem 0.2] to describe AF embeddability in terms of a K-theory condition, and then analysing this condition to find a more concrete condition which can be described by the equation (iv) of the Theorem 2.
+Note that C*-algebras of Deaconu-Renault groupoids generalise graph algebras [19] and are examples of topological graph algebras [15, Proposition 10.9].
+When $X$ is compact, then the corresponding topological graph algebra is unital and, in this case, Theorem 2 is an example of [36, Theorem 6.7], that characterises when a unital topological graph algebra is AF embeddable.
 In order to prove Theorem 2, we apply the following result by Brown [4], that gives an equivalent condition to AF embeddability of crossed products of AF algebras by the integers:
@@
 Theorem 1.2.
-(Brown) [8, Theorem 0.2] If $B$ is an AF algebra and $\beta\in\mathrm{Aut}(B)$, then the following are equivalent:
+(Brown) [4, Theorem 0.2] If $B$ is an AF algebra and $\beta\in\mathrm{Aut}(B)$, then the following are equivalent:
   1. (a)
@@
     $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$.
-Note that AF embeddability, quasidiagonality and the stable finite properties are preserved by stable isomorphisms.
+Note that AF embeddability, quasidiagonality and stable finiteness are preserved by stable isomorphisms.
 If we choose an AF algebra $B$ such that $B\rtimes_{\beta}\mathbb{Z}$ is stably isomorphic to $C^{*}(\mathcal{G})$, then we can apply Brown’s theorem to get a characterisation of the AF embeddability of $C^{*}(\mathcal{G})$.
 Indeed, we let $c:\mathcal{G}\rightarrow\mathbb{Z}$ be the canonical cocycle, and we let $B=C^{*}(\mathcal{G}(c))$ be the C*-algebra of the skew product $\mathcal{G}(c)$.
-This C*-algebra is AF by [20, Lemma 6.1].
+This C*-algebra is AF by [12, Lemma 6.1].
 Theorem III.5.7 of [30] gives the isomorphism
@@
 Note that item (d) depends on the K-theory, which is very abstract.
+Moreover, in order to understand (d), we need to know the positive elements of $K_{0}(B)$.
 We use the theory of homology groups of groupoids in order to show that (d) is equivalent to condition (iv) of Theorem 2.
@@
 […]
-The existence of the isomorphism $K_{0}(B)\rightarrow H_{0}(\mathcal{G}(c))$ is given by [20, Corollary 5.2], which is an application of Matui’s HK conjecture [39, 40, 41], and is such that $[1_{V}]_{0}\mapsto[1_{V}]$ for every compact open $V\subset\mathcal{G}^{(0)}$.
+It follows from [12, Corollary 5.2] that $\mathcal{G}(c)$ is AF groupoid.
+Then Theorem 1 gives the ordered group isomorphism $K_{0}(B)\rightarrow H_{0}(\mathcal{G}(c))$.
 The groupoid $c^{-1}(0)$ is the kernel of the cocycle $c:\mathcal{G}\rightarrow\mathbb{Z}$ defined by $c(x,k,y)=k$.
-We define the isomorphism $H_{0}(\rho_{\varphi}):H_{0}(\mathcal{G})(c))\rightarrow H_{0}(c^{-1}(0))$ in Section 6.
+We define the isomorphism $H_{0}(\rho_{\varphi}):H_{0}(\mathcal{G})(c))\rightarrow H_{0}(c^{-1}(0))$ in Section 5.3.
 The homomorphisms $K_{0}(\beta)$ and $[\widetilde{\beta}]_{\mathcal{G}(c)}$ are group homomorphisms induced by $\beta$, and the group homomorphism $[\sigma_{\ast}]$ is induced by the map $\sigma_{\ast}$ of Theorem 2.
 By studying the diagram and by applying properties of homology groups of groupoids, we can show that items (d) and (iv) are equivalent, proving Theorem 2.
-We need to know the positive elements of $K_{0}(B)$ in order to study item (d).
-So, when we apply the commutative diagram, we need to guarantee that the isomorphisms above takes positive elements from one group to the another.
-Although [20, Corollary 5.2] gives the existence of the group isomorphism $K_{0}(B)\cong H_{0}(\mathcal{G}(c))$, we need this isomorphism to be an ordered group isomorphism.
-We prove that this isomorphism is indeed a ordered group isomorphism using different techniques, and we give an explicit formula for the map, stated as follows:
-Definition.
-Let $\sigma:X\rightarrow Y$ be a surjective local homeomorphism on locally compact, Hausdorff, second countable, totally disconnected spaces, and let $\varphi:Y\rightarrow X$ be a continuous section of $\sigma$.
-We define the map $\mathrm{tr}_{\varphi}:\mathcal{P}_{\infty}(C^{*}(R(\sigma)))\rightarrow C_{c}(X,\mathbb{N})$ by
-  • $\mathrm{tr}_{\varphi}(p)(x)=1_{X_{\varphi}}(x)\displaystyle\sum_{y:\sigma(x)=\sigma(y)}p(y)$,       for $p\in C^{*}(R(\sigma))$, $x\in X$, and
-  • $\mathrm{tr}_{\varphi}(p)(x)=\displaystyle\sum_{i=1}^{n}\mathrm{tr}_{\varphi}(p_{ii})(x)$,                  for $p\in\mathcal{P}_{n}(C^{*}(R(\sigma)))$, $n\geq 1$,
-where we identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$ by applying the $j$ map from [48, Proposition II.4.2] and [55, Proposition 9.3.3].
-Theorem 2.
-Let $G$ be an AF groupoid, written as the inductive limit $G=\displaystyle\lim_{\rightarrow}G_{n}$ of elementary groupoids.
-Then, for every $n$, there exists a locally compact, Hausdorff, second countable, totally disconnected space $Y_{n}$, and there is a surjective local homeomorphism $\sigma_{n}:G_{n}^{(0)}\rightarrow Y_{n}$ such that $G_{n}=R(\sigma_{n})$.
-Moreover,
-  1. 1.
-    there exists a unique isomorphism $\nu:K_{0}(C^{*}(G))\rightarrow H_{0}(G)$ defined by
-    $\nu([p]_{0})=[\mathrm{tr}_{\varphi_{n}}(p)]$
-    for $p\in\mathcal{P}_{n}(C^{*}(G_{n}))$, and such that $\varphi_{n}$ is an arbitrary continuous section of $\sigma_{n}$;
-  2. 2.
-    the map $\nu$ is an isomorphism of ordered groups, i.e., $\nu(K_{0}(C^{*}(G))^{+})=H_{0}(G)^{+}$; and
-  3. 3.
-    the isomorphism $\nu$ is precisely the map given by [20, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
-This theorem also generalises Matui’s [39, Theorem 4.10], which also gave an order isomorphism for AF groupoids with compact unit spaces, but did not provide a formula for the isomorphism.
-So, we extend Matui’s result to AF groupoids with non-compact unit spaces.
-Moreover, by showing that $H_{0}(G)^{+}\cong K_{0}(C^{*}(G))^{+}$, Theorem 2 suggests a stronger version of Matui’s HK conjecture, which considered only group homomorphisms.
-We organise the paper as follows: in Section 2, we define Deaconu-Renault groupoids, inductive limits and AF groupoids, presenting some results from the literature.
-In Section 3, we state the isomorphism by [48, Proposition II.5.1] between the C*-algebra of a skew-product groupoid and its corresponding crossed product.
-We also define the map $\beta$, used later in the commutative diagram.
-Then we study the theory of homology groups for groupoids in Section 4.
-In Section 5, we prove Theorem 2, giving an isomorphism between the homology groups for AF groupoids and the K-theory of the C*-algebras of these groupoids.
-Similar results have been proved in [20, Corollary 5.2], [39, Theorem 4.10].
-Our innovation is that we give an explicit formula for this isomorphism, and we show that it preserves positive elements from one group to another.
-In Section 6 we apply the results from Sections 4 and 5 to prove the commutative diagram (1.1).
-In Section 7 we prove Theorem 1, our main result, that describes under certain conditions when the C*-algebra of a Deaconu-Renault groupoid is AF embeddable.
-Finally, we show in Section 8 that our main theorem generalises some known results in the literature.
+We organise the paper as follows: in Section 2, we study inductive limits, AF groupoids, abstract homology groups, and homology groups for groupoids.
+In Section 3, we study the homology groups of AF groupoids, while in Section 4 we prove Theorem 1, our main result.
+In Section 5, we apply the isomorphism from Theorem 1 and the tools from homology groups to characterise the AF embeddability of the C*-algebra of the Deaconu-Renault groupoid by Theorem 2.
+Finally, in Section 6, we show that this theorem generalises known results in the literature.
 The author would like to thank Astrid an Huef and Lisa Orloff Clark for their supervision, Camila F.
@@
 2. Preliminaries
-2.1. Deaconu-Renault groupoids
-Given the topological spaces $X$ and $Y$, a map $\sigma:X\rightarrow Y$ is a local homeomorphism if every $x\in X$ has an open neighbourhood $\mathcal{U}$ such that $\sigma(\mathcal{U})$ is open and $\sigma|_{\mathcal{U}}:\mathcal{U}\rightarrow\sigma(\mathcal{U})$ is a homeomorphism.
-A continuous map $\varphi:Y\rightarrow X$ is a continuous section of $\sigma$ if $y=\sigma(\varphi(y))$ for all $y$.
-Using the argument from the proof of [20, Theorem 4.10], one can show that every surjective local homeomorphism has a continuous section.
-Let $X$ be a locally compact, Hausdorff, second countable space, and $\sigma:X\rightarrow X$ is a surjective local homeomorphism. We define the set
-$\mathcal{G}=\{(x,k,y)\in X\times\mathbb{Z}\times X:\exists m,n\in\mathbb{N},k=m-n,\sigma^{m}(x)=\sigma^{n}(y)\},$
-and we endow it with the range and source maps $r,s:\mathcal{G}\rightarrow\mathcal{G}$ with $r(x,k,y)=(x,0,x)$ and $s(x,k,y)=(y,0,y)$.
-Let $\mathcal{G}^{(2)}=\{((x,k_{1},y),(y,k_{2},z)):(x,k_{1},y),(y,k_{2},z)\in\mathcal{G}\}$ and we define the product $\mathcal{G}^{(2)}\rightarrow\mathcal{G}$ and inverse operations $\mathcal{G}\rightarrow\mathcal{G}$ by
-$(x,k_{1},y)(y,k_{2},z)=(x,k_{1}+k_{2},z),\text{\hskip 10.0pt}(x,k,y)^{-1}=(y,-k,x).$
-With these operations, $\mathcal{G}$ becomes a groupoid.
-The unit space is $\mathcal{G}^{(0)}=\{(x,0,x):x\in X\}$, which is identified with $X$ by the map $(x,0,x)\mapsto x$.
-For the open subsets $A,B\subset X$ and for the natural numbers $m,n$ such that $\sigma|_{A}^{m}$ and $\sigma|_{B}^{n}$ are injective with $\sigma^{m}(A)=\sigma^{n}(B)$, define the subset
-$\mathcal{U}_{A,B}^{m,n}=\{(x,m-n,y)\in A\times\mathbb{Z}\times B:\sigma^{m}(x)=\sigma^{n}(y)\}.$
-Here we use the notation from [58].
-It follows from [14, Theorem 1] that the sets $\mathcal{U}_{A,B}^{m,n}$ generate a topology that makes $\mathcal{G}$ a topological groupoids which is locally compact, Hausdorff, second countable, étale.
-We denote the groupoid $\mathcal{G}$ endowed with this topology a Deaconu-Renault groupoid.
-The first example of this groupoid was described in Renault’s book [48, Section III.2] by a class of groupoids whose C*-algebras are Cuntz algebras, where $X$ is the set of sequences $(x_{i})_{i\in\mathbb{N}}$, with $x_{i}\in\{0,1,\dots,n\}$ (or $x_{i}\in\mathbb{N}$ for the Cuntz algebra $\mathcal{O}_{\infty}$), and $\sigma$ was given by $\sigma(x_{0}x_{1}x_{2}\dots)=x_{1}x_{2}\dots$ Deaconu [14] generalises Renault’s definition by considering compact Hausdorff spaces $X$ and assumes that $\sigma:X\rightarrow X$ is a covering map.
-Under these assumptions, he proved that when $\sigma$ is a homeomorphism, then $C^{*}(\mathcal{G})$ is isomorphic to the crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$.
-Anantharaman-Delaroche [1, Example 1.2 (c)] generalised it to the case where $\sigma$ is a surjective local homeomorphism, and $X$ is locally compact.
-She also gave a sufficient conditions to make the C*-algebra of this groupoid purely infinite [1, Propositions 4.2 and 4.3].
-Then Renault [49] generalised this definition by considering a map $\sigma$ that is defined only on an open subset of $X$.
-For the purpose of our paper, here we consider the definition by Anantharaman-Delaroche, and we impose the additional conditions on the space $X$ to be Hausdorff and second countable.
-By [49, Proposition 2.9], the groupoid is also amenable.
-Note that there are different notions of amenability for groupoids: topological amenable and measurewise amenable.
-By Theorem 3.3.7 of [2], these two notions coincide for groupoids that are locally compact, Hausdorff and étale.
-In this case, [2, Proposition 6.1.8] implies that the reduced and full C*-algebras of the groupoid coincide.
 2.1. Inductive limits
-An inductive sequence in a category $\mathscr{C}$ is a sequence $\{A_{n}\}_{n=1}^{\infty}$ of objects in $\mathscr{C}$ and a sequence $\varphi_{n}:A_{n}\rightarrow A_{n+1}$ of morphisms in $\mathscr{C}$, usually written
+An inductive sequence in a category $\mathscr{C}$ is a sequence $\{A_{n}\}_{n=1}^{\infty}$ of objects in $\mathscr{C}$ and a sequence $\varphi_{n}:A_{n}\rightarrow A_{n+1}$ of morphisms in $\mathscr{C}$, usually written as
 […]
@@
 Inductive limits, when they exist, are essentially unique in the sense that if $(A,\{\mu_{n}\})$ and $(B,\{\lambda_{n}\})$ are inductive limits of the same inductive sequence, then there is an isomorphism $\lambda:A\rightarrow B$ making the diagram of item (ii) above commutative.
-See [51, page 93] for details.
+See [34, page 93] for details.
 Because of the essential uniqueness of inductive limits (when they exist), we shall refer to the inductive limit (rather than an inductive limit).
@@
 There are different definitions for AF groupoids in the literature.
-The first was given by Renault [48, Definition III.1.1] in his book, and he proved in [48, Proposition III.1.15] that AF groupoids determine AF algebras up to isomorphism.
-Because of this result, we can see AF groupoids as the groupoid analogs of AF algebras.
-Other definitions for such groupoids have been introduced in the literature [50], [21, Definition 3.7], [39, Definition 2.2], and [20, Definition 4.9].
-In [10] there is a proof of the equivalence of the definitions [48, Definition III.1.1] and [20, Definition 4.9].
+The first was given by Renault [30, Definition III.1.1] in his book, and he proved in [30, Proposition III.1.15] that AF groupoids determine AF algebras up to isomorphism.
+Because of this result, we can see AF groupoids as groupoid analogues of AF algebras.
+Other definitions for such groupoids have been introduced in the literature [32], [13, Definition 3.7], [20, Definition 2.2], and [12, Definition 4.9].
+In [6] there is a proof of the equivalence of the definitions [30, Definition III.1.1] and [12, Definition 4.9].
 In this paper, we consider the Definition 4.9 of [12], by Farsi, Kumjian, Pask, and Sims.
+Given the topological spaces $X$ and $Y$, a map $\sigma:X\rightarrow Y$ is a local homeomorphism if every $x\in X$ has an open neighbourhood $\mathcal{U}$ such that $\sigma(\mathcal{U})$ is open and $\sigma|_{\mathcal{U}}:\mathcal{U}\rightarrow\sigma(\mathcal{U})$ is a homeomorphism.
+A continuous map $\varphi:Y\rightarrow X$ is a continuous section of $\sigma$ if $y=\sigma(\varphi(y))$ for all $y$.
+Using the argument from the proof of [12, Theorem 4.10], one can show that every surjective local homeomorphism has a continuous section.
 Definition 2.2.
@@
 If $G$ an AF groupoid is the increasing union of the open elementary subgroupoids $G_{n}$.
-Then $G$ is the inductive limit $G=\displaystyle\lim_{\longrightarrow}G_{n}$ [48, Section III.1].
+Then $G$ is the inductive limit $G=\displaystyle\lim_{\longrightarrow}G_{n}$ [30, Section III.1].
 Remark 2.4.
-Given an elementary groupoid $G\cong R(\sigma)$ with $\sigma:X\rightarrow Y$, one we apply the isomorphism to find a local homeomorphism $\widetilde{\sigma}:G^{(0)}\rightarrow Y$ such that $G=R(\widetilde{\sigma})$.
+Given an elementary groupoid $G\cong R(\sigma)$ with $\sigma:X\rightarrow Y$, we apply the isomorphism to find a local homeomorphism $\widetilde{\sigma}:G^{(0)}\rightarrow Y$ such that $G=R(\widetilde{\sigma})$.
 So, we can suppose that $G=R(\sigma)$ without loss of generality.
@@
 In fact, it is the increasing union of the open elementary subgroupoids $R(\sigma^{n})$.
-We use the notation $c^{-1}(0)$ because of the following.
-If $\mathcal{G}$ is the Deaconu-Renault groupoid and $c:\mathcal{G}\rightarrow\mathbb{Z}$ is the canonical cocycle, then $c^{-1}(0)$ is isomorphic to the subgroupoid $\{(x,0,y)\in\mathcal{G}\}$ of $\mathcal{G}$.
-The isomorphism is defined by $(x,y)\mapsto(x,0,y)$.
-3. Skew-products and crossed products
-As explained in the introduction, we use an isomorphism between the C*-algebras a skew-product groupoid and the corresponding crossed product to understand when the C*-algebra of a Deaconu-Renault groupoid $\mathcal{G}$ is AFE.
-This isomorphism allows us to apply the results of Brown’s Theorem (Theorem 1.1).
-In this section, we use Renault’s book [48] as a reference.
-Definition 3.1.
-[48, Definition I.1.6] Let $G$ be a groupoid and $c:G\rightarrow\mathbb{Z}$ a homomorphism. Define the sets
-$G(c)=\{(g,a):g\in G,a\in\mathbb{Z}\}\hskip 10.0pt\text{and}\hskip 10.0ptG(c)^{(0)}=G^{(0)}\times\mathbb{Z}.$
-Define the maps $r,s:G(c)\rightarrow G^{(0)}\times\mathbb{Z}$ by $r(g,a)=(r(g),a)$ and $s(g,a)=(s(g),a+c(g))$. Consider the set
-(3.1) $G(c)^{(2)}=\{((g,a),(h,a+c(g))):(g,h)\in G^{(2)},a\in\mathbb{Z}\},$
-and define the product $G(c)^{(2)}\rightarrow G(c)$ and the inverse $G(c)\rightarrow G(c)$ by
-(3.2) $(g,a)(h,a+c(g))=(gh,a)\hskip 15.0pt\text{and}\hskip 15.0pt(g,a)^{-1}=(g^{-1},a+c(g)),$
-respectively. Then $G(c)$ becomes a groupoid, called a skew-product.
-If $c$ is a continuous cocycle (i.e., if $c:G\rightarrow\mathbb{Z}$ is a continuous homomorphism) and we equip $G(c)$ with the induced topology from $G\times\mathbb{Z}$, then $G(c)$ inherits the usual topogical properties of $G$.
-Recall that we assume that $G$ is locally compact, Hausdorff, second countable and étale.
-Lemma 3.2.
-Equip $G(c)$ with the topology induced from $G\times\mathbb{Z}$.
-Then $G(c)$ is also a locally compact, Hausdorff, second countable, étale groupoid.
-Moreover, if $G$ is totally disconnected, then $G(c)$ is totally disconnected, too.
-If $G$ is measurewise amenable, then $G(c)$ is also measurewise amenable.
-Note that can identify the unity space $G(c)^{(0)}$ with $\mathcal{X}\times\mathbb{Z}$.
-For $f,f_{1},f_{2}\in C_{c}(G(c))$ and $(g,a)\in G\times\mathbb{Z}$, the operations on $C_{c}(G(c))$ are given by
-$(f_{1}\cdot f_{2})(g,a)=\sum_{g_{1}g_{2}=g}f_{1}(g_{1},a)f_{2}(g_{2},a+c(g_{1})),\text{ and }f^{*}(g,a)=\overline{f(g^{-1},a+c(g))}.$
-Assume that $G$ is a locally compact, Hausdorff, second countable, étale, measurewise amenable groupoid, and suppose that $c:G\rightarrow\mathbb{Z}$ is a continuous cocycle.
-We define a dynamical system $\alpha^{c}$ induced by $c$.
-For every $z\in\mathbb{T}$, $f\in C_{c}(G)$ and $g\in G$, we set $\alpha_{z}^{c}(f)(g)=z^{c(g)}f(g)$.
-By Proposition II.5.1 of Renault’s book [48], there exists a unique dynamical system defined by the equation above.
-We make an abuse of notation by denoting this dynamical system $\alpha^{c}:\mathbb{T}\rightarrow\mathrm{Aut}(C^{*}(G))$.
-The following proposition by Renault [48, Proposition II.5.7] gives an isomorphism between the C*-algebra of skew-product groupoids and the corresponding crossed products.
-For an introduction to crossed products, see Chapters 1 and 2 of Dana William’s book [59].
-Proposition 3.3.
-There exists a unique isomorphism $\rho\rtimes u:C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rightarrow C^{*}(G(c))$ such that
-[…]
-for $f\in C_{c}(\mathbb{T},C_{c}(G))$ and $(g,a)\in G(c)$. Moreover, the inverse of $\rho\rtimes u$ is given by
-$(\rho\rtimes u)^{-1}(F)(z)(g)$ $=\frac{1}{2\pi}\sum_{a\in\mathbb{Z}}z^{c(g)+a}F(g,a),$
-for $F\in C_{c}(G(c))$, $z\in\mathbb{T}$, $g\in G$.
-Now we find a formula for the action $\beta$ on $C^{*}(G(c))$ induced by the dual action $\widehat{\alpha}^{c}$.
-We will apply this formula in Chapter 7 for Deaconu-Renault groupoids in order to understand in more detail the condition $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$, given by Brown’s theorem.
-Proposition 3.4.
-Let $\rho\rtimes u:C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rightarrow C^{*}(G(c))$ be the isomorphism of Proposition 3.3.
-Then there exists a unique dynamical system $\beta:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G(c)))$ such that
-$\beta(F)(g,a)=F(g,a+1)\hskip 15.0pt\text{for $F\in C_{c}(G(c))$, $(g,a)\in G(c)$,}$
-and $\beta=(\rho\rtimes u)\circ\widehat{\alpha}^{c}\circ(\rho\rtimes u)^{-1}$ for all $n\in\mathbb{Z}$.
-Moreover, $C^{*}(G(c))\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic.
-Proof.
-Let $\widehat{\alpha}^{c}:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T})$ be the dual action, which is given by $\widehat{\alpha}^{c}_{n}(f)(z)=z^{-n}f(z)$, for $f\in C(\mathbb{T},C^{*}(G))$, $z\in\mathbb{T}$, $n\in\mathbb{Z}$.
-For each $n\in\mathbb{Z}$, let $\beta_{n}=(\rho\rtimes u)^{-1}\circ\widehat{\alpha}^{c}_{n}\circ\rho\rtimes u$.
-Since $\rho\rtimes u$ is an isomorphism and $\widehat{\alpha}^{c}$ is an automorphism, we have that $\beta:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G(c)))$ is a dynamical system.
-Fix $n\in\mathbb{Z}$.
-By continuity of $\beta_{n}$, finding the formula of this automorphism on $C_{c}(G(c))$ is sufficient to determine $\beta_{n}$ uniquely.
-Let $F\in C_{c}(G(c))$ and $(g,a)\in G(c)$.
-Then, Proposition 3.3,
-$\beta_{n}(F)(g,a)$ $=(\rho\rtimes u)\circ\widehat{\alpha^{c}}_{n}\circ(\rho\rtimes u)^{-1}(F)(g,a)$
-[…]
-[…]
-$=\frac{1}{2\pi}\int_{\mathbb{T}}z^{-c(g)-a-n}\sum_{b\in\mathbb{Z}}z^{c(g)+b}F(g,b)dz$
-[…]
-$=F(g,a+n).$
-By definition, we have that $\rho\rtimes u\circ\widehat{\alpha}^{c}=\beta\circ\rho\rtimes u$.
-Then $\rho\rtimes u$ is an equivariant isomorphism from $(C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T},\widehat{\alpha}^{c},\mathbb{Z})$ onto $(C^{*}(G(c)),\beta,\mathbb{Z})$.
-By [59, Lemma 2.65], $C^{*}(G(c))\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic.
-∎
-4. Homology Groups and K-theory
-Our main result, Theorem 1 is an application of Brown’s theorem (Theorem 1.1), which characterises the AF embeddability of a C*-algebra $B\rtimes_{\beta}\mathbb{Z}$ in terms of the K-theory of $B$.
-Note that condition (iv) is very abstract so, in order to provide a more concrete condition, we need to understand the K-theory of the groupoid C*-algebra $C^{*}(\mathcal{G}(c))$.
-The definition of K-theory is very abstract and not very manageable.
-So, we are going to apply an isomorphism between the K-theory $K_{0}(B)$ and its corresponding homology group $H_{0}(\mathcal{G}(c))$, which allows us to apply several helpful results from the theory of homology groups that will allow us to understand what condition (iv) in Brown’s theorem tells us about the groupoid $\mathcal{G}$.
-There is already in the literature an isomorphism between the K-theory and the corresponding homology group for AF algebras, described in [20, Corollary 5.2] by Farsi, Kumjian, Pask and Sims.
-This isomorphism is not sufficient to study condition (iv) of Brown’s theorem in terms of homology groups because it is not clear if this isomorphism preserves positive elements.
-So, we are going to extend [20, Corollary 5.2] by showing in Theorem 2 that the isomorphism preserves positive elements, and we give a more explicit formula for the isomorphism.
-Our idea was to consider the building blocks $R(\sigma)$ of AF groupoids and show that the C*-algebras $C^{*}(R(\sigma))$ are approximated by C*-algebras of the form $C_{0}(V,M_{n})$, whose K-groups are much easier to understand.
-Theorem 2 also generalises [39, Theorem 4.10] which considers only AF groupoids with compact unit spaces.
 2.3. Homology in algebra
-First we present the definition of homology groups from abstract algebra. A possible reference is [16, Section 17.1].
+First we present the definition of homology groups from abstract algebra. A possible reference is [11, Section 17.1].
 Definition 2.6.
@@
 2.4. Homology for groupoids
-We define homology groups for Haudorff, second countable, ample groupoids, as in [20, 39]. We fix $X,Y$ to be locally compact, Hausdorff, second countable, totally disconnected spaces.
-Here $C_{c}(X,\mathbb{Z})$ denotes the set of continuous and compactly supported functions on $X$ assuming integer values.
+We define homology groups for Hausdorff, second countable, ample groupoids, as in [12, 20]. We fix $X,Y$ to be locally compact, Hausdorff, second countable, totally disconnected spaces.
+Here, $C_{c}(X,\mathbb{Z})$ denotes the set of continuous and compactly supported functions on $X$ assuming integer values.
 Given $f\in C_{c}(X,\mathbb{Z})$ and a local homeomorphism $\sigma:X\rightarrow Y$, define for every $y\in Y$,
@@
 The operations on $H_{0}(G)$ are given by $[f_{1}]+[f_{2}]=[f_{1}+f_{2}]$ and $-[f]=[-f]$.
-The group $H_{0}(G)$ is not necessarily an ordered group because the intesection $H_{0}(G)^{+}\cap(-H_{0}(G)^{+})$ may be different from $\{0\}$. See [39, Remark 3.2].
+The group $H_{0}(G)$ is not necessarily an ordered group because the intesection $H_{0}(G)^{+}\cap(-H_{0}(G)^{+})$ may be different from $\{0\}$. See [20, Remark 3.2].
 Given a homomorphism $\rho:G\rightarrow H$ between étale groupoids, we define $\rho^{(0)}:G^{(0)}\rightarrow H^{(0)}$ by $\rho^{(0)}=\rho|_{G^{(0)}}$.
@@
 $\rho^{(n)}(g_{1},\ldots,g_{n})=(\rho(g_{1}),\ldots,\rho(g_{n})),\text{\hskip 10.0ptfor $(g_{1},\ldots,g_{n})\in G^{(n)}$.}$
-We prove the lemma below, stated in [39, Subsection 3.2] for étale groupoids.
+We prove the lemma below, stated in [20, Subsection 3.2] for étale groupoids.
 Lemma 2.10.
@@
 Proposition 2.11.
-Let $\rho:\rightarrow H$ be a homomorphism that is also a local homeomorphism. Then, for each $n\geq 0$, $\rho$ induces a homomorphism $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ defined by
+Let $\rho:G\rightarrow H$ be a homomorphism that is also a local homeomorphism. Then, for each $n\geq 0$, $\rho$ induces a homomorphism $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ defined by
 $H_{n}(\rho)([f])=[\rho_{\ast}^{(n)}(f)]\hskip 7.0pt\text{for }f\in C_{c}(G^{(n)},\mathbb{Z}).$
 Remark 2.12.
-In this paper, for a function $\tau:C_{c}(G^{(0)},\mathbb{Z})\rightarrow C_{c}(H^{(0)},\mathbb{Z})$ we use the notation $[\tau]$ to denote the function $[\tau]:H_{0}(G)\rightarrow H_{0}(H)$ given by $[\tau]([f])=[\tau(f)]$.
+In this paper, for a function $\tau:C_{c}(G^{(0)},\mathbb{Z})\rightarrow C_{c}(H^{(0)},\mathbb{Z})$ such that $f_{1}\sim f_{2}$ in $H_{0}(G)$ imples $\tau(f_{1})\sim\tau(f_{2})$ in $H_{0}(H)$, we use the notation $[\tau]$ to denote the function $[\tau]:H_{0}(G)\rightarrow H_{0}(H)$ given by $[\tau]([f])=[\tau(f)]$.
 Moreover, for a homomorphism $\rho:G\rightarrow H$ that is also a local homeomorphism, we have $H_{0}(\rho)=[\rho_{\ast}^{(0)}]$.
@@
 Proposition 2.14.
-[39, Proposition 3.5] Let $n\geq 1$. If $\rho,\sigma:G\rightarrow H$ are similar homomorphisms and local homeomorphisms, then $H_{n}(\rho)=H_{n}(\sigma)$.
+[20, Proposition 3.5] Let $n\geq 1$. If $\rho,\sigma:G\rightarrow H$ are similar homomorphisms and local homeomorphisms, then $H_{n}(\rho)=H_{n}(\sigma)$.
 Corollary 2.15.
-[39, Proposition 3.5] If $G$ and $H$ are homologically similar, then they have isomorphic homology groups.
+[20, Proposition 3.5] If $G$ and $H$ are homologically similar, then they have isomorphic homology groups.
 In particular, if $n\geq 1$, $\rho:G\rightarrow H$ and $\sigma:H\rightarrow G$ are local homeomorphisms and homomorphisms satisfying (2.4), then $H_{n}(\rho):H_{n}(G)\rightarrow H_{n}(H)$ is an isomorphism with inverse $H_{n}(\sigma)$.
 Moreover, both $H_{0}(\rho)$ and $H_{0}(\sigma)$ preserve positive elements.
-4.3. Homology for AF groupoids
+3. Homology for AF groupoids
+In this section, we study the properties of the homology groups of AF groupoids.
 Like K-theory, homology groups are continuous with respect to inductive limits.
@@
 ∎
-5. K-theory for AF Groupoids
-Here we extend the isomorphism [20, Corollary 5.2] between the K-theory and the homology group for AF groupoids, showing that the map preserves positive elements, and giving a formula for the isomorphism.
+4. Main result
+Here we extend the isomorphism [12, Corollary 5.2] between the K-theory and the homology group for AF groupoids, showing that the map preserves positive elements, and giving a formula for the isomorphism.
 Although it is often difficult to find the K-theory of an arbitrary C*-algebra, finding the K-theory of $M_{n}$ is a easy.
-For this case, we get the isomorphism $K_{0}(M_{n})\rightarrow\mathbb{Z}$ given by $[p]_{0}\mapsto\mathrm{tr}\hphantom{.}p$, where $\mathrm{tr}\hphantom{.}p$ is the matrix trace of $p\in\mathcal{P}_{\infty}(M_{n})$.
+In this case, we have the isomorphism $K_{0}(M_{n})\rightarrow\mathbb{Z}$ given by $[p]_{0}\mapsto\mathrm{tr}\hphantom{.}p$, where $\mathrm{tr}\hphantom{.}p$ is the matrix trace of $p\in\mathcal{P}_{\infty}(M_{n})$.
 Similarly, for a locally compact, Hausdorff, second countable, totally disconnected space $X$, the K-theory of $C_{0}(X,M_{n})$ is also described in terms of a trace operator.
-To find the K-theory of an AF groupoid, we first consider the building blocks of the groupoid.
+To find the K-theory of the C*-algebra of an AF groupoid, we first consider the building blocks of the groupoid.
 In other words, we consider an elementary groupoid of the form $R(\sigma)$.
-We show that $R(\sigma)$ is locally approximated by groupoids of the form $C_{0}(V,M_{n})$.
+We show that $C^{*}(R(\sigma))$ is locally approximated by groupoids of the form $C_{0}(V,M_{n})$.
 Then we apply approximation to find a formula for the $K_{0}$-group of $C^{*}(R(\sigma))$ and an isomorphism from this group to the corresponding homology group.
@@
 Let $p\in C_{0}(X,M_{n})$ be a projection.Let $K\subset X_{\varphi}$ be a compact subset such that $\|p(x)\|<1$ for every $x\notin K$.
-It follows from [51, Propositions 2.2.4 and 2.2.7] that $\mathrm{tr}\hphantom{.}p(x)=0$ for all $x\notin K$.
+It follows from [34, Propositions 2.2.4 and 2.2.7] that $\mathrm{tr}\hphantom{.}p(x)=0$ for all $x\notin K$.
 Then $\mathrm{tr}\hphantom{.}p\in C_{c}(X_{\varphi},\mathbb{N})$ and $p\in C_{c}(X,\mathbb{N})$.
@@
 Also, $\widetilde{p}=v^{*}\widetilde{q}v$ and $v^{*}v=vv^{*}=\mathrm{id}1_{K}$.
-Considering the unital C*-algebra $C(K,M_{n})\subset C_{0}(X,M_{n})$, it follows from [51, Propositions 2.2.4 and 2.2.7] that there exists $u_{p},u_{q}\in C_{0}(X,M_{n})$ with support in $K$ such that
+Considering the unital C*-algebra $C(K,M_{n})\subset C_{0}(X,M_{n})$, it follows from [34, Propositions 2.2.4 and 2.2.7] that there exists $u_{p},u_{q}\in C_{0}(X,M_{n})$ with support in $K$ such that
 […]
@@
     $\mathrm{tr}\hphantom{.}\mu(f)(x)=\sum_{u:\sigma(u)=\sigma(x)}f(u),$
-    where we apply Renualt’s $j$ map [48, Proposition II.4.2] to identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$.
+    where we apply Renualt’s $j$ map [30, Proposition II.4.2] to identify $C^{*}(R(\sigma))$ as a subset of $C_{0}(R(\sigma))$.
   2. (ii)
@@
   • $\mu$ is an isometry
-    By [55, Proposition 9.3.1], for all $x\in X$, the $\ast$-representation $\pi_{x}:C_{c}(R(\sigma))\rightarrow B(\ell^{2}(R(\sigma)_{x}))$ is given by
+    By [40, Proposition 9.3.1], for all $x\in X$, the $\ast$-representation $\pi_{x}:C_{c}(R(\sigma))\rightarrow B(\ell^{2}(R(\sigma)_{x}))$ is given by
     $\pi_{x}(f)\delta_{(v,x)}=\sum_{(u,v)\in R(\sigma)}f(u,v)\delta_{(u,x)}=\sum_{u:\sigma(u)=\sigma(x)}f(u,v)\delta_{(u,x)},$
@@
     Given $x\in X$, let $u=\varphi(\sigma(x))\in X_{\varphi}$.
-[55, Proposition 9.3.1] implies that there exists a unitary $U_{(x,u)}:\ell^{2}(R(\sigma)_{u})\rightarrow\ell^{2}(R(\sigma)_{x})$ such that $\pi_{x}=U_{(x,u)}\pi_{u}U_{(x,u)}^{*}$.
+[40, Proposition 9.3.1] implies that there exists a unitary $U_{(x,u)}:\ell^{2}(R(\sigma)_{u})\rightarrow\ell^{2}(R(\sigma)_{x})$ such that $\pi_{x}=U_{(x,u)}\pi_{u}U_{(x,u)}^{*}$.
 Then $\|\pi_{x}(f)\|=\|\pi_{u}(f)\|$.
@@
 The following theorem gives an isomorphism of the ordered groups $K_{0}(C^{*}(G))$ and $H_{0}(G)$.
-Corollary 5.2 by Farsi, Kumjian, Pask and Sims [20] and Matui’s [39, Theorem 4.10] prove isomorphisms of homology groups of groupoids and the corresponding $K_{0}$-group.
-However, [20, Corollary 5.2] does not prove whether such isomorphism preseves positive elements from one group to another, and Matui only considers groupoids with compact unit spaces, without providing a detailed proof.
+Corollary 5.2 by Farsi, Kumjian, Pask and Sims [12] and Matui’s [20, Theorem 4.10] prove isomorphisms of homology groups of groupoids and the corresponding $K_{0}$-group.
+However, [12, Corollary 5.2] does not prove whether such isomorphism preseves positive elements from one group to another, and Matui only considers groupoids with compact unit spaces, without providing a detailed proof.
 Our theorem generalises both results, and we describe the isomorphism by an explicit formula.
@@
   3. (3)
-    the isomorphism $\nu$ is precisely the map given by [20, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
+    the isomorphism $\nu$ is precisely the map given by [12, Corollary 5.2]. In other words, $\nu([1_{V}]_{0})=[1_{V}]$ for all $V\subset G^{(0)}$ compact open.
 We divide the proof into three parts:
@@
 ∎
-6. The Commutative Diagram
+5. Application to Deaconu-Renault groupoids
+Here, we apply the ordered group isomorphism from Theorem 1 to characterise the AF embeddability of the C*-algebra of a Deaconu-Renault groupoid.
+First, we define this class of groupoid in Subsection 5.1, then we prove the commutative diagram 1.1 in Section 5.3, which we used in the proof of Theorem 2 in Subsection 5.4.
+5.1. Deaconu-Renault groupoids
+Let $X$ be a locally compact, Hausdorff, second countable space, and $\sigma:X\rightarrow X$ is a surjective local homeomorphism. We define the set
+$\mathcal{G}=\{(x,k,y)\in X\times\mathbb{Z}\times X:\exists m,n\in\mathbb{N},k=m-n,\sigma^{m}(x)=\sigma^{n}(y)\},$
+and we endow it with the range and source maps $r,s:\mathcal{G}\rightarrow\mathcal{G}$ with $r(x,k,y)=(x,0,x)$ and $s(x,k,y)=(y,0,y)$.
+Let $\mathcal{G}^{(2)}=\{((x,k_{1},y),(y,k_{2},z)):(x,k_{1},y),(y,k_{2},z)\in\mathcal{G}\}$ and we define the product $\mathcal{G}^{(2)}\rightarrow\mathcal{G}$ and inverse operations $\mathcal{G}\rightarrow\mathcal{G}$ by
+$(x,k_{1},y)(y,k_{2},z)=(x,k_{1}+k_{2},z),\text{\hskip 10.0pt}(x,k,y)^{-1}=(y,-k,x).$
+With these operations, $\mathcal{G}$ becomes a groupoid.
+The unit space is $\mathcal{G}^{(0)}=\{(x,0,x):x\in X\}$, which is identified with $X$ by the map $(x,0,x)\mapsto x$.
+For the open subsets $A,B\subset X$ and for the natural numbers $m,n$ such that $\sigma|_{A}^{m}$ and $\sigma|_{B}^{n}$ are injective with $\sigma^{m}(A)=\sigma^{n}(B)$, define the subset
+$\mathcal{U}_{A,B}^{m,n}=\{(x,m-n,y)\in A\times\mathbb{Z}\times B:\sigma^{m}(x)=\sigma^{n}(y)\}.$
+Here we use the notation from [42].
+It follows from [9, Theorem 1] that the sets $\mathcal{U}_{A,B}^{m,n}$ generate a topology that makes $\mathcal{G}$ a topological groupoids which is locally compact, Hausdorff, second countable, étale.
+We denote the groupoid $\mathcal{G}$ endowed with this topology a Deaconu-Renault groupoid.
+The first example of this groupoid was described in Renault’s book [30, Section III.2] by a class of groupoids whose C*-algebras are Cuntz algebras, where $X$ is the set of sequences $(x_{i})_{i\in\mathbb{N}}$, with $x_{i}\in\{0,1,\dots,n\}$ (or $x_{i}\in\mathbb{N}$ for the Cuntz algebra $\mathcal{O}_{\infty}$), and $\sigma$ was given by $\sigma(x_{0}x_{1}x_{2}\dots)=x_{1}x_{2}\dots$ Deaconu [9] generalises Renault’s definition by considering compact Hausdorff spaces $X$ and assumes that $\sigma:X\rightarrow X$ is a covering map.
+Under these assumptions, he proved that when $\sigma$ is a homeomorphism, then $C^{*}(\mathcal{G})$ is isomorphic to the crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$.
+Anantharaman-Delaroche [1, Example 1.2 (c)] generalised it to the case where $\sigma$ is a surjective local homeomorphism, and $X$ is locally compact.
+She also gave a sufficient conditions to make the C*-algebra of this groupoid purely infinite [1, Propositions 4.2 and 4.3].
+Then Renault [31] generalised this definition by considering a map $\sigma$ that is defined only on an open subset of $X$.
+For the purpose of our paper, here we consider the definition by Anantharaman-Delaroche, and we impose the additional conditions on the space $X$ to be Hausdorff and second countable.
+By [31, Proposition 2.9], the groupoid is also amenable.
+Note that there are different notions of amenability for groupoids: topological amenable and measurewise amenable.
+By Theorem 3.3.7 of [2], these two notions coincide for groupoids that are locally compact, Hausdorff and étale.
+In this case, [2, Proposition 6.1.8] implies that the reduced and full C*-algebras of the groupoid coincide.
+Remark 5.1.
+We use the notation $c^{-1}(0)$ for the groupoid $\bigcup_{n=1}^{\infty}R(\sigma^{n})$ in Example 2.5 because of the following.
+If $\mathcal{G}$ is the Deaconu-Renault groupoid and $c:\mathcal{G}\rightarrow\mathbb{Z}$ is given by $c(x,k,y)=k$ (in this case, $c$ is called the canonical cocycle), then $c^{-1}(0)$ is isomorphic to the subgroupoid $\{(x,0,y)\in\mathcal{G}\}$ of $\mathcal{G}$.
+The isomorphism is defined by $(x,y)\mapsto(x,0,y)$.
+5.2. Skew-products and crossed products
+As explained in the introduction, we use an isomorphism between the C*-algebras a skew-product groupoid and the corresponding crossed product to understand when the C*-algebra of a Deaconu-Renault groupoid $\mathcal{G}$ is AFE.
+This isomorphism allows us to apply the results of Brown’s Theorem (Theorem 1.2).
+In this section, we use Renault’s book [30] as a reference.
+Definition 5.2.
+[30, Definition I.1.6] Let $G$ be a groupoid and $c:G\rightarrow\mathbb{Z}$ a homomorphism. Define the sets
+$G(c)=\{(g,a):g\in G,a\in\mathbb{Z}\}\hskip 10.0pt\text{and}\hskip 10.0ptG(c)^{(0)}=G^{(0)}\times\mathbb{Z}.$
+Define the maps $r,s:G(c)\rightarrow G^{(0)}\times\mathbb{Z}$ by $r(g,a)=(r(g),a)$ and $s(g,a)=(s(g),a+c(g))$. Consider the set
+(5.1) $G(c)^{(2)}=\{((g,a),(h,a+c(g))):(g,h)\in G^{(2)},a\in\mathbb{Z}\},$
+and define the product $G(c)^{(2)}\rightarrow G(c)$ and the inverse $G(c)\rightarrow G(c)$ by
+(5.2) $(g,a)(h,a+c(g))=(gh,a)\hskip 15.0pt\text{and}\hskip 15.0pt(g,a)^{-1}=(g^{-1},a+c(g)),$
+respectively. Then $G(c)$ becomes a groupoid, called a skew-product.
+If $c$ is a continuous cocycle (i.e., if $c:G\rightarrow\mathbb{Z}$ is a continuous homomorphism) and we equip $G(c)$ with the induced topology from $G\times\mathbb{Z}$, then $G(c)$ inherits the usual topogical properties of $G$.
+Recall that we assume that $G$ is locally compact, Hausdorff, second countable and étale.
+Lemma 5.3.
+Equip $G(c)$ with the topology induced from $G\times\mathbb{Z}$.
+Then $G(c)$ is also a locally compact, Hausdorff, second countable, étale groupoid.
+Moreover, if $G$ is totally disconnected, then $G(c)$ is totally disconnected, too.
+If $G$ is measurewise amenable, then $G(c)$ is also measurewise amenable.
+Note that can identify the unity space $G(c)^{(0)}$ with $\mathcal{X}\times\mathbb{Z}$.
+For $f,f_{1},f_{2}\in C_{c}(G(c))$ and $(g,a)\in G\times\mathbb{Z}$, the operations on $C_{c}(G(c))$ are given by
+$(f_{1}\cdot f_{2})(g,a)=\sum_{g_{1}g_{2}=g}f_{1}(g_{1},a)f_{2}(g_{2},a+c(g_{1})),\text{ and }f^{*}(g,a)=\overline{f(g^{-1},a+c(g))}.$
+Assume that $G$ is a locally compact, Hausdorff, second countable, étale, measurewise amenable groupoid, and suppose that $c:G\rightarrow\mathbb{Z}$ is a continuous cocycle.
+We define a dynamical system $\alpha^{c}$ induced by $c$.
+For every $z\in\mathbb{T}$, $f\in C_{c}(G)$ and $g\in G$, we set $\alpha_{z}^{c}(f)(g)=z^{c(g)}f(g)$.
+By Proposition II.5.1 of Renault’s book [30], there exists a unique dynamical system defined by the equation above.
+We make an abuse of notation by denoting this dynamical system $\alpha^{c}:\mathbb{T}\rightarrow\mathrm{Aut}(C^{*}(G))$.
+The following proposition by Renault [30, Proposition II.5.7] gives an isomorphism between the C*-algebra of skew-product groupoids and the corresponding crossed products.
+For an introduction to crossed products, see Chapters 1 and 2 of Dana William’s book [43].
+Proposition 5.4.
+There exists a unique isomorphism $\rho\rtimes u:C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rightarrow C^{*}(G(c))$ such that
+[…]
+for $f\in C_{c}(\mathbb{T},C_{c}(G))$ and $(g,a)\in G(c)$. Moreover, the inverse of $\rho\rtimes u$ is given by
+$(\rho\rtimes u)^{-1}(F)(z)(g)$ $=\frac{1}{2\pi}\sum_{a\in\mathbb{Z}}z^{c(g)+a}F(g,a),$
+for $F\in C_{c}(G(c))$, $z\in\mathbb{T}$, $g\in G$.
+Now we find a formula for the action $\beta$ on $C^{*}(G(c))$ induced by the dual action $\widehat{\alpha}^{c}$.
+We will apply this formula in Chapter 5.4 for Deaconu-Renault groupoids in order to understand in more detail the condition $H_{\beta}\cap K_{0}(B)^{+}=\{0\}$, given by Brown’s theorem.
+Proposition 5.5.
+Let $\rho\rtimes u:C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rightarrow C^{*}(G(c))$ be the isomorphism of Proposition 5.4.
+Then there exists a unique dynamical system $\beta:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G(c)))$ such that
+$\beta(F)(g,a)=F(g,a+1)\hskip 15.0pt\text{for $F\in C_{c}(G(c))$, $(g,a)\in G(c)$,}$
+and $\beta=(\rho\rtimes u)\circ\widehat{\alpha}^{c}\circ(\rho\rtimes u)^{-1}$ for all $n\in\mathbb{Z}$.
+Moreover, $C^{*}(G(c))\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic.
+Proof.
+Let $\widehat{\alpha}^{c}:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T})$ be the dual action, which is given by $\widehat{\alpha}^{c}_{n}(f)(z)=z^{-n}f(z)$, for $f\in C(\mathbb{T},C^{*}(G))$, $z\in\mathbb{T}$, $n\in\mathbb{Z}$.
+For each $n\in\mathbb{Z}$, let $\beta_{n}=(\rho\rtimes u)^{-1}\circ\widehat{\alpha}^{c}_{n}\circ\rho\rtimes u$.
+Since $\rho\rtimes u$ is an isomorphism and $\widehat{\alpha}^{c}$ is an automorphism, we have that $\beta:\mathbb{Z}\rightarrow\mathrm{Aut}(C^{*}(G(c)))$ is a dynamical system.
+Fix $n\in\mathbb{Z}$.
+By continuity of $\beta_{n}$, finding the formula of this automorphism on $C_{c}(G(c))$ is sufficient to determine $\beta_{n}$ uniquely.
+Let $F\in C_{c}(G(c))$ and $(g,a)\in G(c)$.
+Then, Proposition 5.4,
+$\beta_{n}(F)(g,a)$ $=(\rho\rtimes u)\circ\widehat{\alpha^{c}}_{n}\circ(\rho\rtimes u)^{-1}(F)(g,a)$
+[…]
+[…]
+$=\frac{1}{2\pi}\int_{\mathbb{T}}z^{-c(g)-a-n}\sum_{b\in\mathbb{Z}}z^{c(g)+b}F(g,b)dz$
+[…]
+$=F(g,a+n).$
+By definition, we have that $\rho\rtimes u\circ\widehat{\alpha}^{c}=\beta\circ\rho\rtimes u$.
+Then $\rho\rtimes u$ is an equivariant isomorphism from $(C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T},\widehat{\alpha}^{c},\mathbb{Z})$ onto $(C^{*}(G(c)),\beta,\mathbb{Z})$.
+By [43, Lemma 2.65], $C^{*}(G(c))\rtimes_{\beta}\mathbb{Z}$ and $C^{*}(G)\rtimes_{\alpha^{c}}\mathbb{T}\rtimes_{\widehat{\alpha}^{c}}\mathbb{Z}$ are isomorphic.
+∎
+5.3. The Commutative Diagram
 Recall from the introduction that we want to find a group which is isomorphic to $K_{0}(B)$ and more manageable, where $\mathcal{G}$ is a Deaconu-Renault groupoid, $B$ is the C*-algebra of the skew product groupoid $\mathcal{G}(c)$, and $c:\mathcal{G}\rightarrow\mathbb{Z}$ is the continuous cocycle given by $c(x,k,y)=k$, called the canonical cocycle.
@@
 […]
-6.1. Part 1 of the diagram
+5.3.1. Part 1 of the diagram
 First we define the map $\widetilde{\beta}$ of the diagram.
@@
 Therefore the result holds. ∎
-6.2. Part 2 of the diagram
+5.3.2. Part 2 of the diagram
 In the second part of the diagram, we have an isomorphism between homology groups of different groupoids.
 We will obtain this isomorphism by applying the techniques of homological similarity of groupoids.
-Let us define the groupoid $c^{-1}(0)$.
 Until the rest of this section, we fix $\varphi:X\rightarrow X$ be a continuous section of the surjective local homeomorphism $\sigma:X\rightarrow X$ that defines $\mathcal{G}$.
@@
 This implies that $[\sigma_{\ast}]=H_{0}(\rho)[\widetilde{\beta}]H_{0}(\eta)$. Therefore, the diagram commutes. ∎
-7. Main Result
-The main result of this paper is Theorem 1, which gives a condition on the map $\sigma:X\rightarrow X$, characterising when the corresponding groupoid C*-algebra $C^{*}(\mathcal{G})$ is AF embeddable.
-We prove this result by applying the commutative diagram of Section 4 in combination with Theorem 1.1.
+5.4. AF embeddability for Deaconu-Renault groupoids
+Here we show Theorem 2, which gives a condition on the map $\sigma:X\rightarrow X$, characterising when the corresponding groupoid C*-algebra $C^{*}(\mathcal{G})$ is AF embeddable.
+We prove this result by applying the commutative diagram of Section 5.3 in combination with Theorem 1.2.
 We begin this section by defining $\sigma_{\ast}(F)$ for functions $F\in C_{c}(c^{-1}(0),\mathbb{Z})$, using the fact that $c^{-1}(0)$ is the increasing union of the sets $R(\sigma^{n})$.
@@
 Let $c:\mathcal{G}\rightarrow\mathbb{Z}$ be the continuous cocycle given by $c(x,k,y)=k$ for $(x,k,y)\in\mathcal{G}$.
-By [20, Corollary 5.2], we have that $\mathcal{G}(c)$ is an AF groupoid, and therefore the C*-algebra $C^{*}(G(c))$ is AF.
+By [12, Corollary 5.2], we have that $\mathcal{G}(c)$ is an AF groupoid, and therefore the C*-algebra $C^{*}(G(c))$ is AF.
 For simplicity of notation, we denote this C*-algebra by $B$.
@@
 For an introduction to graph algebras, see the book [29] by Raeburn.
-Given a row-finite graph with no sources, [33, Theorem 4.2] gives an isomorphism from the graph algebra $C^{*}(E)$ to the C*-algebra $C^{*}(G_{E})$ defined by $1_{e}\mapsto 1_{Z(e,r(e))}$, where
+Given a row-finite graph with no sources, [19, Theorem 4.2] gives an isomorphism from the graph algebra $C^{*}(E)$ to the C*-algebra $C^{*}(G_{E})$ defined by $1_{e}\mapsto 1_{Z(e,r(e))}$, where
 $G_{E}=\{(x_{0}z,|x_{0}|-|y_{0}|,y_{0}z):x_{0},y_{0}\in E^{\ast},z\in E^{\infty}\},$
@@
 and the topology on $E^{\infty}$ is generated by the cylinders $Z(\mu)$ for $\mu\in E^{*}$, where each $Z(\mu)$ is the set of all infinite sequences starting with $\mu$.
-The groupoid is almost identical to the groupoid in [33, Definition 2.3], but we made some changes to make it compatible with the notation in Raeburn’s book [46].
+The groupoid is almost identical to the groupoid in [19, Definition 2.3], but we made some changes to make it compatible with the notation in Raeburn’s book [29].
 In constrast to our definition, in [19] two consecutive edges $x_{i},x_{i+1}$ satisfy $r(x_{i})=s(x_{i+1})$.
@@
 By (6.1), $f(y)=0$ for all $y\neq z_{1}$ such that $\sigma^{n}(y)=z_{0}=\sigma^{n}(x)$.
-Since $z_{1}$ satisfies condition $(2)$ of 8.1 for 00 and $p$, we have that $x\neq z_{1}$.
+Since $z_{1}$ satisfies condition $(2)$ of 6.1 for 00 and $p$, we have that $x\neq z_{1}$.
 Therefore $f(x)=0$.
@@
 By Theorem 2, we have the equivalence $(i)\Leftrightarrow(ii)\Leftrightarrow(iii)\Leftrightarrow(v)$.
-The implication $(iii)\Rightarrow(iv)$ is straightforward, and $(iv)\Rightarrow(v)$ follows from the same argument used in [54, Theorem 1.1].
+The implication $(iii)\Rightarrow(iv)$ is straightforward, and $(iv)\Rightarrow(v)$ follows from the same argument used in [38, Theorem 1.1].
 ∎
 6.2. The crossed product $C(X)\rtimes_{\sigma}\mathbb{Z}$
-Now we apply Theorem 1 to show that it generalises Pimsner’s theorem [45, Theorem 9] (see also [9, Theorem 11.5] by Brown).
+Now we apply Theorem 2 to show that it generalises Pimsner’s theorem [26, Theorem 9] (see also [5, Theorem 11.5] by Brown).
 In contrast to these references, we do not assume that $X$ is a compact metric space.
@@
 For our purposes, both definitions are equivalent when $X$ is assumed to be compact.
-Now we show that Theorem 1 can be applied to prove [45, Theorem 9] (or [9, Theorem 11.5]).
+Now we show that Theorem 2 can be applied to prove [26, Theorem 9] (or [5, Theorem 11.5]).
 Corollary 6.9.
@@
 Theorem 6.12.
-[52, Theorem 6.7] if $E$ is a compact topological graph with no sinks, then the following are equivalent:
+[36, Theorem 6.7] if $E$ is a compact topological graph with no sinks, then the following are equivalent:
   1. (i)
@@
     $r$ is injective and for every $v\in E^{0}$ and $\varepsilon>0$, there exists an $\varepsilon$-pseudoloop in $E$ based at $v$.
-Note that, for a Deaconu-Renault groupoid $\mathcal{G}$ corresponding to $\sigma:X\rightarrow X$, then $E=(E^{0},E^{1},s,r)\coloneqq(X,X,\sigma,\mathrm{id})$ is a topological graph with $C^{*}(E)\cong C^{*}(\mathcal{G})$ by [27, Proposition 10.9].
+Note that, for a Deaconu-Renault groupoid $\mathcal{G}$ corresponding to $\sigma:X\rightarrow X$, then $E=(E^{0},E^{1},s,r)\coloneqq(X,X,\sigma,\mathrm{id})$ is a topological graph with $C^{*}(E)\cong C^{*}(\mathcal{G})$ by [15, Proposition 10.9].
 For Deaconu-Renault groupoids, we prove the equivalence of items $(i),(ii),(iii)$ and $(v)$ in Theorem 6.12.
@@
 Then $\Omega$ is compact open because $X$ is compact and totally disconnected.
-By 8.4, we have $\sigma(\Omega)\subset\Omega$ and $\sigma(\Omega)\cap V=\emptyset$.
+By 6.4, we have $\sigma(\Omega)\subset\Omega$ and $\sigma(\Omega)\cap V=\emptyset$.
 So $\sigma(\Omega)\subsetneq\Omega$.
@@
 We believe this is an open question.
+6.4. Cuntz-Pimsner algebras
+Cuntz-Pimsner algebras form a large class of C*-algebras that includes, among others, the C*-algebra associated with a Deaconu-Renault groupoid $\mathcal{G}$.
+By applying the techniques described in [33], one can show that $C^{*}(\mathcal{G})$ is a Cuntz-Pimsner algebra, for which Schafhauser characterised the AF embeddability in [39, Theorem C].
+We refer to Chapter 8 of [29] for an introduction to Cuntz-Pimsner algebras.
+Definition 6.16.
+Let $G$ be a locally compact, Hausdorff, second countable, étale groupoid, and let $c:G\rightarrow\mathbb{Z}$ be a continuous cocycle.
+We say that $c$ is unperforated if, for every $g\in G$ such that $c(g)=n>0$, there exist composable $g_{1},g_{2},\dots,g_{n}$ such that each $c(g_{i})=1$ and $g=g_{1}g_{2}\dots g_{n}$.
+Lemma 6.17.
+The canonical cocycle $c:\mathcal{G}\rightarrow\mathbb{Z}$ on the Deaconu-Renault groupoid $\mathcal{G}$ is unperforated.
+Proof.
+Let $(x,k,y)\in\mathcal{G}$ such that $c(x,k,y)=k>0$.
+Then there exists an $m$ such that $\sigma^{m}+k(x)=\sigma^{m}(y)$.
+Since $\sigma$ is surjective, we can choose $z\in X$ with $\sigma^{k-1}(z)=y$.
+This implies that
+$\sigma^{m}+k(x)=\sigma^{m}(y)=\sigma^{m+k-1}(z).$
+This implies that $(x,1,z)\in\mathcal{G}$ and $c(x,1,z)=1$.
+Moreover, for $i=1,\dots,k-1$, we have $(\sigma^{i-1}(z),1,\sigma^{i}(z))\in\mathcal{G}$ and $c(\sigma^{i-1}(z),1,\sigma^{i}(z))=1$, where $\sigma^{0}(z)=z$.
+Hence,
+$(x,k,y)=(x,1,z)(z,1,\sigma(z))\dots(\sigma^{i-3}(z),1,\sigma^{i-2}(z))(\sigma^{i-2}(z),1,y).$
+Therefore, $c$ is unperforated. ∎
+We will apply the following theorem by Rennie, Robertson, and Sims [33] to write $C^{*}(\mathcal{G})$ as a Cuntz-Pimsner algebra.
+Theorem 6.18.
+[33, Lemma 4 and Theorem 11] Suppose $G$ is a locally compact, Hausdorff, second countable, étale groupoid, and that $c:G\rightarrow\mathbb{Z}$ is an unperforated continuous cocycle.
+Let $X(G)$ be the completion in $C_{r}^{*}(G)$ of the set
+$\{f\in C_{c}(G):\mathrm{supp}\hphantom{.}f\subset c^{-1}(1)\}.$
+Then the inclusion $I_{0}:C_{c}(c^{-1}(0))\rightarrow C_{c}(G)$ extends to an embedding $I_{0}:C_{r}^{*}(c^{-1}(0))\rightarrow C_{r}^{*}(G)$ and the inclusion $I_{1}:C_{c}(c^{-1}(1))\rightarrow C_{c}(G)$ extends to a linear map $I_{1}:X(G)\rightarrow C_{r}^{*}(G)$.
+The set $X(G)$ is a C*-correspondence over $C_{r}^{*}(G)$ with module actions
+$a\cdot\xi=I_{0}(a)\xi\hskip 15.0pt\text{and}\hskip 15.0pt\xi\cdot a=\xi I_{0}(a)\hskip 15.0pt\text{for $\xi\in X(G)$, $a\in C_{r}^{*}(c^{-1}(0))$.}$
+Moreover, the pair $(I_{1},I_{0})$ is a Cuntz-Pimsner covariant representation of $X(G)$, and the integrated form $I_{1}\times I_{0}$ is an isomorphism of the Cuntz-Pimsner algebra $\mathcal{O}_{C^{*}_{r}(c^{-1}(0))}(X(G))$ onto $C_{r}^{*}(G)$.
+Remark 6.19.
+We make a change in notation, writing $\mathcal{O}_{C^{*}_{r}(c^{-1}(0))}(X(G))$ instead of $\mathcal{O}_{X(G)}$ to maintain consistency with the notation used in Schafhauser’s paper [39].
+Now we write $C^{*}(\mathcal{G})$ as a Cuntz-Pimsner algebra.
+Corollary 6.20.
+The C*-algebra $C^{*}(\mathcal{G})$ is isomorphic to $\mathcal{O}_{C^{*}(c^{-1}(0))}(X(\mathcal{G}))$.
+Proof.
+Since the canonical cocyle $c:\mathcal{G}\rightarrow\mathbb{Z}$ is unperforated, Theorem 6.18 implies that $C^{*}_{r}(\mathcal{G})$ is isomorphic to $\mathcal{O}_{C_{r}^{*}(c^{-1}(0))}(X(\mathcal{G}))$.
+The groupoids $\mathcal{G}$ and $c^{-1}(0)$ are amenable, then $\mathcal{O}_{C_{r}^{*}(c^{-1}(0))}(X(\mathcal{G}))=\mathcal{O}_{C_{*}(c^{-1}(0))}(X(\mathcal{G}))$ and $C^{*}_{r}(\mathcal{G})=C^{*}(\mathcal{G})$.
+∎
+Proposition 6.21.
+[39, Proposition 4.3] Suppose $A$ and $B$ are C*-algebras and $H$ is a row-finite $A\text{--}B$ C*-correspondence.
+If $T$ is an $A$-Ferdholm operator, then $T\otimes 1_{H}$ is a $B$-Fredholm operator and the map $T\mapsto T\otimes 1_{H}$ induces a positive group homomorphism $[H]:K_{0}(A)\rightarrow K_{0}(B)$.
+The following theorem by Schafhauser characterises the AF embeddability for a large class of Cuntz-Pimsner algebras.
+Theorem 6.22.
+[39, Theorem C] Suppose $A$ is an AF algebra and $H$ is a separable C*-correspondence over $A$. Then the following are equivalent:
+  1. (1)
+    $\mathcal{O}_{A}(H)$ is AF embeddable,
+  2. (2)
+    $\mathcal{O}_{A}(H)$ is quasidiagonal,
+  3. (3)
+    $\mathcal{O}_{A}(H)$ is stably finite.
+Moreover, if $H$ is row-finite and faithful, the above conditions are equivalent to
+  1. (4)
+    if $x\in K_{0}(A)$ and $[H](x)\leq x$, then $[H](x)=x$.
+Note that, when $X$ is totally disconnected, the C*-algebra $C^{*}(c^{-1}(0))$ is AF.
+Hence, for the C*-algebra $C^{*}(\mathcal{G})=\mathcal{O}_{C^{*}(c^{-1}(0))}(X(\mathcal{G}))$, conditions (1)–(3) are equivalent.
+It is not clear to the author, whether $X(\mathcal{G})$ is row-finite or faithful.